A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 7200–7249

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.02

Title word cross-reference

1 [611]. 2 [122, 59]. 3 [666, 845, 613, 399, 122, 609, 861, 665, 80]. 3 × 3 [193]. _{2} [792]. AX = B [58]. H_{1} [272]. D^{2} [808]. η_{T} [617]. F [564]. H [660, 728]. K [72, 705, 199, 203, 148, 170, 195]. M [114]. N [611, 610]. ω [35]. R [732]. SGBF [112]. T [610].

-Bounded [564]. -Buffer [728]. -Clauses [272]. -Nearest [170, 195]. -NN [199, 203]. -Optimal [148]. -out-of- [611, 610]. -Regular [35]. -Resilient [72]. -Tree [732, 660].

0.13um [445].

10GE [40]. 10th [925, 928]. 11th [920]. 12-DOF [462]. 12th [927]. 14th [919]. 15th [900, 926, 899]. 16th [888]. 17th [921, 922, 923]. 18th [905, 901].

2.0 [414]. 21st [894, 897, 898, 911]. 2T1R [400].

30th [895, 896]. 34th [910].

4th [912, 924].

5th [913].

60th [915].

7th [895, 896].

8-Bit [856]. 8th [917].

9798 [311]. 9th [904, 893, 909, 890, 891].

Abductive [155]. Aberto [773]. Abstract [267, 512, 228]. Abstraction [37, 354, 203]. Accelerating [854]. Accelerator [663]. Access [742, 612]. Accompanying [52]. according [132]. Accuracy [686, 665]. Accurate [736]. Acids [460]. Acquisition [767]. across [91, 373]. ACSL [524]. Action [903, 525, 129]. Active [868]. Activities [538, 177]. Ad [72, 482, 73]. Ad-Hoc [72, 482]. AdaBoost [63, 64]. AdaBoost-Based [64]. Adaptable [425]. Adaptation [257, 88, 421, 800]. Adaptations [832]. Adapted [167]. Adaptive [887, 742, 605, 249, 639, 199, 873, 217, 728, 610]. Addiction [544]. Address [307]. Adini [76]. Adjectives [376, 378, 382, 374]. Administrator [648]. Adoption [514]. Advanced [216, 921, 922, 923]. Advances [910]. Advantages [187]. Adverbs [762]. Advertising [736]. Aesthetic [845, 856, 857]. Affairs [580]. Affect [387]. Affine [452, 396]. Affordability [332]. Afghanistan [539]. against [734, 605, 434]. Agent [166, 359, 904, 360, 163, 356, 164, 347, 362, 351, 165, 350, 358]. Agent-Programming [351]. Agents [352]. Aggregate [232]. Aggregated [688, 732]. Aggregating [537, 486, 198]. Aggregation [687]. Aggression [176]. Agreement [360, 132]. Agreements [417]. AJAX [644]. Alcoholism [532]. Algebra [234, 81, 524]. Algebraically [560]. Algebras [559]. Algorithm [818, 560, 760, 636, 154, 828, 805, 63, 814, 56, 86, 649, 49, 823, 87, 203, 79, 148, 835, 54, 877, 627, 96, 61, 405, 53]. Algorithms [810, 819, 882, 850, 821, 901, 812, 81, 524, 866, 180, 816, 170, 798, 189, 876]. Alignment [409]. Allocation [806]. Almost [273]. Alternating [86]. Alternative [584, 548]. Alternatives [368]. Alternatives-Based [368]. Ambiguity [383]. American [374]. Amino [460]. Amount [82]. Amplification [797]. Amsterdam [905]. Analysing [306, 876]. Analysis [533, 552, 680, 10, 136, 760, 322, 446, 458, 883, 310, 287, 44, 461, 295, 550, 901, 288, 492, 45, 513, 596, 520, 31, 834, 720, 216, 386, 597, 584, 867, 412, 591, 634, 631, 48, 86, 859, 775, 207, 330, 794, 504, 208, 763, 684, 507, 755, 761, 302, 615, 318, 212, 169, 435, 28]. Analytical [17, 38, 210]. Analyzers [521]. Analyzing [671, 658, 874, 433]. Android [527, 782]. Animal [202]. Animating [844]. Annotation [139]. Anonymity [318]. anonymization [316]. Anonymous [432]. Answer [153, 113]. Answerer [481]. Answering [481]. Ant [793, 798, 172]. Anti [58]. Anti-bisymmetric [58]. AnyOut [701]. Anytime [701, 819]. AODV [295]. Apoptosis [463]. Appearance [691]. Appearance-Order-Based [691]. Append [729]. Append-only [729]. Application [754, 671, 170, 165, 189, 62]. Applications [193, 192, 98, 672, 619, 407, 921, 922, 620, 411, 748, 919, 918, 212, 923, 930]. Applicative [269]. Applied [870, 890, 891]. Applying [258, 418]. Appraisal [188]. Approach [21, 523, 881, 5, 179, 171, 231, 658, 360, 140, 585, 699, 566, 188, 412, 688, 547, 362, 425, 599, 535, 22, 803, 795, 490, 877, 481, 177, 80]. Approaches [8, 181, 757, 899]. Appropriate [476]. Approximate [148]. Approximation [516, 518]. Approximations [294]. April [910, 900, 925, 902, 930, 920, 901, 928, 912, 927, 916, 921, 922, 909, 929, 926, 897, 917, 898, 919, 918, 913, 923, 899]. APWeb [919, 918]. Arbitrary [649]. ARCAMA [666]. ARCAMA- [666]. Architectural [846, 522, 242]. Architecture [685, 184, 12, 690, 95]. Architectures [301, 85, 88, 344]. Area [33]. Argumentation [115, 164]. Arithmetic [526]. ARM [428]. Arrays [179]. Art [855, 857, 929, 848, 849]. ArtDeco [887]. Articulated [399]. Artifact [413, 690]. Artifact-Centric [413, 690]. Artificial [895, 896, 456, 832, 839, 835, 181, 450]. Artistic [847]. Ascent [880]. ASD [38]. ASD-Based [38]. Asia [919]. ASL [372, 365]. Aspect [742, 292]. Aspect-Oriented [292]. Assembly [836]. Assertions [419]. Assessing [132]. Assessments [548]. Assigned [177]. Assignment [751, 791, 825]. Assisted [121]. Associated [548]. Association [201, 636, 596, 651]. Associations [840]. Associative [467]. Associativity [313]. Assumptions [611]. Asynchronous [668, 415, 465, 639]. Athens [893]. ATL [354]. Atoms [60]. Attack [605, 434]. Attacker [314]. Attacking [624]. Attacks [607, 246, 608, 621, 433]. Attribute [583]. Attribution [370]. Auction [360, 661]. Auction-Based [360]. Audio [168, 125, 556, 555]. Audits [424]. Augmented [666]. Augmenting [823]. August [894]. Australia [588]. Authenticated [441]. Authentication [311, 445, 173]. Autism [763]. Auto [61]. Autoimmune [469]. Automata [290, 496, 229, 37, 288, 249, 289, 270, 172, 191]. Automata-Based [496]. Automated [301, 225, 885, 295, 599, 863, 64, 586]. Automatic [268, 136, 453, 778, 568, 823, 176, 490, 798, 206, 761, 139]. Automatically [42, 529, 769]. Automating [567]. Automation [591]. Automaton [711]. Automaton-Based [711]. Autonomic [10, 8, 9, 639]. Availability [27, 246]. AVANTSSAR [301]. Avatars [844]. Average [484]. Aware [666, 416, 15, 675, 741, 298, 707, 16, 423, 710, 195, 418, 650, 652, 284]. Awareness [538].

Back [294, 25, 119, 218, 363, 448, 508, 557, 600, 682, 758, 864, 69]. Backgrounds [59]. Balance [537]. Balancing [93]. Banded [77]. Bandit [873]. Bandit-Based [873]. Barcelona [910, 892]. Barcode [59]. Base [599, 774, 118]. Based [39, 808, 494, 836, 804, 523, 103, 222, 779, 243, 675, 271, 496, 733, 626, 322, 590, 656, 446, 636, 368, 171, 671, 359, 691, 640, 154, 130, 360, 163, 729, 214, 399, 724, 851, 514, 288, 554, 663, 731, 641, 632, 438, 721, 443, 657, 478, 253, 43, 57, 403, 160, 740, 404, 598, 164, 611, 837, 63, 735, 814, 872, 591, 688, 638, 738, 868, 400, 630, 60, 49, 642, 860, 677, 12, 788, 702, 791, 829, 727, 199, 873, 51, 739, 737, 535, 208, 38, 842, 653, 65, 617, 389, 757, 711, 94, 510, 95, 876, 202, 831, 743]. Based [728, 302, 52, 480, 64, 652, 635, 426, 395, 637, 433, 421, 767, 401, 62, 61, 586, 53, 593, 697, 124, 911]. Bases [695]. Basic [29]. Basis [444]. Batch [447]. Battery [34]. Bayes [160]. Bayesian [111, 777, 734, 168, 145, 828]. BCM [41]. BDDs [160]. BDI [347, 352]. Be [855]. Bee [832, 835]. Been [855]. Behavior [533, 597, 743]. Behavioral [230, 399, 913, 391]. Behaviors [334]. Behaviour [384, 876]. Behaviours [227]. Beijing [908, 916]. Being [378]. Belief [381]. Benchmark [813]. Benign [65]. Berth [806]. Best [114, 705, 60]. between [638, 65, 333]. Beyond [143, 548]. BGP [302]. Bi [819]. Bi-objective [819]. Bias [788]. Bicategories [263]. Big [756]. Bilateral [182]. Bilingual [766]. Binary [175]. Binders [278]. Binding [455]. Bioinformatics [928]. Biological [834, 829]. Biologically [840, 929]. Biology [458]. Biometric [552, 553]. Bioprocesses [165]. Birdsongs [852]. Birthday [915]. Bisimulations [269]. bisymmetric [58]. Bit [286, 856]. BLAS [75]. Blastoderm [838]. Blinking [136]. Block [193, 609, 630]. Block-Based [630]. Blocking [81]. Blog [740, 637]. blogs [738]. Board [878]. Bogtrotters [361]. Books [388]. Boolean [112, 293]. Boosting [42, 65]. Botnet [317]. Bound [92]. Bounded [564, 286]. Bounds [287, 29, 210]. Box [752]. BPMN [414, 413]. Brainwave [386]. Branch [92, 214]. Branch-and-Bound [92]. Brazilian [777, 775]. Breakthroughs [623]. Breast [65]. Breeding [202]. Bringing [133]. Broadcast [73, 711]. Brookes [273]. Browser [612]. Brzozowski [560]. Büchi [290, 271]. Bucket [114]. Budapest [924]. Buffer [728, 728]. Building [429, 582]. Built [137]. Built-In [137]. Bullies [547]. Bump [152]. Burst [465]. Busan [921, 922, 923]. Business [9, 343, 107, 251, 409, 690, 41]. BWT [650].

C [463, 751, 524, 182, 220, 212]. C-means [182]. C/ACSL [524]. Cache [81, 727]. CAD [586]. Calcium [466]. Calcium/Calmodulin [466]. Calculator [132]. Calculus [254, 226, 276, 752, 32]. Calibration [96]. Call [226]. Call-by-Need [226]. Calls [102]. Calmodulin [466]. Cambridge [909]. Camellia [607]. Camera [554]. CaMKII [466]. Can [394, 487, 382]. Cancer [837]. Cancers [65]. Candidate [488]. CAP [613]. Capability [166]. Capsules [566]. CAPTCHA [613]. Cardiac [101]. Cardinality [882]. Care [580]. Cartesian [784, 796]. Cascades [459]. Case [494, 2, 309, 523, 667, 469, 31, 33, 164, 391]. Case-Based [164]. Categories [131, 252, 757]. Categorization [175]. Category [567]. Causal [111]. CBRN [538]. CC [897]. CEFP [924]. Cell [463, 457]. Cells [909]. Cellular [594, 191, 878]. Central [924]. Centric [413, 690]. Certain [436]. Certificate [323]. Certification [408, 292]. Cessation [541]. Chain [24]. Chains [444, 513]. Challenges [4, 587]. Chamber [296]. Channel [39]. Chaotic [874, 450, 194]. Characteristics [27, 386]. Characterizations [111, 270]. Chattering [174]. Checked [323]. Checking [150, 46, 298, 356, 354, 293, 525, 291, 510, 241]. Checkonly [254]. Checkpointing [215]. chemical [460]. Chemistry [468]. Child [391]. Children [597, 129, 382, 476, 763]. China [908, 916, 917, 919, 918, 889]. Chinese [663, 642]. Chip [74]. Choice [367, 379, 573, 275]. Cholesky [75]. Choosing [123]. Choreography [844]. Chosen [605, 604]. CHR [504]. CHRiSM [506]. Cipher [609, 434]. Ciphertext [605, 604]. Circle [713]. Circuits [784]. Class [262, 520, 197, 186, 63]. Class-Escape [520]. Class-Modular [520]. Classification [168, 851, 190, 196, 184, 702, 793, 203, 195, 595, 635]. Classifier [63, 199]. Classifier-Based [63]. Classifiers [201, 198]. Classroom [547]. Clauses [254, 272]. Cleaning [772]. Climbing [880]. Clinical [581]. Clones [498]. Closures [566]. Cloud [309, 408, 107, 774, 716, 625, 426]. Clouds [671, 411]. Cloze [778]. CLP [494]. CLP-Based [494]. Cluster [731, 199]. Cluster-Based [731, 199]. Clustered [443]. Clustering [672, 457, 157, 699, 182, 661, 651, 806]. Clusters [679]. Co [560, 885, 200, 400, 60, 505, 792]. Co-evolution [885]. Co-evolutionary [200]. Co-located [400]. Co-Logic [505]. Co-operative [885, 60]. Coalgebra [568]. Coalgebraic [265]. Coalgebras [267]. Coarse [84]. Coating [404]. Code [611]. Code-Based [611]. Coded [54]. Coding [618]. Coevolution [796]. Cognates [766]. Cognition [129]. Coimbra [925]. Collaboration [4]. Collaborative [854, 109, 395]. Collective [85, 876]. College [913]. Colloquium [905]. Colony [832, 835, 172]. Color [590, 403]. Colored [712]. Coloring [815, 576]. Combination [490, 64]. Combinatorial [927]. Combining [71, 268, 883, 200, 138, 807, 572, 187, 381, 260]. ComMapReduce [715]. Commercial [726]. Commitment [348]. Common [774, 696]. Communicating [229, 515]. Communication [715, 893, 331, 85, 872, 326, 334]. Communications [332]. Community [536, 532, 740, 481]. Commutativity [313]. Comonad [266]. Compact [146, 841]. Comparative [882, 854, 51, 541]. Comparing [802, 832, 477, 452, 101]. Comparison [93, 834, 196, 333, 124]. Compartmentalized [838]. Compass [563]. Compatibility [105]. Competitive [196, 362]. Compilation [116]. Compiler [214, 215, 897]. Complementation [271]. Complete [314, 230, 570, 571]. Completeness [569]. Complex [8, 543, 155, 456, 596, 830, 128, 841, 755, 450, 405]. Complexity [561, 574, 507]. Compliance [408]. Complicated [59]. Component [243, 227, 211, 169]. Component-Based [243]. Composed [515]. Composing [850]. Composition [410, 245, 517, 415, 420, 425, 710, 426]. Compositional [227, 522, 299, 259]. Comprehensive [455]. Compression [627]. Compressive [57]. Compressor [100]. Computable [606, 602]. Computation [496, 811, 930, 115, 928, 927, 493, 620, 707, 660, 697]. Computational [900, 305, 606, 852, 548, 455, 925]. Computationally [314, 435]. Computations [108]. Computer [893, 854, 404, 386, 547, 78]. Computer-in-the-Loop [547]. Computing [309, 416, 429, 417, 93, 639, 907, 888, 69, 913]. Concentration [466]. Concept [638]. Conceptual [257, 112, 157, 582]. Concurrent [98, 281, 224, 279, 253, 753, 263, 577]. Conditional [323, 300, 147]. Conditionals [369]. Conditions [529]. Conference [910, 900, 925, 908, 895, 896, 902, 901, 928, 927, 916, 921, 922, 909, 929, 926, 894, 897, 917, 888, 919, 890, 891, 913, 923, 899]. Conferences [900, 902, 901, 897, 898, 899]. ConfiChair [309]. Confidence [660]. Configuration [139]. Conflict [113, 287, 253]. Conflict-Driven [287]. Conformance [241]. Conjunction [372]. Connectivity [67]. Consensus [537, 531]. Consequences [372]. Conservative [456]. Considerations [556, 389]. Considers [714]. Consistency [113, 245, 525, 79]. Consistent [223]. Constraint [537, 562]. Constraints [73, 91, 658, 882, 531, 251, 144]. Construction [901, 897, 809, 212]. Constructions [767]. Consumption [866]. Contact [464, 842]. Container [266]. Contamination [812]. Content [675, 603]. Contest [358]. Context [879, 666, 15, 17, 675, 16, 710, 652, 366, 387]. Context-Aware [666, 15, 675, 16, 710, 652]. Context-Sensitive [387]. Contextual [736]. Contingency [584]. Continue [716]. Continuous [806]. Contract [258]. Contracts [243, 230]. Contrasting [546]. Contribution [734]. Contributors [698]. Control [269, 742, 231, 783, 456, 799, 632, 524, 108, 165, 625, 450, 612, 174]. Controlled [525]. Controller [33, 54]. Controllers [871, 853]. Converge [369]. Convey [133]. Cooperation [830, 318]. Coordinate [669]. Coordination [762, 359, 410]. Coq [518]. cores [100]. Corpora [139]. Corpus [765, 380, 772, 767]. Corpus-Based [767]. Corrected [447]. Correctness [224, 235]. Correlation [811, 632, 638, 546, 61]. Corroded [404]. COST [903, 90, 379, 125, 65, 342, 433]. Cost-Based [433]. Cost-Efficient [342]. Cost-Sensitive [65]. Could [855]. Counter [297]. Counter-Example [297]. Counterexamples [516]. Countries [333]. Coverage [536, 654]. Covering [30, 816]. CPM [61]. CPM-Signal [61]. CPUs [205]. Crack [397]. Crane [804]. Creation [731]. Creative [863]. Credibility [647]. Credibility-Oriented [647]. CRF [641]. CRFs [642]. CRISPR [454]. Criticality [464]. Cross [813, 421]. Cross-Domain [813]. Cross-Layer [421]. Crossovers [786]. Crowd [536, 486, 743]. Crowdsourced [535]. Cryptanalysis [609]. Cryptographic [193, 225]. Cryptography [616]. Cryptosystems [617]. CSP [36]. CSPs [116].

CUDA [101]. Cultural [537, 913, 540]. Cumulative [552]. Curves [315]. Customer [163, 635]. Cut [593]. Cyber [535, 743]. Cyber-Physical [743]. CyberDesign [106]. Cyberinfrastructure [106]. Cyclic [77]. Cyprus [907]. Cytomics [591].

D [666, 845, 613, 399, 122, 609, 861, 59, 665, 80]. DAA [447]. Dance [130]. Dancing [844]. DASFAA [921, 922, 923]. Data [21, 701, 706, 192, 679, 453, 131, 461, 157, 678, 928, 190, 596, 585, 834, 720, 646, 598, 584, 705, 634, 707, 398, 686, 884, 100, 197, 482, 718, 16, 674, 332, 774, 186, 22, 653, 914, 684, 581, 424, 711, 143, 719, 74, 206, 202, 687, 284, 889, 659, 692, 255]. Data-Oriented [22]. Database [123, 727, 921, 922, 923]. Databases [688, 703, 696, 709, 660]. Datasets [316, 839]. Date [791]. DBMS [721]. De-anonymization [316]. Debt [188]. Debugger [497]. Debugging [517]. December [905, 907]. Decentralised [887]. Decentralized [410, 172]. Decidable [221]. Deciding [319, 300]. Decision [724, 158, 146, 187, 184, 251, 186, 842, 581]. Declarative [495, 497]. Declassification [319]. Decline [544]. Decomposition [808, 733, 823, 186]. Dedicated [915]. Defect [404]. Defence [624]. Defense [626, 535]. Defined [575]. Defining [423, 748]. Definite [516]. Definition [137]. Definitions [616]. Deformation [397]. Degradation [179]. Degree [546, 374]. Delay [39, 32]. Delays [33]. Delimited [269, 231]. Delimited-Control [269]. Deliveries [807]. Demand [340, 711]. Dementia [64]. Democratic [543]. Deniable [605]. Denial [246, 433]. Denmark [911]. Density [62]. Deontic [367]. Department [580]. Dependability [888]. Dependence [198]. Dependencies [113, 828, 149]. Dependency [776]. Dependent [31, 29, 86]. Deployment [337]. Deployments [333]. Depression [67]. Derivation [615]. Derivatives [879, 593]. Describe [138]. Description [136, 624]. Descriptive [561]. Descriptors [847]. Design [20, 243, 5, 846, 8, 885, 42, 412, 868, 929, 639, 861, 23, 389, 797, 798, 258, 390]. Design-Driven [5]. Designing [17, 393]. Desired [454]. Detect [498]. Detecting [192, 840, 839, 547]. Detection [701, 543, 436, 598, 837, 211, 434, 631, 787, 176, 841, 476, 653, 291, 59, 61]. Determination [90, 80]. Determines [466]. Deterministic [290, 289]. Determinization [265]. Developing [550, 411, 350]. Development [523, 514, 428]. Device [445]. Devices [614, 812, 868, 109, 728, 589]. Dexter [915, 578]. DFT [888]. Diagonally [77]. Diagrams [262, 116]. Dicionário [773]. Different [802, 55, 467, 197]. Differential [607, 304, 452, 609, 434, 608]. Differentiation [463]. Difficult [474]. Diffie [442]. Diffusion [542, 99, 546, 468]. Digilog [388]. Digital [393, 286, 125, 868, 398, 335, 194, 191]. Dimensional [731, 398, 195, 585, 834, 723]. Dimensions [316]. Directed [500, 149, 877]. Directions [86]. DIS [367]. Disambiguation [779, 483]. Disasters [626]. Disc [592]. Discharge [196]. Discourse [371]. Discovering [680, 832, 629, 655]. Discovery [666, 152, 735, 867, 689, 241]. Discrete [452, 445, 829]. Discrete-Time [829]. Discretization [76, 198, 194].

Discriminating [656]. Discrimination [65]. Discriminative [51]. Discussion [589]. Discussions [556]. Disease [828, 599]. Disease-Related [828]. Disease-Symptom [599]. Disequalities [272]. Dishonest [312]. Disjoint [198]. Disjunction [372]. Disk [714]. Disorder [763]. Display [400]. Disputed [647]. Dissimilar [333]. Distance [811, 790, 653, 576]. Distance-Based [653]. Distances [195]. Distributed [103, 104, 610, 105, 620, 108, 109, 241]. Distribution [812, 646, 653, 181, 869, 687, 172]. Diverse [800, 418]. Diversification-Driven [815]. Diversity [871, 654, 849]. Diversity-Enhancing [871]. DNA [832]. DNF [159]. Do [129, 133]. Document [479, 388]. Documents [486, 488, 651]. DOF [462]. Dogs [394]. Domain [670, 641, 493, 437, 813, 809, 327, 489, 769]. Domain-Specific [493]. Domains [80]. Dominance [808]. Dominant [77]. Double [56]. Down [697]. DQDI [923]. Drawing [853, 577]. Driven [494, 5, 287, 242, 380, 814, 815, 207, 362, 425, 340, 41, 107]. Drosophila [838]. Dual [83, 296, 730]. Dual-Grained [730]. Due [791]. Due-Date [791]. Dumb [283]. Duplicate [627]. Duplication [565, 833]. during [499]. Dynamic [166, 105, 493, 216, 107, 553, 129, 874, 838]. Dynamical [453, 456]. Dynamically [416]. Dynamically-Green [416]. Dynamics [457, 540, 464, 139, 539, 544].

E-health [588]. E-voting [310]. EAP [435]. EAP-TNC [435]. Early [838]. Easy [277]. ECG [720]. Echelon [803]. ECIR [910]. ECMAF [421]. Ecological [795]. Econ [903]. Economic [426]. Economics [325, 903, 339, 336]. Ecosystem [792]. Ecosystems [326, 327]. Eden [750]. Edge [825]. Edutainment [906]. Effect [457, 467, 546]. Effective [394, 517, 638, 466, 270, 708]. Effectiveness [177]. Effects [264, 469, 125, 127, 381, 390, 395, 665]. Efficiency [731, 604]. Efficient [233, 602, 289, 649, 739, 439, 703, 710, 797, 625, 342]. Efficiently [698]. eID [619]. Electric [634]. Electrical [810]. elegans [463]. Elevation [398]. Elicitation [12]. Elimination [837, 87, 174]. Elliptic [315]. Email [473, 490]. Embedded [249, 38]. Embedding [495, 555]. Embodied [893]. Embryos [838]. Emergence [878]. Empirical [133, 529, 876]. Emulation [428]. Enable [517]. Enabled [432]. Enabling [24, 684]. Encephalomyelitis [469]. Encoding [465, 275]. Encrypting [191]. Encryption [446, 604, 313]. Encryptions [605]. End [222, 764]. Ended [875]. Ending [762]. Endpoint [61]. Energy [416, 866]. Energy-Aware [416]. Enforceable [320]. Enforcement [742]. Enforcing [678, 430]. Engagement [539]. Engine [392, 717, 403, 719]. Engineering [416, 407, 106, 831, 889, 899]. English [377, 531]. Enhance [242]. Enhanced [613]. Enhancement [200]. Enhancing [519, 6, 871, 884, 395]. Enough [371]. Enrichment [478]. Ensemble [175, 851, 702, 185, 53]. Ensembles [179, 169]. Enterprise [6, 4, 15]. Enterprises [887, 5, 10, 8, 9]. Entities [482]. Entity [311, 695, 641, 583, 482]. Entity-Attribute-Value [583].

Entries [637]. Entropy [552, 776, 614]. Entropy-Guided [776]. Enumeration [692]. Environment [852, 93, 45, 838]. Environmental [666, 341, 623]. Environments [166, 121, 870, 107, 582]. Enzymatic [459]. Epidemic [598]. Epigenetic [450]. Episode [719]. Epistasis [839]. Epistemic [572]. Equality [570]. Equation [452, 293, 58]. Equational [313]. Equations [101]. Equilibria [280]. Equilibrium [496]. Equirecursive [564]. Equivalence [313, 225, 616, 35]. Equivalences [274]. Equivalent [279]. Erlang [755]. ERP [381]. ERP-Effects [381]. Escape [520]. ESOP [898]. Essays [915]. Established [469]. Estimate [646]. Estimating [67, 716]. Estimation [478, 829, 617, 62]. Estimator [198, 62]. Estonia [900, 902, 901, 897, 898, 899]. ETAPS [900, 902, 901, 897, 898, 899]. EuroGP [926]. European [910, 900, 902, 901, 928, 927, 926, 897, 898, 924, 899, 332, 327, 333]. Evaluation [485, 40, 445, 122, 693, 55, 710, 38, 888]. Evaluative [378]. Event [276, 155, 674, 421]. Event-Based [421]. Events [475, 812, 29, 14, 907]. Eventually [223]. Evidence [376, 380]. EVIS [719]. EvoApplications [930]. EvoBIO [928]. EvoCOMNET [930]. EvoCOMPLEX [930]. EvoCOP [927]. EvoFIN [930]. EvoGAMES [930]. EvoHOT [930]. EvoIASP [930]. Evolution [536, 845, 540, 871, 856, 860, 326, 792, 830, 327, 798, 875, 878, 885]. Evolutionary [818, 810, 846, 930, 882, 850, 852, 565, 855, 800, 814, 872, 866, 857, 868, 824, 884, 861, 797, 849, 881, 200, 928, 927, 929]. Evolvability [821]. Evolved [456]. Evolving [879, 781, 847, 853, 462, 791, 792, 848]. EvoMUSART [929]. EvoNUM [930]. EvoPAR [930]. EvoRISK [930]. EvoSTIM [930]. EvoSTOC [930]. Exact [811, 807, 526]. Examination [539]. Example [297, 556]. Examples [197]. Exchange [792]. Exchanges [532]. Excitatory [467]. Exclusion [439]. Executing [411]. Execution [105, 802, 235, 31, 180, 259, 515]. Existing [42]. Expanded [680]. Expansion [488, 809, 723]. Expectation [352]. Experience [917, 124]. Experimental [561, 93, 469, 384]. Experimentation [20]. Experimenting [375]. Experiments [591]. Expert [626, 171]. Explain [544]. Explaining [471]. Explanatory [159]. Explicit [98, 574]. Exploitation [387]. Exploiting [552, 829, 206, 144]. Exploration [469, 648, 582]. Explorations [42]. Exploring [862, 394, 473, 743]. Expression [97, 453, 860, 176, 838]. Expression-Based [860]. Expressive [158]. Extended [267, 347]. Extending [469, 756, 272]. Extension [629, 437, 635]. Extensions [501, 612]. Extracting [21, 770, 654, 599, 22, 769]. Extraction [766, 641, 161, 768, 14, 13]. Extremal [94]. Extreme [50]. Extrusion [276].

Face [787]. Faces [860]. Facial [176]. Facilitate [394]. Factor [536, 455]. Factoring [604]. Factorization [75]. Factors [545]. Failure [31]. Failure-Dependent [31]. Family [613]. FASE [899]. Fast [290, 693, 78, 203, 719, 692]. Fault [434, 888]. Favorite [656]. FDG [64]. FDG-PET [64]. Feature [836, 656, 200, 776, 554, 837, 258, 202, 64, 66].

Feature-Based [554]. Feature-Oriented [258]. Features [403, 127, 857, 553]. Feedback [125]. Feeds [476]. Feldspar [754]. FEM [88]. Female [539]. Fence [297]. Few [545]. FFTs [85]. Fibrational [264]. Fictional [238]. Field [20, 479, 624]. Fields [734]. Fifth [417]. Filter [396]. Filtering [182, 659]. Filters [286]. Final [568, 661]. Financial [881]. Find [672]. Finding [856, 738, 79]. Fine [102, 612, 215]. Fine-Grain [215]. Fine-Grained [612]. Fingerprinting [555]. Fingertips [685]. Finite [28]. Firing [465]. First [268, 902, 158, 916, 929, 186]. First-Order [158]. Fitness [811, 855, 790, 874]. Fitness-Distance [811]. Flash [733, 729, 727, 732, 728, 730]. Flash-Based [733, 729, 727, 728]. FlashDB [923]. Flexible [630, 439, 428]. Flickr [741, 478]. Flickr-Aware [741]. Flight [174, 826]. Flow [323, 424, 390, 80]. Flow-Oriented [424]. Flows [190, 14]. Flowshop [92]. Fly [16, 74]. FNN [626]. FOAF [549]. Focused [785]. Fold [505]. Foldering [490]. Folksonomies [652]. Force [125]. Force-Feedback [125]. Forecasting [879, 170]. foreign [889]. Forensic [554]. Forests [178]. Forex [884]. Formal [310, 512, 246, 514, 140, 912, 365, 302, 624]. Formalizing [304]. Formally [222]. Formation [464]. Formed [781]. Formulas [112]. FOSSACS [900]. Foundations [347, 220, 900]. Four [306]. Fractal [401]. Fractional [593]. Fragment [836]. Fragments [270]. Frames [118]. Framework [257, 782, 200, 249, 596, 43, 740, 216, 567, 634, 251, 690, 813, 831, 615, 428, 633, 433, 421]. Frameworks [420]. France [375]. Fraudsters [178]. Free [367, 379, 681, 868]. French [136]. Frequency [466, 869]. Frequent [640, 703, 651]. Friend [744, 713]. Front [3, 7, 385, 406, 427, 449, 470, 491, 509, 530, 551, 558, 579, 601, 622, 643, 664, 683, 704, 725, 746, 759, 780, 801, 822, 843, 865, 886, 11, 18, 26, 47, 68, 89, 110, 120, 141, 162, 183, 204, 219, 240, 261, 282, 303, 324, 345, 346, 349, 353, 357, 364]. FTL [730]. Function [90, 102]. Functional [495, 268, 749, 417, 67, 750, 757, 924]. Functional-Logic [495]. Functionally [418]. Functions [811, 438, 437, 615]. Fundamental [883, 899]. Furthest [649]. Fusion [552, 66]. Future [475, 329, 344, 336]. Fuzzy [851, 182, 174].

GA [883, 202, 64]. GA-Based [202]. Gaia [362]. Gaia-Driven [362]. Game [392, 443, 534, 535, 389, 390, 318, 391]. Game-Based [535]. Game-Design [390]. Game-Play [390]. Game-Theoretic [318]. Games [281, 280, 279, 394, 386, 573, 263, 577]. Gas [192, 624]. Gaussian [87, 653]. GC [213]. GC-Safe [213]. Gendered [539]. Gene [453, 451, 565, 452, 838, 831, 595]. General [254, 429, 588, 439]. Generalised [783]. Generalized [94, 189, 62]. Generate [232]. Generated [521]. Generation [494, 136, 778, 134, 776, 553, 599, 338, 401]. Generative [844]. Generator [675]. Generators [78]. GeNet [831]. Genetic [879, 781, 782, 826, 783, 821, 812, 799, 839, 596, 784, 785, 856, 786, 859, 827, 788, 789, 791, 794, 180, 170, 796, 54, 831, 877, 96, 926]. Genome [833]. Genotype [795]. Genre [851]. Geo [478, 713]. Geo-Location [478].

Geo-Social [713]. Geographic [671, 678, 686]. Geographical [920]. Geological [626]. Georeferenced [671]. Germany [888]. Gesture [893, 140, 123, 122, 127, 124, 893]. Gestures [121, 126, 133]. Gettier [370]. GI [888]. GI/ITG [888]. GIS [667]. GKR [892]. Global [811, 27, 735]. Goal [348]. Gold [594]. Good [275]. Goodness [548]. Governance [339, 336]. GPS [673]. GPU [90, 663, 829, 206]. GPU-Based [663, 829]. GPUs [235]. Grain [215]. Grained [102, 730, 612]. Grammar [137, 788]. Grammar/Prosody [137]. Grammars [253, 861]. Grammatical [792, 798]. Graph [113, 154, 503, 115, 438, 253, 500, 699, 688, 49, 815, 861, 831, 255, 892]. Graph-Based [154, 438, 831]. Graphical [114]. Graphics [848]. Graphs [537, 112, 55, 118]. GRASP [809]. Great [531, 894]. Greece [893]. Greedy [816]. Greek [137]. Green [416, 805, 340]. Grid [870, 103, 92, 283, 94]. Grids [96]. Ground [741]. Group [543, 631]. Groups [744]. Growing [192]. Guaranteed [644]. Guarantees [250, 72]. Guardol [285]. Guided [297, 776]. GW [893].

HactarV2 [359]. HAIS [895, 896]. Halt [716]. Hand [399, 123, 139]. Handwriting [553]. Hangzhou [917]. Haptic [125, 400]. Haptic-Visual [400]. Hard [277]. Hardware [802, 621]. Harmony [185]. Hash [444, 438, 437, 553]. Hashing [315, 636]. Haskell [747, 750, 753]. Having [189]. Health [587, 588, 916]. Healthcare [585, 583, 582]. Healthcare-Related [583]. Heap [250, 207]. Heat [99]. Hedonic [124]. Held [900, 902, 901, 897, 898, 899]. Helicopter [783, 174]. Hellman [442]. Help [487]. Hennessy [572]. Heterogeneous [386, 866, 16, 172]. Heterosis [451]. Heuristic [807, 814, 813, 806]. Heuristics [37]. Hidden [401]. Hiding [914]. Hierarchical [670, 88, 74, 96]. Hierarchically [515]. Hierarchy [724]. Hierarchy-Based [724]. High [781, 614, 751, 731, 83, 591, 654, 108, 195, 402]. High-Dimensional [731, 195]. High-Entropy [614]. High-Level [781]. High-Throughput [591]. Highly [760, 718, 838]. Hill [880]. Histograms [152, 686]. History [284]. History-Aware [284]. Hoc [72, 482, 73]. Holistic [676, 321]. Homes [810]. hop [20]. Horizontal [565]. Hospitals [581]. Hotel [170]. Hovering [783]. HP [817]. Hubness [195]. Hubness-Aware [195]. Human [893, 839, 854, 857, 128, 595, 52, 177]. Human-Accompanying [52]. Human-Computer [893, 854]. Human-Motion [128]. Hungary [924]. Hunting [152, 178]. Hybrid [862, 871, 184, 199, 388, 181, 402, 660, 895, 896]. HyFlex [813]. Hyper [814]. Hyper-Heuristic [814]. Hyperparameter [873]. Hyperproperties [321]. Hypotheses [159].

I/O [749, 40]. IBPM [43]. Iconic [133]. Iconicity [365]. ICSOC [907]. ICT [341]. Identification [614, 554, 442, 197, 388, 535, 64, 169]. Identify [538]. Identifying [698]. Identity [446]. Identity-Based [446]. IDP [918]. IEC [311]. IEKB [918]. IFS [401]. II [896, 805, 922, 891]. ILP [894, 156]. ILS [804, 815]. ILS-Based [804]. Image [590, 403, 591, 638, 182, 594, 593].

Images [478, 592, 65, 401, 191]. Imaging [64]. Imbalanced [197]. Immune [839]. Immunochromatographic [594]. Impact [341, 475, 731]. Imperative [781]. Imperfect [577]. Implantable [296]. Implementation [20, 754, 764, 493, 732]. Implementations [101]. Implemented [755]. Implementing [360]. Implications [331, 335, 333, 317]. Implicit [359]. Impossible [607, 608]. Improve [180]. Improved [271, 604, 504, 633]. Improvement [715, 34]. Improvements [433]. Improving [201, 205, 768, 787, 686, 553, 538, 789, 727, 687]. Improvisational [387]. Improviser [863]. In/Out [546]. In/Out-Degree [546]. Inclusion [439]. Incomplete [453, 77]. Incremental [636, 658, 163]. Incrementalization [321]. Indefinite [82]. Independent [256, 603, 696]. Index [731, 711]. Indexes [695]. Indicator [330]. Indicators [332]. Indices [365]. Indifferentiability [437]. Indifferentiable [315]. Indirect [214]. Indistinguishability [606]. Individual [714]. Indoor [668]. Induction [264, 777, 163, 159]. Inductive [150, 153, 155, 894]. Inductive-Abductive [155]. Industrial [6, 330]. Industry [4]. Infer [486]. Inference [453, 145, 142, 160, 148]. Inferring [516, 828, 565, 835]. InfiniBand [43]. Inflective [760]. Influence [546]. Informally [485]. Information [323, 910, 685, 770, 671, 920, 580, 587, 657, 916, 583, 546, 327, 133, 577, 700, 909, 917]. Informational [532]. Informative [656]. Informed [840]. Infra [655]. Infra-skyline [655]. Infrastructure [323, 164]. Infrastructures [887]. Inhabitation [574]. Inheritance [212]. Inhibitor [458]. Inhibitory [467]. Initial [885]. Initialisation [788]. Innovation [5]. Inquisitive [367, 370]. Inquisitiveness [368]. Insensitivity [248]. Insertion [297]. Inspirations [391]. Inspired [929]. Instance [200]. Instances [414, 302]. Institutionalized [394]. Instrumentation [216]. Integrated [762, 585]. Integrating [150, 679, 583, 352]. Integration [140, 16, 259]. Integrity [618, 644]. Integrity-Guaranteed [644]. Intelligence [181]. Intelligent [895, 896, 889, 889]. Intensification [815, 374]. Intensification/Diversification [815]. Intensification/Diversification-Driven [815]. Inter [206]. Inter-thread [206]. Interacting [118]. Interaction [672, 393, 453, 667, 893, 125, 867, 841, 387, 124]. Interactions [30, 538, 842, 459]. Interactive [676, 268, 846, 474, 513, 856, 861, 109]. interchange [889]. Interface [846, 123, 400, 117, 389]. Interfaces [124]. Interleaved [155]. Intermodeling [252]. Internal [712, 377]. International [900, 925, 908, 895, 896, 892, 902, 904, 920, 893, 328, 407, 901, 912, 916, 921, 922, 722, 909, 929, 894, 897, 907, 917, 888, 911, 918, 890, 891, 913, 923, 899]. Internet [44, 167, 339]. Interoperability [242]. Interpolating [590, 792]. Interpolation [679]. Interpretation [512, 228, 380]. Interpreted [354]. Interprocedural [213]. Intersection [574]. Intervention [337, 541]. Intraday [881, 884]. Introduction [1, 417, 331, 325, 326, 344, 336]. INTRUST [908]. Inventories [412]. Inverse [519]. Inversion [209]. Investigating [857]. Investigation [790]. Investment [882]. Invocation [639]. IOV [40]. IPCAT [909]. iPIN [619]. IPTV [39]. IR [910]. IScIDE [889]. ISO [311, 507]. ISO/IEC [311]. ISO605 [903]. ISOcat [131]. ISPEC [917]. Issue [914]. Italy [920]. iTasks [748]. ITEMS [923, 656, 476, 708]. Iterated [814, 816]. Iterative [58]. ITG [888].

JaCaMo [350]. Jason [351]. Java [298, 255]. Job [870, 90, 791, 94]. Jobs [96]. Joint [900, 902, 901, 897, 898, 899]. July [892, 894, 911]. June [924].

Kairos [721]. Kaiserslautern [888]. Kernel [621, 50, 169, 62]. Kernels [75, 206]. Key [322, 403, 396, 441, 617, 615]. Keynote [623]. Keyphrase [654]. Keys [312, 689]. Keyword [698, 709, 692]. Keywords [659]. Kinase [458]. King [375]. Kleene [559]. Kleisli [252]. Knowledge [15, 892, 695, 42, 629, 14, 12, 599, 737, 13, 118, 581, 143, 350, 370, 624]. Korea [921, 922, 923]. Kozen [915, 578]. Kunming [919, 918].

Labanotation [130]. Label [641, 184, 201]. Labeled [116]. Labeling [168, 637]. Labelled [293]. Labels [486]. Lagrange [590]. Lambda [226]. Landscapes [874]. Language [495, 905, 777, 136, 760, 925, 131, 134, 137, 893, 503, 138, 135, 140, 139, 285, 493, 347, 161, 525, 373, 763, 256, 13, 773, 374]. Languages [242, 492, 520, 278, 569, 501, 756, 898]. LARES [45]. Large [27, 316, 772, 630, 535]. Large-Scale [535]. Law [642]. Laws [36]. Layer [421, 54]. LBlock [608]. LDA [700, 637]. LDA-Based [637]. Leaders [640, 334]. Leadership [70]. Leadership-Scale [70]. Learn [394]. Learned [42]. Learners [175]. Learning [536, 151, 393, 287, 154, 155, 776, 156, 142, 928, 415, 854, 646, 196, 188, 146, 161, 55, 702, 149, 863, 50, 172]. Learnt [514]. Least [58]. Length [308, 617]. Lens [533]. Lessons [514, 535]. Leukemia [458]. Level [781, 417, 517, 75]. Level-3 [75]. Leveraging [681]. Lévy [826]. LFG [775]. Libraries [524]. Library [224, 847, 99, 289]. Licensing [377]. Lifecycle [411]. Lifetime [34]. Light [765]. Lightweight [715, 618, 599]. Like [372]. Lince [764]. Line [103, 871, 772, 878]. Linear [81, 524, 77, 795, 507]. Lines [260, 259]. Linguistic [138, 737]. Link [545, 700]. Linked [707, 774]. Linking [442]. Links [549]. Lipreading [173]. List [697]. Lists [308]. Little [756]. Live [867, 535]. Lived [20]. Liveness [73, 173]. Living [121]. Load [810, 93]. Loadings [211]. Local [111, 846, 27, 819, 814, 184]. Locality [795, 206]. Localization [681]. located [400]. Location [478, 739, 816, 803, 743]. Location-Based [743]. Location-Routing [803]. Lock [236]. Locomotion [462]. Logic [495, 150, 151, 153, 113, 498, 492, 784, 160, 238, 146, 572, 501, 894, 270, 118, 505, 905, 915, 911]. Logic-Based [160, 911]. Logical [158, 347, 567, 570]. Logistics [181]. Long [20, 585, 827]. Long-Lived [20]. Long-Term [585]. Look [69]. Looks [372]. Loop [547]. Loose [247]. LOPSTR [911]. Losing [577]. Loss [565]. Lossless [398]. Love [545]. Low [125]. LR [239]. LSA [741]. LTE [869]. LTL [290]. Lung [837]. LVCSR [51].

MacGuffin [434]. Machine [323, 928, 436]. Machine-Checked [323]. Machines [515]. Machining [586]. Main [721, 838]. Maintaining [849]. Major [67]. Majority [657, 189]. Makes [536]. Making [673, 581]. Málaga [930, 928, 927, 929, 926]. Malignant [65]. Malleability [707]. Malleability-Aware [707]. Malware [291]. Manage [734]. Management [810, 5, 322, 417, 163, 729, 443, 720, 591, 718, 344, 350, 728]. Manifold [57]. Manifold-Based [57]. Manipulating [250]. Manipulation [289]. Many [536]. Map [116, 842, 194, 671]. Mapping [550, 722]. MapReduce [715]. Maps [676, 192, 168, 795]. March [900, 895, 896, 902, 901, 909, 897, 888, 898, 899]. Maritime [734]. Marker [500]. Marker-Directed [500]. Market [883]. Markov [513, 146, 401]. MAS [355]. Mass [431]. Massive [626]. Massively [719]. Matching [419, 105, 628, 691, 60, 49, 388, 719]. Material [404]. Mathematics [890, 891]. Matrices [193, 87, 79]. Matrix [836, 58, 630, 794]. Matter [25, 119, 218, 363, 448, 508, 557, 600, 682, 758, 864, 3, 7, 385, 406, 427, 449, 470, 491, 509, 530, 551, 558, 579, 601, 622, 643, 664, 683, 704, 725, 746, 759, 780, 801, 822, 843, 865, 886, 11, 18, 26, 47, 68, 89, 110, 120, 141, 162, 183, 204, 219, 240, 261, 282, 303, 324, 345, 346, 349, 353, 357, 364]. Maude [497]. Maximal [816]. Maximizes [177]. May [904, 893, 544]. MBC [918]. MCE [51]. McEliece [611]. MD [913]. Mean [170]. Meaning [905]. Meaningful [485]. Means [866, 132, 172, 182]. Measure [676]. Measurement [888]. Measures [856, 789]. Measuring [795, 838]. Mechanism [454, 556, 465]. Mechanisms [8, 715, 678, 344]. MedHelp [532]. Media [533, 550, 541]. Medial [786]. Median [584]. Mediated [464]. Medical [182, 489, 589, 586]. Medication [67]. Meditations [562]. Meet [229]. Meets [264, 527, 751]. Melodies [856]. Membrane [457]. Memetic [56, 177]. Memory [221, 712, 224, 104, 721, 298, 467, 732, 730, 650]. Memory-Aware [650]. Memristive [784]. mente [762]. Merging [519]. Mesh [234]. Meshes [76]. Meshless [397]. Message [797]. Messages [490]. Meta [834, 502]. Meta-dimensional [834]. Meta-predicate [502]. Metabolite [828]. Metabolomic [461]. Metadata [722]. Metaheuristic [804]. Metaheuristics [802]. Metal [59]. MetaMap [599]. Metaphor [670]. Metaverses [582]. Method [519, 840, 590, 714, 209, 837, 58, 629, 829, 423, 388, 595, 52, 769]. Methodologies [887]. Methodology [802, 31]. Methods [138, 912, 807, 834, 51, 825, 124]. Metrics [485, 107, 790]. Metrics-driven [107]. MFF [760]. mHealth [333]. Micro [738, 838]. Micro-blogs [738]. Micro-environment [838]. Middle [222]. Middle-End [222]. Migratability [414]. Migration [871]. Milner [572]. Miner [156]. Mini [114]. Mini-bucket [114]. Minimal [636]. Minimization [265]. Mining [636, 640, 928, 699, 744, 793, 703, 581, 424, 202, 708, 651]. Minority [657, 197]. Mirror [868]. Missing [679]. Mitochondrial [836]. Mixed [765, 393, 443, 820, 275]. Mixed-Strategy [443]. ML [180]. MMB [888]. MMOs [545]. Mobile [666, 72, 667, 432, 431, 328, 668, 720, 329, 109, 332, 327, 52, 665]. Modal [254]. Modalities [373]. Mode [174]. Model [150, 523, 845, 46, 224, 733, 656, 242, 163, 399, 514, 138, 817, 554, 565, 235, 253, 785, 298, 356, 479, 354, 293, 857, 583, 398, 411, 425, 199, 129, 291, 94, 510, 624, 212, 426, 455, 401]. Model-Aware [298]. Model-Based [523, 514]. Model-Driven [242, 425]. Modeling [881, 472, 123, 167, 296, 531, 70, 413, 482, 256, 700, 913, 637]. Modelling [463, 9, 134, 137, 45, 888]. Models [879, 542, 862, 221, 777, 305, 46, 518, 496, 522, 155, 114, 283, 854, 160, 452, 884, 791, 792, 546, 149, 548, 317, 586]. Modern [591]. Modes [437]. Modestly [46]. Modification [374]. Modifications [471]. Modifiers [376]. Modifying [517]. Modular [520, 403, 354, 501]. Molecular [410]. Molecules [454]. Monitor [292]. Monitoring [19, 20, 103, 43, 251, 674, 352, 165, 421]. Monitors [230, 436]. Monodomain [101]. MOPSO [808]. Morphology [761]. Morphosyntactic [763]. Most [655]. Motifs [832]. Motion [128]. Motivation [390]. Motor [135]. Movement [130, 634]. Moving [243]. MPI [98]. MQQ [440]. MQQ-SIG [440]. MRF [592]. mTAN [619]. Mu [254]. Multi [818, 808, 20, 870, 358, 152, 393, 201, 904, 882, 360, 157, 585, 237, 63, 693, 60, 184, 752, 362, 100, 829, 793, 748, 299, 165, 54, 723, 350, 74, 878, 395, 172]. Multi-Agent [904, 360, 362, 165, 350, 358]. Multi-cellular [878]. Multi-class [63]. Multi-cores [100]. Multi-dimensional [585, 723]. Multi-hop [20]. Multi-Label [184, 201]. Multi-layer [54]. Multi-Objective [808, 870, 793, 882]. Multi-Pareto-Ranking [818]. Multi-process [752]. Multi-Processors [74]. Multi-Relational [157, 152]. Multi-source [693]. Multi-stage [237]. Multi-swarm [60, 829]. Multi-symbols [395]. Multi-tasks [172]. Multi-threaded [299]. Multi-touch [393]. Multi-user [748]. Multiagent [875]. Multicollisions [438]. Multifactor [441]. Multimedia [914]. Multimodal [667]. Multiobjective [832, 835]. Multiobjectivizing [817]. Multiparty [229]. Multiple [91, 231, 850, 486, 85, 396, 66]. Multiplication [630]. Multiprocessor [85]. Multiprocessor-Node [85]. Multispecific [459]. Music [851, 855, 929, 863]. Musical [859]. Mutation [826, 56, 794]. Mutations [833]. Mythology [534].

Names [642]. Naples [920]. Narcissists [277]. NASA [912]. Natural [123, 161, 525, 13, 389, 875]. Nearest [200, 739, 170, 195]. NEAT [823]. Need [226, 328]. Needy [332]. Negation [368]. Negative [368, 304]. Neighbor [200, 649, 739, 195]. Neighborhood [803]. Neighbors [696]. Neighbours [170]. Net [792]. Netherlands [905]. Nets [319]. Network [20, 680, 626, 742, 171, 451, 190, 632, 40, 33, 618, 867, 29, 631, 339, 535, 460, 54, 743, 52, 624, 648, 405, 53]. Network-Based [171]. Networked [887, 6, 5, 4, 10, 15, 8, 9, 540, 830]. Networks [19, 745, 111, 72, 73, 179, 734, 549, 543, 27, 234, 456, 443, 618, 452, 146, 187, 147, 293, 866, 744, 868, 713, 467, 539, 23, 340, 655, 841, 344, 336, 450, 831, 877, 338, 769, 318, 594, 623].

Neural [192, 179, 187, 54, 52, 594, 53]. Neuro [881]. Neuro-evolutionary [881]. NeuroEvolution [823]. Neurons [468]. Neutrality [339]. News [475, 476]. Next [726, 338, 708]. Next-Items [708]. NFC [432, 431]. NFC-Enabled [432]. NFM [912]. NFPSLAM [417, 907]. NFPSLAM-SOC [417, 907]. NGN [342]. NGNs [337]. NICs [40]. Nigeria [550]. NN [199, 203]. No [379]. Node [785, 85]. Node-Focused [785]. Nodes [306]. Noise [203, 186, 838, 194]. Noise-Tolerant [203]. Nominal [278]. Non [610, 417, 289, 198, 389]. Non-adaptive [610]. Non-deterministic [289]. Non-Disjoint [198]. Non-Functional [417]. Non-touch [389]. Nonconforming [76]. Nondeterministic [274, 46]. Nonlinear [78, 165, 50]. Norfolk [912]. Norm [765]. Normalized [593]. Norwegian [310]. Notes [351]. Novel [613, 629, 87, 148, 61, 586]. November [908]. Novice [390]. NP [571]. NP-Complete [571]. NSGA [805]. NSGA-II [805]. Number [78]. Numeric [287]. Numerical [56, 80]. Nuprl [567]. Nur [384]. Nursery [597].

O [749, 40]. Object [520, 735, 775, 212, 66]. Object-Oriented [520]. Objective [808, 870, 793, 96, 819, 882]. Objectives [281]. Objects [699, 686]. Oblivious [610, 611]. Observations [597]. Occasion [915]. October [889]. Odense [911]. of-mente [762]. Offshore [734]. Oil [734, 624]. OLAP [717]. Old [394]. Older [394]. On-Board [878]. On-Demand [711]. On-Line [103, 871, 772, 878]. On-the-Fly [16]. One [811, 198, 308, 838]. One-Dependence [198]. Ones [656]. Online [145, 411, 217, 687, 50, 661, 391]. Only [215, 729]. Ontology [779, 130, 638, 12]. Ontology-Based [779, 638, 12]. Open [392, 43, 707, 774, 875]. Open-Ended [875]. Open-Source-Based [43]. OpenCL [97, 205, 101]. OpenGL [101]. Operation [812, 791]. Operation-Based [791]. Operational [251, 507]. operative [885, 60]. Operator [800, 873]. Operators [269, 800]. Opinion [540, 539]. Opportunities [587]. Optic [592]. Optical [868]. Optimal [23, 825, 210, 148]. Optimisation [862]. Optimistic [71]. Optimization [870, 819, 828, 927, 500, 872, 56, 825, 94, 869, 877, 172]. Optimizations [40]. Optimize [882, 170]. Optimizer [808]. Optimizing [866, 825]. Optimum [811]. Option [175]. Orbit [28]. Order [268, 691, 158]. Ordered [281]. Organisations [887]. Organisms [878]. Organization [106, 460, 875]. Organizational [4, 538]. Organizing [192, 168]. Oriented [21, 301, 6, 417, 717, 407, 292, 520, 662, 690, 423, 22, 621, 424, 258, 41, 647, 907]. Orléans [99]. Orthogonal [48]. Orthographic [765]. Oscillations [466]. Out-Degree [546]. Outbreak [598]. Outlier [701, 653]. Outliers [192]. Output [189]. Over-Approximations [294]. Overfitting [799]. Overload [625]. Overloading [571]. Overview [841].

Pacemaker [296]. Pacific [919]. Page [388]. Pairing [617]. Pairing-Based [617]. Panel [589]. Papers [890, 891, 905, 908, 892, 904, 893, 894, 907, 911, 889, 924]. Paphos [907]. Paradigm [685, 826, 207]. Parallel [90, 276, 86, 630, 750, 76, 753, 87, 88, 79, 208, 94, 719, 95, 877, 96, 890, 891]. Parallelism [102]. Parallelization [84, 100, 217]. Parameter [829]. Parameters [745, 145]. Parametric [307]. Parent [391]. Parent-Child [391]. Pareto [818, 819]. Parity [279, 288]. Park [894, 913]. Parsers [239]. Parsing [606, 776]. Part [900, 895, 896, 902, 901, 921, 922, 897, 898, 890, 891, 899]. Partial [552, 196, 293, 256]. Particle [808, 862, 396, 825]. Partition [697]. Partitioned [695]. Partitioning [91]. Partnerships [338]. Parts [59]. Path [739, 272, 52]. Patient [580, 584]. Pattern [523, 628, 640, 464, 639, 914, 708]. Patterns [672, 246, 473, 117, 703]. PDL [528]. Pen [400]. Pen-Based [400]. People [483]. Perception [135]. Percussionist [858]. Perfect [636, 577]. Perfection [280]. Perform [839]. Performance [446, 717, 751, 83, 43, 40, 205, 122, 86, 467, 108, 727, 390, 402, 687]. Performance-Oriented [717]. Permutation [92]. Personal [675, 642]. Personalized [652, 708]. Perspective [265, 870, 588, 460]. Pessimistic [71]. PET [64]. Petri [319]. Phase [809, 627]. Phd [422]. Phenotype [795]. Pheromone [872]. Pheromone-Based [872]. Phones [432, 431, 668]. Phylogenetic [835]. Physical [743]. Physico [460]. Physico-chemical [460]. Physiological [596]. Pi [276]. Pi-Calculus [276]. Pickups [807]. Pictorial [673]. Piecewise [452]. PIES [80]. Pipeline [626]. Piracy [734]. Pivoting [82]. PKCS#11 [322]. Placements [236]. Plagiarised [488]. Planning [142, 585, 809]. Plastic [397]. Plasticity [467]. Plate [76]. Platform [301, 666, 853, 591]. Platformer [885]. Platforms [6]. Platforms-Oriented [6]. Plausibility [556]. Play [390]. Players [391]. Playing [386]. Plink [700]. Plink-LDA [700]. PointBurst [633]. Pointed [267]. Pointing [122]. Points [520, 396, 208]. Points-To [208, 520]. Poland [890, 891]. Policies [320, 871, 430, 327]. Policy [6, 331, 330, 335]. Polish [584]. Polite [378]. Polygonal [80]. Polynomial [518, 590, 154]. Polyphonic [862]. Poor [332]. Pop [848]. Population [876, 849]. Population-Based [876]. Populations [850, 830, 467]. Portable [98]. Portfolio [880, 188]. Portfolios [882]. Portugal [925]. Portuguese [925, 762, 777, 765, 770, 771, 776, 764, 772, 768, 775, 774, 773, 761, 767]. Positioning [668]. Positive [702]. Possible [469]. POST [902]. Posts [637]. Potential [655, 877]. Power [681, 304, 158, 714, 584]. PPAM [890, 891]. Practice [900, 902, 901, 897, 917, 898, 899]. Practitioner [588]. Practitioners [841]. Pragmatic [372, 412, 351, 124]. Pre [602]. Pre-computable [602]. Precise [286]. Precision [484]. Preconditioner [76]. Predicate [517, 502]. Predicate-Level [517]. Predicting [549, 475, 661]. Prediction [545, 542, 836, 214, 817, 585, 420, 187, 180, 489, 595, 52, 50, 913]. Predictive [463, 156]. Predictor [842]. Preface [422]. Preference [380, 857, 710]. Preference-Driven [380]. Preferences [147]. Preprocessing [772].

Preserves [35]. Preserving [36]. Prevalence [544]. Prevent [621]. Prices [661]. Pricing [879]. Principal [169]. Principles [902]. Prior [700]. PRISM [145]. Privacy [309, 304]. Private [338]. Probabilism [366]. Probabilistic [706, 151, 274, 228, 37, 36, 31, 160, 705, 35, 506, 659, 660]. Probability [705]. Problem [804, 90, 126, 807, 805, 442, 815, 76, 809, 874, 816, 803, 181, 143, 370, 806]. Problems [187, 86]. ProbLog [148]. Procedure [722, 173]. Proceedings [895, 896, 921, 922, 910, 900, 925, 902, 930, 920, 901, 928, 912, 927, 916, 909, 929, 926, 897, 917, 888, 898, 919, 918, 913, 923, 899]. Process [414, 179, 9, 234, 424, 194, 241, 752]. Processes [274, 658, 409, 690]. Processign [909]. Processing [706, 103, 925, 685, 454, 718, 739, 674, 13, 773, 890, 891, 379]. Processors [74]. Product [260, 259]. Production [454]. Productivity [751]. Products [726, 735, 260]. Profiles [458, 677, 652]. Program [781, 915, 499, 209, 216, 911, 212]. Programming [879, 150, 358, 98, 781, 153, 782, 826, 783, 799, 784, 785, 747, 856, 347, 786, 750, 859, 827, 753, 207, 926, 894, 788, 247, 789, 791, 793, 794, 351, 898, 621, 796, 757, 258, 831, 924, 904, 898]. Programs [250, 151, 268, 749, 113, 521, 145, 498, 492, 237, 356, 299, 505, 506, 95, 541]. Progress [716]. Project [177]. Projection [154, 48]. Projective [384]. Projectors [84]. Prokaryotes [455]. Prolog [507]. ProMAS [904]. Prompts [231]. Pronouns [775]. Proof [447, 497, 312, 439]. Proofs [323, 567, 299]. Propagation [640, 724, 632]. Properties [225, 606, 453, 27, 417, 505, 460, 35, 32]. Property [73]. Prophetic [270]. PROPHETS [247]. PROPOR [925]. Propositional [118]. Prosody [137]. Protection [678]. Protein [836, 458, 817, 841, 842, 460]. Protein-Protein [841]. Protocol [305, 444, 72, 310, 867, 435]. Protocols [73, 225, 610, 306, 308, 797]. Provable [316]. Provably [311]. Proving [505, 529]. Provisioning [107]. Pruning [636, 185]. Pseudoboolean [811]. Pseudorandom [78]. PSO [60, 829]. Psycho [545]. Psycho-social [545]. PU [646]. Public [337, 338]. PUF [445]. Pure [820]. Purpose [429]. Pushdown [291]. PVS [524]. Python [360].

QoS [418]. QoS-Aware [418]. Quadrotor [174]. Quadruped [462]. Qualitative [355]. Qualities [124]. Quality [552, 420]. Quantificational [373]. Quantified [562]. Quantifiers [564]. Quantitative [280, 513, 283, 355, 384]. Quantized [286]. Quantum [79, 54]. Quasi [877]. Quasi-Parallel [877]. Queries [17, 472, 474, 252, 705, 693, 649, 655, 710, 716, 692]. Query [706, 103, 473, 471, 713, 488, 739, 723, 665]. Querying [680, 15, 262, 696]. Question [778, 369, 481]. Questions [497]. Queue [28]. Quo [726]. QVT [254]. QVT-R [254].

R [254]. Ramsey [288, 271]. Ramsey-Based [288, 271]. Random [799, 178, 487]. Randomly [521]. Range [712, 681, 842]. Range-Free [681]. Ranking [818, 745]. Rapid [454]. Rare [840]. Rating [404]. Rational [347]. RDBMS [591, 693]. RDF [695]. Re [313]. Re-encryption [313]. Reachability [280]. Reaction [824, 575, 468]. Reaction-Diffusion [468]. Reactive [444]. Readings [377]. Real [533, 844, 685, 499, 720, 411, 526, 674, 684, 54]. Real-Coded [54]. Real-Time [844, 685, 720, 411, 674]. Real-World [533]. Reality [666, 395]. Realizing [690]. Reasoning [268, 749, 892, 155, 236, 237, 752, 355]. Receptor [457]. Reclamation [729]. Recognition [642, 176, 914, 586]. Recommendation [171, 662, 481, 708]. Recommendations [677, 633]. Recommending [741]. Reconstruction [57, 592]. Recurrent [821]. Recursion [528]. Recursive [254, 837, 570, 102, 149]. Redactable [603]. Reduced [608]. Reduced-Round [608]. Reducing [39, 82]. Reduction [313, 201, 233, 190, 356, 248, 77, 809, 302]. Reduction-Based [302]. Reference [2, 669]. Referents [371]. Refinement [37, 256]. Reflection [743]. Reflections [578]. Reform [764]. Region [677]. Regional [59]. Regions [95]. Regression [187]. Regressor [170, 50]. Regular [492, 278, 569, 35]. Regularization [480]. Regularized [49]. Regularly [703]. Regulation [328, 329, 326, 327]. Regulatory [453, 451, 452, 450, 831]. Reinforcement [863]. Reject [175]. Related [828, 738, 583, 260]. Relating [381, 348]. Relation [641, 161, 631, 599, 480]. Relational [154, 157, 161, 147, 688, 149, 709, 152]. Relational-Based [688]. Relationship [163, 633]. Relative [552]. Relaxation [49]. Relaxations [499]. Relaxed [273]. Relevance [486, 479, 789]. Reliability [29, 553, 30, 489]. Reliable [175, 73]. Remarks [54]. Remote [436]. Removal [603]. Renewal [441]. Repair [284]. Repairing [311]. Replacement [405]. Replication [71, 208]. Replication-Based [208]. Representation [892, 115, 861]. Representations [113, 383]. Representative [741]. Representing [130, 116]. Requirements [525]. Research [910, 392, 641, 587, 631, 384, 635]. Researchers [841]. Reservation [170]. Residuated [559]. Resilient [72]. Resolution [253]. Resolver [669]. Resource [494, 250, 107, 652]. Resource-Driven [494]. Resources [131, 773, 177]. Respect [869]. Response [812]. Resting [67]. Resting-State [67]. Restricted [36, 442]. Restructuring [206]. Result [483, 692]. Results [885, 903]. Résumé [358]. Retinal [592]. Retrial [28]. Retrieval [910, 485, 479]. Retrieving [488]. Return [621]. Return-Oriented [621]. Reusable [791]. Reverse [706, 649, 831]. Review [737]. Revised [905, 908, 892, 904, 893, 83, 894, 907, 911, 890, 891, 889, 924]. Revisited [320, 226]. Revisiting [274, 317]. Reward [143]. Rewriting [503]. RFID [602, 735]. Ridgelet [50]. Rigorous [518]. Risks [734]. Roaming [328]. Robot [871, 462, 878, 52, 172]. Robotics [563]. Robots [853]. Robust [169, 62]. Robustness [279, 869]. Role [126, 833, 460]. Rollout [342]. Root [620]. Rotation [714]. Rough [163]. Round [608]. Routing [444, 306, 807, 805, 803, 181]. RPCL [51]. RT [515]. Rule [163, 851, 596]. Rules [201, 636, 137, 641, 793]. Running [628, 802]. Runtime [527, 292, 630, 217].

Saaty [79]. SAC [751]. Safe [158, 213]. Safety [523, 408]. Salamanca [895, 896]. Sales [877]. Saliency [128]. Sambamba [217]. Sampling [799]. SAT [284]. Satellite [907]. Satisfaction [562]. Saturators [376]. Saving [714]. SBP [913]. SCADA [626, 624]. Scalable [717, 735, 644, 718, 108, 719, 709, 848]. Scale [378, 630, 70, 535]. Scaling [194]. Scenarios [565]. Scene [847]. Scenes [128]. Schedule [248]. Schedulers [36]. Scheduling [870, 90, 91, 812, 34, 791, 94, 95, 177, 96]. Schema [691]. Scheme [447, 616, 602, 445, 442, 739, 711, 728, 402]. Schemes [446, 114]. School [597, 924]. Science [900, 916, 889]. Scientific [69]. Scope [371, 383]. Screen [614]. Screening [591, 825]. SDDSfL [104]. Search [483, 846, 698, 667, 474, 695, 819, 473, 657, 854, 403, 740, 814, 688, 638, 482, 813, 327, 803, 185, 650, 652, 709, 647, 665, 28, 806]. Searching [712, 765, 60]. Seasonal [879]. Second [892, 889]. Secrecy [797]. Sector [341, 343]. Secure [444, 619, 605, 620, 604]. Security [301, 192, 320, 527, 305, 429, 312, 902, 556, 308, 917, 914, 424, 645, 615, 914]. SecurityPolicy [419]. Segmentation [139, 663, 182, 132, 592, 635, 594, 593]. Segmentations [132]. Segmenting [650]. Selected [905, 908, 892, 904, 893, 903, 894, 907, 911, 890, 891, 889, 924]. Selection [880, 836, 200, 882, 837, 553, 884, 873, 210, 202, 418, 64]. Selective [319]. Self [192, 168, 800, 504, 875]. Self-adaptation [800]. Self-organization [875]. Self-Organizing [192, 168]. Self-sustainability [504]. Semantic [676, 779, 419, 641, 638, 722, 669, 790, 117, 769]. Semantics [367, 915, 276, 140, 502, 382, 507, 348, 723]. Semi [139, 646, 699, 55]. Semi-automatic [139]. Semi-supervised [646, 699, 55]. SenDe [918]. Sense [774]. Sensing [57, 387]. Sensitive [65, 387]. Sensitivity [466]. Sensor [19, 20, 680, 679, 632, 443, 618, 866, 23, 623]. Sensors [681]. Sentence [377, 647]. Sentence-Internal [377]. Separation [307, 238]. September [890, 891]. Sequence [156, 390]. Sequences [568, 575, 650, 651]. Sequential [50, 708]. Series [884, 702, 829, 50, 661]. Service [21, 301, 6, 417, 246, 407, 245, 167, 415, 735, 420, 412, 425, 690, 423, 340, 333, 418, 41, 426, 433, 421, 405, 907]. Service-Oriented [21, 301, 417, 407, 690, 423, 907]. Services [680, 416, 429, 105, 328, 408, 639, 739, 418, 481]. Session [72, 229]. Set [153, 113, 163, 654, 116]. Set-Based [163]. Set-Labeled [116]. Sets [834, 570]. Settings [610]. Seventh [407]. Shared [195]. Sharing [590, 634, 790]. Shift [531]. Shop [90, 791]. Short [842]. Short-Range [842]. SIG [440]. Sign [136, 134, 137, 893, 138, 135, 139, 374]. Signage [393]. Signal [465, 61]. Signalling [457, 456, 459]. Signals [168]. Signature [602]. Signatures [603]. Signed [537, 131]. Significant [744]. Silico [834]. SIM [923]. Similarity [676, 779, 657, 638, 538]. Simple [112, 203, 270]. Simplex [83]. Simplifying [497]. SIMT [235]. Simulate [455]. Simulated [853, 462].

Simulating [468]. Simulation [46, 179, 451, 469, 99, 78, 133, 397]. Simulative [38]. SINBAD [696]. Single [751]. Sino [889]. Sino-foreign-interchange [889]. Site [595]. Sites [91]. Situational [538]. Size [210]. Skeleton [99]. Skewed [687]. Skyline [706, 707, 655]. SLA [423, 32, 41]. SLA-Aware [423]. SLCA [697]. Sliding [174]. Slow [544]. Small [870, 566, 78]. Small-Step [566]. Small-World [870]. Smart [810, 736]. SMEs [6]. Smoking [544, 541]. Smoothing [484]. SMT [529, 510]. SMT-Based [510]. SNSM [923]. SOC [417, 907]. Social [533, 542, 745, 680, 672, 393, 742, 171, 550, 331, 478, 740, 744, 631, 713, 539, 332, 655, 334, 548, 743, 877, 633, 541, 545, 913]. Societies [540, 164]. Society [327, 743]. Socio [540]. Socio-cultural [540]. Socioemotional [582]. Software [511, 900, 902, 901, 408, 343, 524, 30, 897, 38, 898, 259, 755, 428, 215, 899]. Software-Only [215]. Solar [179]. Solution [762, 58]. Solutions [8, 114]. Solve [807]. Solver [83, 529]. Solving [126, 806]. SOM [196]. Somatosensory [394]. Some [526, 757, 433]. Something [382]. Sound [852, 856, 127, 859, 929, 435]. Soundness [314, 606]. Source [392, 43, 773, 28, 693]. Sources [16, 838]. South [921, 922, 923]. Space [361, 233, 729, 307, 84, 42, 356, 868]. Spaces [233]. Spain [910, 895, 896, 892, 930, 928, 927, 929, 926]. Spares [28]. Sparse [316, 87, 50]. Spatial [126, 632, 721, 161, 677, 129, 133, 189, 665]. Spatio [457, 679]. Spatio-temporal [457, 679]. Spatiotemporal [464]. Speaker [169]. Special [914]. Specialization [499, 875]. Specific [493]. Specification [227, 528]. Specifications [243]. Spectral [127]. Spectrum [763]. Speech [667, 485, 167, 718, 623]. Speed [665]. Speeding [199]. Spelling [764]. SPIDAR [400]. SPIDAR-Pen [400]. Splice [595]. Splitting [661]. Spontaneous [126, 176]. Spread [598]. SQL [689]. Square [620]. Squares [58]. SQWRL [680]. SR [40]. SR-IOV [40]. SRAM [445]. SSA [222]. SSA-Based [222]. SSD [726]. Stable [246]. Staccato [132]. Stack [40]. Stacker [804]. Stage [662, 237]. Standard [311]. Standardization [722]. Star [693]. Stars [693, 655]. State [519, 233, 67, 356, 575, 515]. State-Space [233]. Static [521, 211, 755]. Statistical [46, 376, 517, 538]. Statistics [723]. Status [67, 726]. STE3D [613]. STE3D-CAP [613]. Steel [187]. Steepest [880]. Steered [563]. Steganalysis [556]. Stem [463]. Stems [778]. Step [566]. Stepmothers [277]. Steps [526]. Steroids [693]. Sticky [430]. Stigmergic [872]. Stochastic [829, 108, 28]. Stock [883]. Stokes [80]. Storage [733, 91, 717, 729, 731, 714, 70, 583, 728]. Strategic [409]. Strategies [394, 461, 93, 580, 84, 373, 186, 317]. Strategy [4, 359, 503, 443, 820, 625]. Stratified [670]. Stream [706, 720, 634, 718]. Streaming [701, 109, 674, 402]. Streams [628, 719]. Strictly [77]. String [627]. Strip [594]. Strongly [781]. Structural [745, 654, 460]. Structurally [279]. Structure [151, 378, 276, 817, 603, 696, 284, 660]. Structure-Independent [696]. Structured [485, 776, 479]. Structures [836, 900, 892]. Studies [133].

Study [2, 309, 532, 667, 882, 729, 469, 722, 51, 186, 529, 391, 541]. Sub [467]. Sub-populations [467]. Subgame [280]. Subgraph [699]. Subgroup [152]. Subsequence [650]. Subsystem [30]. Subtyping [564]. Sufficient [306]. Suggestion [647]. Suite [517]. Summarization [737]. Summer [924]. supervised [646, 699, 55]. Supply [24]. Support [17, 532, 474, 214, 724, 106, 40, 251, 424, 402, 650, 767, 215]. Supported [95]. Supporting [309, 115]. sustainability [504]. Sustaining [821]. SVM [837, 654]. SVMs [184]. Swarm [808, 862, 825, 60, 829]. SWIP [117]. Symbolic [314, 305, 493, 259, 435, 515]. symbols [395]. Symmetric [82]. Symposium [920, 912, 898, 422, 911]. Symptom [599]. Synaptic [467]. SynchAADL2Maude [244]. Synchronisation [138]. Synchronization [72, 253]. Synchronous [639]. Syntactic [777, 790]. Synthesis [137, 852, 911]. System [626, 171, 360, 851, 456, 812, 285, 720, 404, 598, 187, 70, 465, 570, 718, 117, 165, 217, 869, 350, 624, 586]. Systematic [231]. Systems [39, 82, 10, 8, 908, 458, 895, 896, 904, 920, 724, 901, 115, 839, 587, 249, 354, 293, 634, 921, 922, 639, 824, 752, 77, 362, 669, 381, 829, 173, 423, 756, 874, 575, 888, 898, 755, 875, 74, 450, 41, 923].

Tables [584, 689]. TACAS [901]. Tag [671, 651, 671]. Tags [741, 477]. Tailored [461]. Taipei [904]. Taiwan [904]. Takedown [317]. TAL [525]. Tallinn [900, 902, 901, 897, 898, 899]. Taming [408]. Target [458, 57, 829]. Targeting [874]. Tariffs [340, 332]. Task [249, 823, 395]. Tasks [122, 180, 450, 172]. Taxonomy [166, 662]. Taxonomy-Oriented [662]. Taylor [518]. TDSC [627]. Teaching [392]. Team [359]. Teams [539, 582, 172]. Technical [883]. Technique [233, 799, 436]. Techniques [242, 109, 665]. Technological [623, 4]. Technologies [887, 331, 919, 918]. Technology [580, 731]. Tel [903]. Telecommunication [325, 903]. Telecommunications [343, 330]. Telehealth [580]. Television [335]. Tell [382]. Temperature [187]. Templates [97, 472]. Temporal [496, 670, 770, 658, 525, 270, 457, 679]. Tent [194]. Term [585, 768, 827, 480]. Terminal [29]. Termination [300, 504, 299, 506, 507]. Terms [570]. Test [494, 232, 517, 423]. Testing [274, 521, 871, 30, 259]. Text [175]. Texts [770, 771]. TFETI [84]. Their [317]. Them [498]. Theoretic [318]. Theoretical [133]. Theories [313, 268, 381]. Theory [537, 900, 227, 902, 901, 135, 567, 534, 897, 898, 899]. Think [129]. Thinking [685, 304]. Third [908]. Thomann [132]. Thousands [529]. thread [206]. threaded [299]. Threading [98]. Threat [535]. Three [398]. Three-Dimensional [398]. Throughput [591]. Thumb [134]. Thyroid [593]. Ticketing [432, 431]. Tile [210]. Time [578, 844, 73, 685, 771, 802, 154, 31, 807, 720, 86, 884, 411, 702, 829, 674, 684, 50, 661]. Time-Series [50]. Timed [37, 513]. Times [180]. Timing [31, 127]. Tiny [681]. Tissues [909]. TNC [435]. Together [133]. Tolerance [888]. Tolerant [203]. Tone [862]. Too [286]. Tool [244, 764, 876]. Toolchain [511]. Tools [901]. Top [705, 709, 697]. Top- [705, 709]. Top-Down [697]. Topic [700, 637].

Topicality [371]. Topics [369]. Topological [569]. Topologies [306, 823]. Toponym [779]. Torun [890, 891]. Touch [614, 393, 389]. Tourism [487]. TPM [430]. Trace [313, 274, 233, 673]. Traceability [24]. Tracing [735]. Tracking [675, 399, 396, 830, 66]. Trademarks [330]. Trading [883]. Traditional [124]. Traffic [44, 674, 869]. Trained [54]. Training [51, 53]. Traits [391]. Trajectories [672]. Transactional [472, 703]. Transactions [223, 906, 914]. Transcript [454]. Transcription [862, 455]. Transfer [610, 565, 611, 188]. Transfers [74]. Transform [580]. Transformation [629, 396, 911]. Transformations [254, 500, 584, 505]. Transforming [497, 255]. Transit [431]. Transition [293]. Translation [290, 41]. Transmission [681, 809]. Transport [341]. Travel [673]. Traveling [858]. Treatments [458]. Tree [492, 289, 184, 569, 732, 842, 660]. Tree-Based [842]. Trees [497, 781, 158, 146, 187, 186, 835]. Trend [679]. Trends [542]. Tri [53]. Tri-training [53]. Triangular [193]. Tribe [640]. Tribe-Leaders [640]. Tricks [394]. Trilateration [668]. Triple [253]. Truncated [609, 170]. Trust [545, 301, 902, 443, 645, 633]. Trust-Relationship [633]. Trusted [429, 908]. TrustZone [428]. Truth [741, 113]. Truth-Dependencies [113]. TSO [297, 224]. Tumor [593]. Tuning [873]. Tutor [747]. Tweets [477]. Twig [628, 693]. Twitter [598]. Twitterverse [550]. Two [138, 616, 29, 546, 803, 627, 96]. Two-Echelon [803]. Two-phase [627]. Two-Terminal [29]. Type [322, 127]. Type-Based [322]. Types [229, 564, 467, 197]. Typical [694].

U.S. [580]. UI [425]. UK [909, 894]. Ultrasonic [668]. Ultrasound [65]. UML [262, 515]. UML-RT [515]. Unbounded [308]. Unboxing [213]. UnCAL [500]. Uncertain [706, 653]. Uncertainty [573]. Unconditional [314]. Under-Approximation [516]. Under-Approximations [294]. Underspecified [383]. Understanding [597, 827]. Unfold [505]. Unfold/Fold [505]. Unlabeled [646, 702]. Unlabeled-Data-Distribution [646]. Unrated [656]. Unreliability [27]. Unsafe [211]. Unsupervised [168]. Unsymmetric [87]. Unusual [454]. Update [724, 381]. Updating [694]. UPPAAL [295]. Upper [193]. Urban [743]. URLs [477]. USA [912, 913]. Usage [723]. Use [104, 550]. Used [190, 196, 855]. User [672, 764, 123, 167, 854, 340, 389, 652, 124, 748]. User-Adapted [167]. Uses [548]. Using [810, 542, 779, 447, 518, 152, 742, 295, 499, 492, 139, 469, 585, 834, 720, 57, 837, 122, 293, 634, 413, 583, 411, 538, 488, 789, 654, 792, 727, 109, 504, 117, 270, 180, 809, 186, 256, 835, 118, 185, 259, 617, 389, 621, 529, 350, 450, 194, 64, 700, 169, 661, 284, 405, 80, 555, 191, 848]. Utilizing [184].

VA [912]. Validating [239]. Validation [301, 548]. Valuable [839]. Value [214, 583]. Value-Based [214]. Variability [827]. Variable [803]. Variants [840]. Variational [160]. Variations [665]. Vasopressin [465]. VATA [289]. VCR [402]. Vector [848]. Vehicle [807, 805, 634, 181]. Verb [760, 761, 767]. Verbal [140, 673]. Verification [313, 527, 10, 305, 447, 225, 522, 512, 286, 307, 602, 285, 524, 99, 296, 526, 173, 308, 529]. Verified [250, 511, 222, 315]. Verifying [249, 33]. Version [347]. versus [192]. Vertical [174]. Veterans [394, 580]. via [472, 254, 201, 67, 36, 867, 708]. Video [394, 173, 109]. Viewports [676]. Views [694]. VIII [914, 906]. Vineyard [20]. Vineyards [19]. Virtual [106, 436, 825, 334, 395, 582]. Virtualization [104, 40, 430, 621]. Visible [398]. Vision [404]. Visual [670, 614, 616, 638, 400]. Visualization [670, 109, 648]. Visually [671, 118]. Visuo [133]. Visuo-Spatial [133]. VMAD [216]. Voting [189, 310]. Vowel [531]. vs [376, 665].

W2GIS [920]. Walks [487]. Water [812]. Watermark [398, 555]. Ways [378]. Weak [221, 775]. Weather [879]. Weather-Derivatives [879]. Weaving [742]. Web [919, 416, 483, 675, 671, 920, 474, 93, 478, 638, 631, 639, 425, 482, 748, 117, 919, 645, 918, 647]. Web-Based [671]. Webpages [489]. Weight [49]. Weighted [35, 189]. Weighting [200, 480]. Weights [825]. Well [267]. Well-Pointed [267]. WESOA [407, 907]. Where [369, 129, 254]. Which [4]. Whole [833]. Wide [371]. Wiimote [122]. Wikipedia [766, 657, 769, 637, 648]. WiMAX [39]. WiMAX-Based [39]. Windows [807]. Windsor [894]. Wine [4, 24]. Winning [577]. Wireless [19, 920, 234, 632, 443, 618, 866, 23]. within [5, 580, 581, 450]. without [606, 194]. Word [663]. WordNet [738]. Words [403]. Work [839]. Workbench [44]. Workers [486]. Workflow [9, 105, 658, 410]. Workflows [91]. Workshop [892, 417, 904, 893, 407, 889]. Workshops [907, 918, 923]. World [533, 870, 550]. Worm [632]. Worst [31, 33]. Worst-Case [31]. Wrapper [202]. Write [727]. Written [99]. WS [419]. WS-SecurityPolicy [419]. WSNs [21, 22].

Xi’an [889]. XML [698, 628, 694, 100, 711, 696, 659, 651, 692]. XOR [618]. XQuery [495].

Years [69].

Zapping [39]. Zooplankton [827].

References

Anastasi:2012:I

[1] Giuseppe Anastasi, Emilio Bellini, Elisabetta Di Nitto, Carlo Ghezzi, and Letizia Tanca. Introduction. Lecture Notes in Computer Science, 7200:1–5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

     (electronic). URL http://link.springer.com/chapter/10.1007/978-
     3-642-31739-2_1/.

Anastasi:2012:RCS

[2] Giuseppe Anastasi, Emilio Bellini, Elisabetta Di Nitto, Carlo Ghezzi, and Letizia Tanca. Reference case study. Lecture Notes in Computer Science, 7200:7–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31739-2_2/.

Anonymous:2012:FMa

[3] Anonymous. Front matter. Lecture Notes in Computer Science, 7200: 15–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/bfm:978-
3-642-31739-2/2/1.

Bellini:2012:WCS

[4] Emilio Bellini. Which collaboration strategy for the networked enterprise in wine industry? Technological and organizational challenges. Lecture Notes in Computer Science, 7200:17–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31739-2_3/.

Bellini:2012:DDA

[5] Emilio Bellini, Claudio Dell’ Era, and Roberto Verganti. A designdriven approach for the innovation management within networked enterprises. Lecture Notes in Computer Science, 7200:31–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31739-2_4/.

Basalisco:2012:ENE

[6] Bruno Basalisco and Guido M. Rey. Enhancing the networked enterprise for SMEs: a service platforms-oriented industrial policy. Lecture Notes in Computer Science, 7200:59–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31739-2_5/.

Anonymous:2012:FMb

[7] Anonymous. Front matter. Lecture Notes in Computer Science, 7200: 83–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/bfm:978-
3-642-31739-2/3/1.

Calcavecchia:2012:CAS

[8] Nicolò M. Calcavecchia, Elisabetta Di Nitto, Daniel J. Dubois, and Carlo Ghezzi. Complex autonomic systems for networked enterprises: Mechanisms, solutions and design approaches. Lecture Notes in Computer Science, 7200:85–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31739-2_6/.

Canfora:2012:AWB

[9] Gerardo Canfora, Giancarlo Tretola, and Eugenio Zimeo. Autonomic workflow and business process modelling for networked enterprises. Lecture Notes in Computer Science, 7200:115–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31739-2_7/.

Bertolino:2012:VAA

[10] Antonia Bertolino, Guglielmo De Angelis, Felicita Di Giandomenico, and Eda Marchetti. Verification and analysis of autonomic systems for networked enterprises. Lecture Notes in Computer Science, 7200: 143–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31739-2_8/.

Anonymous:2012:FMc

[11] Anonymous. Front matter. Lecture Notes in Computer Science, 7200: 171–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/bfm:978-
3-642-31739-2/4/1.

Montedoro:2012:OBK

[12] Marcello Montedoro, Giorgio Orsi, Licia Sbattella, and Roberto Tedesco. Ontology-based knowledge elicitation: An architecture. Lecture Notes in Computer Science, 7200:175–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31739-2_9/.

Sbattella:2012:KEN

[13] Licia Sbattella and Roberto Tedesco. Knowledge extraction from natural language processing. Lecture Notes in Computer Science, 7200: 193–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31739-2_10/.

Mahdiraji:2012:KEE

[14] Alireza Rezaei Mahdiraji, Bruno Rossi, Alberto Sillitti, and Giancarlo Succi. Knowledge extraction from events flows. Lecture Notes in Computer Science, 7200:221–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31739-2_11/.

Bolchini:2012:CAK

[15] Cristiana Bolchini, Elisa Quintarelli, Fabio A. Schreiber, and Maria Teresa Baldassarre. Context-aware knowledge querying in a networked enterprise. Lecture Notes in Computer Science, 7200:237–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31739-2_12/.

Orsi:2012:FCA

[16] Giorgio Orsi and Letizia Tanca. On-the-fly and context-aware integration of heterogeneous data sources. Lecture Notes in Computer Science, 7200: 259–276, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31739-2_13/.

Bolchini:2012:CSD

[17] Cristiana Bolchini, Elisa Quintarelli, and Letizia Tanca. Context support for designing analytical queries. Lecture Notes in Computer Science, 7200: 277–289, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31739-2_14/.

Anonymous:2012:FMd

[18] Anonymous. Front matter. Lecture Notes in Computer Science, 7200: 291–293, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/bfm:978-
3-642-31739-2/5/1.

Alippi:2012:WSN

[19] Cesare Alippi, Giacomo Boracchi, Romolo Camplani, and Manuel Roveri. Wireless sensor networks for monitoring vineyards. Lecture Notes in Computer Science, 7200:295–310, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31739-2_15/.

Anastasi:2012:DIF

[20] Giuseppe Anastasi, Marco Conti, Mario Di Francesco, and Ilaria Giannetti. Design, implementation, and field experimentation of a longlived multi-hop sensor network for vineyard monitoring. Lecture Notes in Computer Science, 7200:311–327, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31739-2_16/.

Anastasi:2012:EDW

[21] Gaetano F. Anastasi, Enrico Bini, and Giuseppe Lipari. Extracting data from WSNs: a service-oriented approach. Lecture Notes in Computer Science, 7200:329–356, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31739-2_17/.

Schreiber:2012:EDW

[22] Fabio A. Schreiber, Romolo Camplani, and Guido Rota. Extracting data from WSNs: a data-oriented approach. Lecture Notes in Computer Science, 7200:357–373, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31739-2_18/.

Mura:2012:ODW

[23] Marcello Mura, Simone Campanoni, William Fornaciari, and Mariagiovanna Sami. Optimal design of wireless sensor networks. Lecture Notes in Computer Science, 7200:375–395, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31739-2_19/.

Cimino:2012:ETW

[24] Mario G. C. A. Cimino and Francesco Marcelloni. Enabling traceability in the wine supply chain. Lecture Notes in Computer Science, 7200: 397–412, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31739-2_20/.

Anonymous:2012:BMa

[25] Anonymous. Back matter. Lecture Notes in Computer Science, 7200: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31739-2/1.

Anonymous:2012:FMe

[26] Anonymous. Front matter. Lecture Notes in Computer Science, 7200: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31739-2/1.

Daduna:2012:ALN

[27] Hans Daduna and Lars Peter Saul. Availability in large networks: Global characteristics from local unreliability properties. Lecture Notes in Computer Science, 7201:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28540-0_1/.

Zhang:2012:SAF

[28] Feng Zhang and Jinting Wang. Stochastic analysis of a finite source retrial queue with spares and orbit search. Lecture Notes in Computer Science, 7201:16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28540-0_2/.

Le:2012:BTT

[29] Minh Lê and Max Walter. Bounds for two-terminal network reliability with dependent basic events. Lecture Notes in Computer Science, 7201:31–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28540-0_3/.

Meitner:2012:SRT

[30] Matthias Meitner and Francesca Saglietti. Software reliability testing covering subsystem interactions. Lecture Notes in Computer Science, 7201: 46–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28540-0_4/.

Hofig:2012:FDT

[31] Kai Höfig. Failure-dependent timing analysis — a new methodology for probabilistic worst-case execution time analysis. Lecture Notes in Computer Science, 7201:61–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28540-0_5/.

Vastag:2012:CSD

[32] Sebastian Vastag. A calculus for SLA delay properties. Lecture Notes in Computer Science, 7201:76–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28540-0_6/.

Ivkovic:2012:VWC

[33] Nikola Ivkovic, Dario Kresic, Kai-Steffen Hielscher, and Reinhard German. Verifying worst case delays in controller area network. Lecture Notes in Computer Science, 7201:91–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28540-0_7/.

Jongerden:2012:LIB

[34] Marijn R. Jongerden and Boudewijn R. Haverkort. Lifetime improvement by battery scheduling. Lecture Notes in Computer Science, 7201: 106–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28540-0_8/.

Sharma:2012:WPE

[35] Arpit Sharma. Weighted probabilistic equivalence preserves ω-regular properties. Lecture Notes in Computer Science, 7201:121–135, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-28540-0_9/

.

Georgievska:2012:PCP

[36] Sonja Georgievska and Suzana Andova. Probabilistic CSP: Preserving the laws via restricted schedulers. Lecture Notes in Computer Science, 7201: 136–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28540-0_10/.

Fioriti:2012:HPT

[37] Luis María Ferrer Fioriti and Holger Hermanns. Heuristics for probabilistic timed automata with abstraction refinement. Lecture Notes in Computer Science, 7201:151–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28540-0_11/.

Sadre:2012:SAE

[38] Ramin Sadre, Anne Remke, Sjors Hettinga, and Boudewijn Haverkort. Simulative and analytical evaluation for ASD-based embedded software. Lecture Notes in Computer Science, 7201:166–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28540-0_12/.

Abdollahpouri:2012:RCZ

[39] Alireza Abdollahpouri and Bernd E. Wolfinger. Reducing channel zapping delay in WiMAX-based IPTV systems. Lecture Notes in Computer Science, 7201:182–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28540-0_13/.

Huang:2012:PEN

[40] Shu Huang and Ilia Baldine. Performance evaluation of 10GE NICs with SR-IOV support: I/O virtualization and network stack optimizations. Lecture Notes in Computer Science, 7201:197–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28540-0_14/.

Winkler:2012:BDB

[41] Ulrich Winkler, Wasif Gilani, and Alan Marshall. Business driven BCM SLA translation for service oriented systems. Lecture Notes in Computer Science, 7201:206–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28540-0_15/.

Jahr:2012:BDS

[42] Ralf Jahr, Horia Calborean, Lucian Vintan, and Theo Ungerer. Boosting design space explorations with existing or automatically learned knowledge. Lecture Notes in Computer Science, 7201:221–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28540-0_16/.

Hoefling:2012:IOS

[43] Michael Hoefling, Michael Menth, Christian Kniep, and Marcus Camen. IBPM: An open-source-based framework for InfiniBand performance monitoring. Lecture Notes in Computer Science, 7201:236–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28540-0_17/.

Eittenberger:2012:WIT

[44] Philipp M. Eittenberger and Udo R. Krieger. A workbench for Internet traffic analysis. Lecture Notes in Computer Science, 7201:240–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28540-0_18/.

Gouberman:2012:MAE

[45] Alexander Gouberman, Martin Riedl, Johann Schuster, and Markus Siegle. A modelling and analysis environment for LARES. Lecture Notes in Computer Science, 7201:244–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28540-0_19/.

Bogdoll:2012:SSM

[46] Jonathan Bogdoll, Arnd Hartmanns, and Holger Hermanns. Simulation and statistical model checking for modestly nondeterministic models. Lecture Notes in Computer Science, 7201:249–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28540-0_20/.

Anonymous:2012:FMf

[47] Anonymous. Front matter. Lecture Notes in Computer Science, 7201: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28540-0/1.

Lin:2012:OPA

[48] Binbin Lin, Chiyuan Zhang, and Xiaofei He. Orthogonal projection analysis. Lecture Notes in Computer Science, 7202:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31919-8_1/.

Liu:2012:WRR

[49] Zhi-Yong Liu, Hong Qiao, and Lei Xu. A weight regularized relaxation based graph matching algorithm. Lecture Notes in Computer Science, 7202:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31919-8_2/.

Yang:2012:OSE

[50] Shuyuan Yang, DiJun Zuo, Min Wang, and Licheng Jiao. Online sequential extreme learning of sparse ridgelet kernel regressor for nonlinear time-series prediction. Lecture Notes in Computer Science, 7202: 17–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31919-8_3/.

Pang:2012:CSR

[51] Zaihu Pang, Xihong Wu, and Lei Xu. A comparative study of RPCL and MCE based discriminative training methods for LVCSR. Lecture Notes in Computer Science, 7202:27–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31919-8_4/.

Wu:2012:PPM

[52] Zhiqian Wu, Masafumi Hashimoto, Baolong Guo, and Kazuhiko Takahashi. A path prediction method for human-accompanying mobile robot based on neural network. Lecture Notes in Computer Science, 7202: 35–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31919-8_5/.

Zhang:2012:TTB

[53] Xiaojie Zhang, Bendu Bai, and Ying Li. Tri-training based on neural network ensemble algorithm. Lecture Notes in Computer Science, 7202: 43–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31919-8_6/.

Takahashi:2012:RML

[54] Kazuhiko Takahashi, Motoki Kurokawa, and Masafumi Hashimoto. Remarks on multi-layer quantum neural network controller trained by realcoded genetic algorithm. Lecture Notes in Computer Science, 7202: 50–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31919-8_7/.

Li:2012:EDG

[55] Chun guang Li, Xianbiao Qi, Jun Guo, and Bo Xiao. An evaluation on different graphs for semi-supervised learning. Lecture Notes in Computer Science, 7202:58–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31919-8_8/.

Li:2012:MAD

[56] Yangyang Li, Bo Wu, Lc Jiao, and Ruochen Liu. Memetic algorithm with double mutation for numerical optimization. Lecture Notes in Computer Science, 7202:66–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31919-8_9/.

Hou:2012:TRU

[57] Biao Hou, Xi Cheng, and Hua Qiong Jiang. Target reconstruction using manifold-based compressive sensing. Lecture Notes in Computer Science, 7202:74–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31919-8_10/.

Li:2012:IML

[58] Lin Li, Xiu jiu Yuan, and Hong Liu. An iterative method for the least squares anti-bisymmetric solution of the matrix equation AX = B. Lecture Notes in Computer Science, 7202:81–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31919-8_11/.

Wang:2012:RDB

[59] Wei Wang, Wei ping He, Lei Lei, and Wen tao Li. The regional detection of 2D barcode in complicated backgrounds of metal parts. Lecture Notes in Computer Science, 7202:89–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31919-8_12/.

Liu:2012:SBM

[60] Luoxia Liu and Bendu Bai. Searching for the best matching atoms based on multi-swarm co-operative PSO. Lecture Notes in Computer Science, 7202:97–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31919-8_13/.

Zhang:2012:NAC

[61] Dongwei Zhang and Ying Guo. A novel algorithm for CPM-signal endpoint detection based on auto correlation. Lecture Notes in Computer Science, 7202:104–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31919-8_14/.

Zhang:2012:GKD

[62] Zhen Zhang, Yanning Zhang, Rui Yao, Haisen Li, and Yu Zhu. Generalized kernel density estimation based robust estimator and its application. Lecture Notes in Computer Science, 7202:112–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31919-8_15/.

Kim:2012:MCC

[63] Tae-Hyun Kim, Dong-Chul Park, Dong-Min Woo, Taikyeong Jeong, and Soo-Young Min. Multi-class classifier-based AdaBoost algorithm. Lecture Notes in Computer Science, 7202:122–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31919-8_16/.

Xia:2012:GAB

[64] Yong Xia, Zhe Zhang, Lingfeng Wen, Pei Dong, and David Dagan Feng. GA and AdaBoost-based feature selection and combination for automated identification of dementia using FDG-PET imaging. Lecture Notes in Computer Science, 7202:128–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31919-8_17/.

Shen:2012:DBB

[65] Xing Shen, Shuheng Zhang, Rui Yao, Yaqing Chen, Yue-Min Zhu, and Su Zhang. Discrimination between benign and malignant breast cancers in ultrasound images based on cost-sensitive boosting. Lecture Notes in Computer Science, 7202:136–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31919-8_18/.

Zhou:2012:MFF

[66] Yu Zhou, Cong Rao, Qin Lu, Xiang Bai, and Wenyu Liu. Multiple feature fusion for object tracking. Lecture Notes in Computer Science, 7202: 145–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31919-8_19/.

Gao:2012:EMS

[67] Wenying Gao, Ling-Li Zeng, Hui Shen, and Dewen Hu. Estimating medication status via resting-state functional connectivity in major depression. Lecture Notes in Computer Science, 7202:153–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31919-8_20/.

Anonymous:2012:FMg

[68] Anonymous. Front matter. Lecture Notes in Computer Science, 7202: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31919-8/1.

Wasniewski:2012:LBY

[69] Jerzy Waśniewski. A look back: 57 years of scientific computing. Lecture Notes in Computer Science, 7203:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31464-3_1/.

Liu:2012:MLS

[70] Ning Liu, Christopher Carothers, Jason Cope, Philip Carns, Robert Ross, and Adam Crume. Modeling a leadership-scale storage system. Lecture Notes in Computer Science, 7203:10–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31464-3_2/.

Bazydlo:2012:COP

[71] Marcin Bazydło, Szymon Francuzik, Cezary Sobaniec, and Dariusz Wawrzyniak. Combining optimistic and pessimistic replication. Lecture Notes in Computer Science, 7203:20–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31464-3_3/.

Brzezinski:2012:RSG

[72] Jerzy Brzeziński, Dariusz Dwornikowski, Łukasz Pia̧tkowski, and Grzegorz Sobański. K-resilient session guarantees synchronization protocol for mobile ad-hoc networks. Lecture Notes in Computer Science, 7203: 30–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31464-3_4/.

Brzezinski:2012:TCR

[73] Jerzy Brzeziński, Michał Kalewski, and Dariusz Wawrzyniak. On time constraints of reliable broadcast protocols for Ad Hoc networks with the liveness property. Lecture Notes in Computer Science, 7203:40– 49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31464-3_5/.

Tudruj:2012:DTF

[74] Marek Tudruj and Łukasz Maśko. Data transfers on the fly for hierarchical systems of chip multi-processors. Lecture Notes in Computer Science, 7203:50–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31464-3_6/.

Gustavson:2012:NLB

[75] Fred G. Gustavson, Jerzy Waśniewski, and José R. Herrero. New level-3 BLAS kernels for Cholesky factorization. Lecture Notes in Computer Science, 7203:60–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31464-3_7/.

Marcinkowski:2012:PPN

[76] Leszek Marcinkowski. Parallel preconditioner for nonconforming adini discretization of a plate problem on nonconforming meshes. Lecture Notes in Computer Science, 7203:70–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31464-3_8/.

Mikkelsen:2012:ICR

[77] Carl Christian Kjelgaard Mikkelsen and Bo Kågström. Incomplete cyclic reduction of banded and strictly diagonally dominant linear systems. Lecture Notes in Computer Science, 7203:80–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31464-3_9/.

Neves:2012:FSN

[78] Samuel Neves and Filipe Araujo. Fast and small nonlinear pseudorandom number generators for computer simulation. Lecture Notes in Computer Science, 7203:92–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31464-3_10/.

Piech:2012:PQA

[79] Henryk Piech and Olga Siedlecka-Lamch. Parallel quantum algorithm for finding the consistency of Saaty’s matrices. Lecture Notes in Computer Science, 7203:102–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31464-3_11/.

Zieniuk:2012:NAD

[80] Eugeniusz Zieniuk, Krzysztof Szerszen, and Marta Kapturczak. A numerical approach to the determination of 3D Stokes flow in polygonal domains using PIES. Lecture Notes in Computer Science, 7203: 112–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31464-3_12/.

Gustavson:2012:CBL

[81] Fred G. Gustavson. Cache blocking for linear algebra algorithms. Lecture Notes in Computer Science, 7203:122–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31464-3_13/.

Becker:2012:RAP

[82] Dulceneia Becker, Marc Baboulin, and Jack Dongarra. Reducing the amount of pivoting in symmetric indefinite systems. Lecture Notes in Computer Science, 7203:133–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31464-3_14/.

Hall:2012:HPD

[83] Julian Hall and Qi Huangfu. A high performance dual revised simplex solver. Lecture Notes in Computer Science, 7203:143–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31464-3_15/.

Hapla:2012:TCS

[84] Vaclav Hapla and David Horak. TFETI coarse space projectors parallelization strategies. Lecture Notes in Computer Science, 7203:152– 162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31464-3_16/.

Jocksch:2012:FMC

[85] Andreas Jocksch. FFTs and multiple collective communication on multiprocessor-node architectures. Lecture Notes in Computer Science, 7203:163–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31464-3_17/.

Lirkov:2012:PAP

[86] Ivan Lirkov, Marcin Paprzycki, and Maria Ganzha. Performance analysis of parallel alternating directions algorithm for time dependent problems. Lecture Notes in Computer Science, 7203:173–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31464-3_18/.

Murri:2012:NPA

[87] Riccardo Murri. A novel parallel algorithm for Gaussian elimination of sparse unsymmetric matrices. Lecture Notes in Computer Science, 7203:

183–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31464-3_19/.

Olas:2012:PFA

[88] Tomasz Olas, Roman Wyrzykowski, and Pawel Gepner. Parallel FEM adaptation on hierarchical architectures. Lecture Notes in Computer Science, 7203:194–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31464-3_20/.

Anonymous:2012:FMh

[89] Anonymous. Front matter. Lecture Notes in Computer Science, 7203: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31464-3/1.

Bozejko:2012:PCF

[90] Wojciech Bożejko, Mariusz Uchroński, and Mieczysław Wodecki. Parallel cost function determination on GPU for the job shop scheduling problem. Lecture Notes in Computer Science, 7204:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31500-8_1/.

Chen:2012:PSW

[91] Weiwei Chen and Ewa Deelman. Partitioning and scheduling workflows across multiple sites with storage constraints. Lecture Notes in Computer Science, 7204:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31500-8_2/.

Drozdowski:2012:GBB

[92] Maciej Drozdowski, Paweł Marciniak, Grzegorz Pawlak, and Maciej Płaza. Grid branch-and-bound for permutation flowshop. Lecture Notes in Computer Science, 7204:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31500-8_3/.

Gehweiler:2012:ECL

[93] Joachim Gehweiler, Peter Kling, and Friedhelm Meyer auf der Heide. An experimental comparison of load balancing strategies in a Web computing environment. Lecture Notes in Computer Science, 7204:31–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31500-8_4/.

Switalski:2012:GSB

[94] Piotr Switalski and Franciszek Seredynski. A grid scheduling based on generalized extremal optimization for parallel job model. Lecture Notes in Computer Science, 7204:41–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31500-8_5/.

Tudruj:2012:SPP

[95] Marek Tudruj and Łukasz Maśko. Scheduling parallel programs based on architecture–supported regions. Lecture Notes in Computer Science, 7204:51–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31500-8_6/.

Yaurima-Basaldua:2012:GAC

[96] Victor Hugo Yaurima-Basaldua, Andrei Tchernykh, and Yair Castro-Garcia. Genetic algorithm calibration for two objective scheduling parallel jobs on hierarchical grids. Lecture Notes in Computer Science, 7204: 61–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31500-8_7/.

Bawidamann:2012:ETO

[97] Uwe Bawidamann and Marco Nehmeier. Expression templates and OpenCL. Lecture Notes in Computer Science, 7204:71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31500-8_8/.

Berka:2012:PET

[98] Tobias Berka, Helge Hagenauer, and Marian Vajteršic. Portable explicit threading and concurrent programming for MPI applications. Lecture Notes in Computer Science, 7204:81–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31500-8_9/.

Javed:2012:VHD

[99] Noman Javed and Frédéric Loulergue. Verification of a heat diffusion simulation written with Orléans skeleton library. Lecture Notes in Computer Science, 7204:91–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31500-8_10/.

Muldner:2012:PXD

[100] Tomasz Müldner, Christopher Fry, Tyler Corbin, and Jan Krzysztof Miziołek. Parallelization of an XML data compressor on multi-cores. Lecture Notes in Computer Science, 7204:101–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31500-8_11/.

Oliveira:2012:CCO

[101] Rafael Sachetto Oliveira, Bernardo Martins Rocha, and Ronan Mendonça Amorim. Comparing CUDA, OpenCL and OpenGL implementations of the cardiac monodomain equations. Lecture Notes in Computer Science, 7204:111–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31500-8_12/.

Saougkos:2012:FGP

[102] Dimitris Saougkos, Aristeidis Mastoras, and George Manis. Fine grained parallelism in recursive function calls. Lecture Notes in Computer Science, 7204:121–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31500-8_13/.

Balis:2012:LGM

[103] Bartosz Balis, Grzegorz Dyk, and Marian Bubak. On-line grid monitoring based on distributed query processing. Lecture Notes in Computer Science, 7204:131–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31500-8_14/.

Chrobot:2012:DMV

[104] Arkadiusz Chrobot, Maciej Lasota, Grzegorz Łukawski, and Krzysztof Sapiecha. Distributed memory virtualization with the use of SDDSfL. Lecture Notes in Computer Science, 7204:141–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31500-8_15/.

Czarnul:2012:DCM

[105] Paweł Czarnul and Michał Wójcik. Dynamic compatibility matching of services for distributed workflow execution. Lecture Notes in Computer Science, 7204:151–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31500-8_16/.

Haupt:2012:CSE

[106] Tomasz Haupt, Nitin Sukhija, and Mark F. Horstemeyer. Cyberinfrastructure support for engineering virtual organization for CyberDesign. Lecture Notes in Computer Science, 7204:161–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31500-8_17/.

Koperek:2012:DBM

[107] Paweł Koperek and Włodzimierz Funika. Dynamic business metrics-driven resource provisioning in cloud environments. Lecture Notes in Computer Science, 7204:171–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31500-8_18/.

Onderka:2012:SCS

[108] Zdzislaw Onderka. Stochastic control of the scalable high performance distributed computations. Lecture Notes in Computer Science, 7204: 181–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31500-8_19/.

Panka:2012:DCV

[109] Maciej Panka, Michal Chlebiej, Krzysztof Benedyczak, and Piotr Bała. Distributed collaborative visualization on mobile devices using interactive video streaming techniques. Lecture Notes in Computer Science, 7204: 191–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31500-8_20/.

Anonymous:2012:FMi

[110] Anonymous. Front matter. Lecture Notes in Computer Science, 7204: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31500-8/1.

Bareinboim:2012:LCC

[111] Elias Bareinboim, Carlos Brito, and Judea Pearl. Local characterizations of causal Bayesian networks. Lecture Notes in Computer Science, 7205:1–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29449-5_1/.

Carloni:2012:BFS

[112] Olivier Carloni. Boolean formulas of simple conceptual graphs ( SGBF ). Lecture Notes in Computer Science, 7205:18–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-29449-5_2/.

Costantini:2012:CCT

[113] Stefania Costantini and Alessandro Provetti. Conflict, consistency and truth-dependencies in graph representations of answer set logic programs. Lecture Notes in Computer Science, 7205:68–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29449-5_3/.

Flerova:2012:BMB

[114] Natalia Flerova, Emma Rollon, and Rina Dechter. Bucket and mini-bucket schemes for M best solutions over graphical models. Lecture Notes in Computer Science, 7205:91–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29449-5_4/.

Fortin:2012:SAS

[115] Jérôme Fortin and Rallou Thomopoulos. Supporting argumentation systems by graph representation and computation. Lecture Notes in Computer Science, 7205:119–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29449-5_5/.

Niveau:2012:RCS

[116] Alexandre Niveau and Hélène Fargier. Representing CSPs with set-labeled diagrams: a compilation map. Lecture Notes in Computer Science, 7205: 137–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29449-5_6/.

Pradel:2012:SWI

[117] Camille Pradel and Ollivier Haemmerlé. A semantic Web interface using patterns: The SWIP system. Lecture Notes in Computer Science, 7205: 172–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29449-5_7/.

Schlegel:2012:VIK

[118] Daniel R. Schlegel and Stuart C. Shapiro. Visually interacting with a knowledge base using frames, logic, and propositional graphs. Lecture Notes in Computer Science, 7205:188–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29449-5_8/.

Anonymous:2012:BMb

[119] Anonymous. Back matter. Lecture Notes in Computer Science, 7205: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29449-5/1.

Anonymous:2012:FMj

[120] Anonymous. Front matter. Lecture Notes in Computer Science, 7205: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29449-5/1.

Anastasiou:2012:GAL

[121] Dimitra Anastasiou. Gestures in assisted living environments. Lecture Notes in Computer Science, 7206:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34182-3_1/.

Kouroupetroglou:2012:UWP

[122] Georgios Kouroupetroglou, Alexandros Pino, and Athanasios Balmpakakis. Using Wiimote for 2D and 3D pointing tasks: Gesture performance evaluation. Lecture Notes in Computer Science, 7206:13–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34182-3_2/.

Glomb:2012:CMH

[123] Przemysław Głomb, Michał Romaszewski, Sebastian Opozda, and Arkadiusz Sochan. Choosing and modeling the hand gesture database for a natural user interface. Lecture Notes in Computer Science, 7206: 24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34182-3_3/.

vanBeurden:2012:UEG

[124] Maurice H. P. H. van Beurden, Wijnand A. Ijsselsteijn, and Yvonne A. W. de Kort. User experience of gesture based interfaces: a comparison with traditional interaction methods on pragmatic and hedonic qualities. Lecture Notes in Computer Science, 7206:36–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34182-3_4/.

Kontogeorgakopoulos:2012:LCF

[125] Alexandros Kontogeorgakopoulos and Georgios Kouroupetroglou. Low cost force-feedback interaction with haptic digital audio effects. Lecture Notes in Computer Science, 7206:48–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34182-3_5/.

Chu:2012:RSG

[126] Mingyuan Chu and Sotaro Kita. The role of spontaneous gestures in spatial problem solving. Lecture Notes in Computer Science, 7206: 57–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34182-3_6/.

Kozak:2012:ESF

[127] Mariusz Kozak, Kristian Nymoen, and Rolf Inge Godøy. Effects of spectral features of sound on gesture type and timing. Lecture Notes in Computer Science, 7206:69–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34182-3_7/.

Pirri:2012:HMS

[128] Fiora Pirri, Matia Pizzoli, and Matei Mancas. Human-motion saliency in complex scenes. Lecture Notes in Computer Science, 7206:81–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34182-3_8/.

Panayi:2012:WWW

[129] Marilyn Panayi. What, why, where and how do children think? Towards a dynamic model of spatial cognition as action. Lecture Notes in Computer Science, 7206:93–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34182-3_9/.

ElRaheb:2012:LBO

[130] Katerina El Raheb and Yannis Ioannidis. A Labanotation based ontology for representing dance movement. Lecture Notes in Computer Science, 7206:106–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34182-3_10/.

Crasborn:2012:IDC

[131] Onno Crasborn and Menzo Windhouwer. ISOcat data categories for signed language resources. Lecture Notes in Computer Science, 7206: 118–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34182-3_11/.

Lucking:2012:AAS

[132] Andy Lücking, Sebastian Ptock, and Kirsten Bergmann. Assessing agreement on segmentations by means of Staccato, the segmentation agreement calculator according to Thomann. Lecture Notes in Computer Science, 7206:129–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34182-3_12/.

Rieser:2012:HDI

[133] Hannes Rieser, Kirsten Bergmann, and Stefan Kopp. How do iconic gestures convey visuo-spatial information? Bringing together empirical, theoretical, and simulation studies. Lecture Notes in Computer Science, 7206: 139–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34182-3_13/.

Delorme:2012:TMG

[134] Maxime Delorme, Michael Filhol, and Annelies Braffort. Thumb modelling for the generation of sign language. Lecture Notes in Computer Science, 7206:151–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34182-3_14/.

Gibet:2012:TMT

[135] Sylvie Gibet, Pierre-François Marteau, and Kyle Duarte. Toward a motor theory of sign language perception. Lecture Notes in Computer Science, 7206:161–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34182-3_15/.

Braffort:2012:ADB

[136] Annelies Braffort and Emilie Chételat-Pelé. Analysis and description of blinking in French sign language for automatic generation. Lecture Notes in Computer Science, 7206:173–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34182-3_16/.

Dimou:2012:GPM

[137] Athanasia-Lida Dimou, Theodore Goulas, and Eleni Efthimiou. Grammar/prosody modelling in Greek sign language: Towards the definition of built-in sign synthesis rules. Lecture Notes in Computer Science, 7206: 183–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34182-3_17/.

Filhol:2012:CTS

[138] Michael Filhol. Combining two synchronisation methods in a linguistic model to describe sign language. Lecture Notes in Computer Science, 7206:194–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34182-3_18/.

Gonzalez:2012:SSU

[139] Matilde Gonzalez and Christophe Collet. Sign segmentation using dynamics and hand configuration for semi-automatic annotation of sign language corpora. Lecture Notes in Computer Science, 7206:204–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34182-3_19/.

Giorgolo:2012:IGV

[140] Gianluca Giorgolo. Integration of gesture and verbal language: a formal semantics approach. Lecture Notes in Computer Science, 7206: 216–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34182-3_20/.

Anonymous:2012:FMk

[141] Anonymous. Front matter. Lecture Notes in Computer Science, 7206: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34182-3/1.

Geffner:2012:ILP

[142] Hector Geffner. Inference and learning in planning. Lecture Notes in Computer Science, 7207:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-31951-8_1.

Sutton:2012:BRP

[143] Richard S. Sutton. Beyond reward: The problem of knowledge and data. Lecture Notes in Computer Science, 7207:2–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-31951-8_2/.

Walsh:2012:EC

[144] Toby Walsh. Exploiting constraints. Lecture Notes in Computer Science, 7207:7–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31951-8_3/.

Cussens:2012:OBI

[145] James Cussens. Online Bayesian inference for the parameters of PRISM programs. Lecture Notes in Computer Science, 7207:14–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31951-8_4/.

Khosravi:2012:LCM

[146] Hassan Khosravi, Oliver Schulte, Jianfeng Hu, and Tianxiang Gao. Learning compact Markov logic networks with decision trees. Lecture Notes in Computer Science, 7207:20–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31951-8_5/.

Koriche:2012:RNC

[147] Frédéric Koriche. Relational networks of conditional preferences. Lecture Notes in Computer Science, 7207:26–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31951-8_6/.

Renkens:2012:ONA

[148] Joris Renkens, Guy Van den Broeck, and Siegfried Nijssen. k-Optimal: a novel approximate inference algorithm for ProbLog. Lecture Notes in Computer Science, 7207:33–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31951-8_7/.

Schulte:2012:LDR

[149] Oliver Schulte, Hassan Khosravi, and Tong Man. Learning directed relational models with recursive dependencies. Lecture Notes in Computer Science, 7207:39–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31951-8_8/.

Alrajeh:2012:IMC

[150] Dalal Alrajeh, Alessandra Russo, Sebastian Uchitel, and Jeff Kramer. Integrating model checking and inductive logic programming. Lecture Notes in Computer Science, 7207:45–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31951-8_9/.

Bellodi:2012:LSP

[151] Elena Bellodi and Fabrizio Riguzzi. Learning the structure of probabilistic logic programs. Lecture Notes in Computer Science, 7207: 61–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31951-8_10/.

Cernoch:2012:SDU

[152] Radomír Černoch and Filip Železný. Subgroup discovery using Bump hunting on multi-relational histograms. Lecture Notes in Computer Science, 7207:76–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31951-8_11/.

Corapi:2012:ILP

[153] Domenico Corapi, Alessandra Russo, and Emil Lupu. Inductive logic programming in answer set programming. Lecture Notes in Computer Science, 7207:91–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31951-8_12/.

Douar:2012:GBR

[154] Brahim Douar, Michel Liquiere, Chiraz Latiri, and Yahya Slimani. Graph-based relational learning with a polynomial time projection algorithm. Lecture Notes in Computer Science, 7207:98–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31951-8_13/.

Dubba:2012:IIA

[155] Krishna Dubba, Mehul Bhatt, Frank Dylla, David C. Hogg, and Anthony G. Cohn. Interleaved inductive-abductive reasoning for learning complex event models. Lecture Notes in Computer Science, 7207: 113–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31951-8_14/.

Ferreira:2012:PSM

[156] Carlos Abreu Ferreira, João Gama, and Vítor Santos Costa. Predictive sequence Miner in ILP learning. Lecture Notes in Computer Science, 7207: 130–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31951-8_15/.

Fonseca:2012:CCM

[157] Nuno A. Fonseca, Vítor Santos Costa, and Rui Camacho. Conceptual clustering of multi-relational data. Lecture Notes in Computer Science, 7207:145–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31951-8_16/.

Gillis:2012:EPS

[158] Joris J. M. Gillis and Jan Van den Bussche. Expressive power of safe firstorder logical decision trees. Lecture Notes in Computer Science, 7207: 160–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31951-8_17/.

Inoue:2012:DHE

[159] Katsumi Inoue. DNF hypotheses in explanatory induction. Lecture Notes in Computer Science, 7207:173–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31951-8_18/.

Ishihata:2012:VBI

[160] Masakazu Ishihata, Yoshitaka Kameya, and Taisuke Sato. Variational Bayes inference for logic-based probabilistic models on BDDs. Lecture Notes in Computer Science, 7207:189–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31951-8_19/.

Kordjamshidi:2012:RLS

[161] Parisa Kordjamshidi, Paolo Frasconi, Martijn Van Otterlo, and Marie-Francine Moens. Relational learning for spatial relation extraction from natural language. Lecture Notes in Computer Science, 7207:204–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31951-8_20/.

Anonymous:2012:FMl

[162] Anonymous. Front matter. Lecture Notes in Computer Science, 7207: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31951-8/1.

Fan:2012:AMI

[163] Yu-Neng Fan and Ching-Chin Chern. An agent model for incremental rough set-based rule induction in customer relationship management. Lecture Notes in Computer Science, 7208:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28942-2_1/.

Jordan:2012:CBA

[164] Jaume Jordán, Stella Heras, and Vicente Julián. Case-based argumentation infrastructure for agent societies. Lecture Notes in Computer Science, 7208:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28942-2_2/.

Skupin:2012:AMA

[165] Piotr Skupin and Mieczyslaw Metzger. The application of multi-agent system in monitoring and control of nonlinear bioprocesses. Lecture Notes in Computer Science, 7208:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28942-2_3/.

Aguero:2012:ACT

[166] Jorge Agüero, Miguel Rebollo, Carlos Carrascosa, and Vicente Julián. Agent capability taxonomy for dynamic environments. Lecture Notes in Computer Science, 7208:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28942-2_4/.

Griol:2012:MIU

[167] David Griol, Javier Carbó, and José Manuel Molina. Modeling Internet as a user-adapted speech service. Lecture Notes in Computer Science, 7208:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28942-2_5/.

Cruz:2012:UCA

[168] Ricardo Cruz, Andrés Ortiz, Ana M. Barbancho, and Isabel Barbancho. Unsupervised classification of audio signals by self-organizing maps and Bayesian labeling. Lecture Notes in Computer Science, 7208:61– 70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28942-2_6/.

Yang:2012:RSI

[169] IL-Ho Yang, Min-Seok Kim, Byung-Min So, Myung-Jae Kim, and Ha-Jin Yu. Robust speaker identification using ensembles of kernel principal component analysis. Lecture Notes in Computer Science, 7208: 71–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28942-2_7/.

Sanz-Garcia:2012:AGA

[170] Andrés Sanz-García, Julio Fernández-Ceniceros, and Fernando Antoñanzas-Torres. Application of genetic algorithms to optimize a truncated mean k-nearest neighbours regressor for hotel reservation forecasting. Lecture Notes in Computer Science, 7208:79–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28942-2_8/.

Davoodi:2012:SNB

[171] Elnaz Davoodi, Mohsen Afsharchi, and Keivan Kianmehr. A social network-based approach to expert recommendation system. Lecture Notes in Computer Science, 7208:91–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28942-2_9/.

deLope:2012:DMT

[172] Javier de Lope, Darío Maravall, and Yadira Quiñonez. Decentralized multi-tasks distribution in heterogeneous robot teams by means of ant colony optimization and learning automata. Lecture Notes in Computer Science, 7208:103–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28942-2_10/.

Owczarek:2012:LPL

[173] Agnieszka Owczarek and Krzysztof Ślot. Lipreading procedure for liveness verification in video authentication systems. Lecture Notes in Computer Science, 7208:115–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28942-2_11/.

Zeghlache:2012:FSM

[174] S. Zeghlache, D. Saigaa, K. Kara, Abdelghani Harrag, and A. Bouguerra. Fuzzy sliding mode control with chattering elimination for a quadrotor helicopter in vertical flight. Lecture Notes in Computer Science, 7208: 125–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28942-2_12/.

Armano:2012:EBL

[175] Giuliano Armano, Camelia Chira, and Nima Hatami. Ensemble of binary learners for reliable text categorization with a reject option. Lecture Notes in Computer Science, 7208:137–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28942-2_13/.

Piatkowska:2012:SFE

[176] Ewa Pia̧tkowska and Jerzy Martyna. Spontaneous facial expression recognition: Automatic aggression detection. Lecture Notes in Computer Science, 7208:147–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28942-2_14/.

Yannibelli:2012:MAP

[177] Virginia Yannibelli and Analía Amandi. A memetic approach to project scheduling that maximizes the effectiveness of the human resources assigned to project activities. Lecture Notes in Computer Science, 7208: 159–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28942-2_15/.

Konijn:2012:HFR

[178] R. M. Konijn and W. Kowalczyk. Hunting for fraudsters in random forests. Lecture Notes in Computer Science, 7208:174–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28942-2_16/.

Bukhtoyarov:2012:NNE

[179] Vladimir Bukhtoyarov, Eugene Semenkin, and Andrey Shabalov. Neural networks ensembles approach for simulation of solar arrays degradation process. Lecture Notes in Computer Science, 7208:186–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28942-2_17/.

Priya:2012:UGA

[180] Rattan Priya, Bruno Feres de Souza, André L. D. Rossi, and André C. P. L. F. de Carvalho. Using genetic algorithms to improve prediction of execution times of ML tasks. Lecture Notes in Computer Science, 7208: 196–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28942-2_18/.

Simic:2012:HAI

[181] Dragan Simić and Svetlana Simić. Hybrid artificial intelligence approaches on vehicle routing problem in logistics distribution. Lecture Notes in Computer Science, 7208:208–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28942-2_19/.

Liu:2012:FCM

[182] Yuchen Liu, Kai Xiao, Alei Liang, and Haibing Guan. Fuzzy C-means clustering with bilateral filtering for medical image segmentation. Lecture Notes in Computer Science, 7208:221–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28942-2_20/.

Anonymous:2012:FMm

[183] Anonymous. Front matter. Lecture Notes in Computer Science, 7208: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28942-2/1.

Madjarov:2012:HDT

[184] Gjorgji Madjarov and Dejan Gjorgjevikj. Hybrid decision tree architecture utilizing local SVMs for multi-label classification. Lecture Notes in Computer Science, 7209:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28931-6_1/.

Sheen:2012:EPU

[185] Shina Sheen, S. V. Aishwarya, R. Anitha, S. V. Raghavan, and S. M. Bhaskar. Ensemble pruning using harmony search. Lecture Notes in Computer Science, 7209:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28931-6_2/.

Saez:2012:FSD

[186] José A. Sáez, Mikel Galar, Julián Luengo, and Francisco Herrera. A first study on decomposition strategies with data with class noise using decision trees. Lecture Notes in Computer Science, 7209:25–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28931-6_3/.

Kordos:2012:CAN

[187] Miroslaw Kordos, Piotr Kania, Pawel Budzyna, Marcin Blachnik, and Tadeusz Wieczorek. Combining the advantages of neural networks and decision trees for regression problems in a steel temperature prediction system. Lecture Notes in Computer Science, 7209:36–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28931-6_4/.

Kajdanowicz:2012:TLA

[188] Tomasz Kajdanowicz, Slawomir Plamowski, Przemyslaw Kazienko, and Wojciech Indyk. Transfer learning approach to debt portfolio appraisal. Lecture Notes in Computer Science, 7209:46–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28931-6_5/.

Toman:2012:GWM

[189] Henrietta Toman, Laszlo Kovacs, Agnes Jonas, Lajos Hajdu, and Andras Hajdu. Generalized weighted majority voting with an application to algorithms having spatial output. Lecture Notes in Computer Science, 7209:56–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28931-6_6/.

Grzenda:2012:TRD

[190] Maciej Grzenda. Towards the reduction of data used for the classification of network flows. Lecture Notes in Computer Science, 7209: 68–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28931-6_7/.

delRey:2012:EDI

[191] A. Martín del Rey, G. Rodríguez Sánchez, and A. de la Villa Cuenca. Encrypting digital images using cellular automata. Lecture Notes in Computer Science, 7209:78–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28931-6_8/.

Bankovic:2012:SOM

[192] Zorana Banković, David Fraga, Juan Carlos Vallejo, and José M. Moya. Self-organizing maps versus growing neural gas in detecting data outliers for security applications. Lecture Notes in Computer Science, 7209: 89–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28931-6_9/.

Alvarez:2012:CAB

[193] Rafael Álvarez, Francisco Martínez, José-Francisco Vicent, and Antonio Zamora. Cryptographic applications of 3 × 3 block upper triangular matrices. Lecture Notes in Computer Science, 7209:97–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28931-6_10/.

Vazquez-Medina:2012:DCN

[194] Ruben Vazquez-Medina, José Luis Del-Río-Correa, and César Enrique Rojas-López. Digital chaotic noise using tent map without scaling and discretization process. Lecture Notes in Computer Science, 7209:105– 115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28931-6_11/.

Tomasev:2012:HAS

[195] Nenad Tomašev and Dunja Mladenić. Hubness-aware shared neighbor distances for high-dimensional k-nearest neighbor classification. Lecture Notes in Computer Science, 7209:116–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28931-6_12/.

Jaramillo-Vacio:2012:CCL

[196] Rubén Jaramillo-Vacio, Alberto Ochoa-Zezzatti, and Armando Rios-Lira. Comparison of competitive learning for SOM used in classification of partial discharge. Lecture Notes in Computer Science, 7209: 128–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28931-6_13/.

Napierala:2012:IDT

[197] Krystyna Napierala and Jerzy Stefanowski. Identification of different types of minority class examples in imbalanced data. Lecture Notes in Computer Science, 7209:139–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28931-6_14/.

Martinez:2012:NDD

[198] Ana M. Martínez, Geoffrey I. Webb, M. Julia Flores, and José A. Gámez. Non-disjoint discretization for aggregating one-dependence estimator classifiers. Lecture Notes in Computer Science, 7209:151–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28931-6_15/.

Ougiaroglou:2012:AHC

[199] Stefanos Ougiaroglou, Georgios Evangelidis, and Dimitris A. Dervos. An adaptive hybrid and cluster-based model for speeding up the k-NN classifier. Lecture Notes in Computer Science, 7209:163–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28931-6_16/.

Derrac:2012:CEF

[200] Joaquín Derrac, Isaac Triguero, Salvador García, and Francisco Herrera. A co-evolutionary framework for nearest neighbor enhancement: Combining instance and feature weighting with instance selection. Lecture Notes in Computer Science, 7209:176–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28931-6_17/.

Charte:2012:IML

[201] Francisco Charte, Antonio Rivera, María José del Jesus, and Francisco Herrera. Improving multi-label classifiers via label reduction with association rules. Lecture Notes in Computer Science, 7209:188–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28931-6_18/.

Unold:2012:GBW

[202] Olgierd Unold, Maciej Dobrowolski, Henryk Maciejewski, Pawel Skrobanek, and Ewa Walkowicz. A GA-based wrapper feature selection for animal breeding data mining. Lecture Notes in Computer Science, 7209: 200–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28931-6_19/.

Ougiaroglou:2012:SNT

[203] Stefanos Ougiaroglou and Georgios Evangelidis. A simple noise-tolerant abstraction algorithm for fast k-NN classification. Lecture Notes in Computer Science, 7209:210–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28931-6_20/.

Anonymous:2012:FMn

[204] Anonymous. Front matter. Lecture Notes in Computer Science, 7209: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28931-6/1.

Karrenberg:2012:IPO

[205] Ralf Karrenberg and Sebastian Hack. Improving performance of OpenCL on CPUs. Lecture Notes in Computer Science, 7210:1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28652-0_1/.

Unkule:2012:ARG

[206] Swapneela Unkule, Christopher Shaltz, and Apan Qasem. Automatic restructuring of GPU kernels for exploiting inter-thread data locality. Lecture Notes in Computer Science, 7210:21–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28652-0_2/.

Marron:2012:PPD

[207] Mark Marron, Ondřej Lhoták, and Anindya Banerjee. Programming paradigm driven heap analysis. Lecture Notes in Computer Science, 7210: 41–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28652-0_3/.

Putta:2012:PRB

[208] Sandeep Putta and Rupesh Nasre. Parallel replication-based points-to analysis. Lecture Notes in Computer Science, 7210:61–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28652-0_4/.

Hou:2012:NMP

[209] Cong Hou, George Vulov, Daniel Quinlan, David Jefferson, and Richard Fujimoto. A new method for program inversion. Lecture Notes in Computer Science, 7210:81–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28652-0_5/.

Shirako:2012:ABO

[210] Jun Shirako, Kamal Sharma, Naznin Fauzia, Louis-Noël Pouchet, and J. Ramanujam. Analytical bounds for optimal tile size selection. Lecture Notes in Computer Science, 7210:101–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28652-0_6/.

Kwon:2012:SDU

[211] Taeho Kwon and Zhendong Su. Static detection of unsafe component loadings. Lecture Notes in Computer Science, 7210:122–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28652-0_7/.

Yang:2012:OMC

[212] Jing Yang, Gogul Balakrishnan, Naoto Maeda, Franjo Ivančić, Aarti Gupta, and Nishant Sinha. Object model construction for inheritance in C++ and its applications to program analysis. Lecture Notes in Computer Science, 7210:144–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28652-0_8/.

Petersen:2012:GSI

[213] Leaf Petersen and Neal Glew. GC-safe interprocedural unboxing. Lecture Notes in Computer Science, 7210:165–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28652-0_9/.

Farooq:2012:CSV

[214] Muhammad Umar Farooq, Lei Chen, and Lizy Kurian John. Compiler support for value-based indirect branch prediction. Lecture Notes in Computer Science, 7210:185–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28652-0_10/.

Zhao:2012:CSF

[215] Chuck (Chengyan) Zhao, J. Gregory Steffan, Cristiana Amza, and Allan Kielstra. Compiler support for fine-grain software-only checkpointing. Lecture Notes in Computer Science, 7210:200–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28652-0_11/.

Jimborean:2012:VAD

[216] Alexandra Jimborean, Luis Mastrangelo, Vincent Loechner, and Philippe Clauss. VMAD: An advanced dynamic program analysis and instrumentation framework. Lecture Notes in Computer Science, 7210:220– 239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28652-0_12/.

Streit:2012:SRS

[217] Kevin Streit, Clemens Hammacher, Andreas Zeller, and Sebastian Hack. Sambamba: a runtime system for online adaptive parallelization. Lecture Notes in Computer Science, 7210:240–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28652-0_13/.

Anonymous:2012:BMc

[218] Anonymous. Back matter. Lecture Notes in Computer Science, 7210: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28652-0/1.

Anonymous:2012:FMo

[219] Anonymous. Front matter. Lecture Notes in Computer Science, 7210: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28652-0/1.

Stroustrup:2012:FC

[220] Bjarne Stroustrup. Foundations of C++. Lecture Notes in Computer Science, 7211:1–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28869-2_1/.

Atig:2012:WDA

[221] Mohamed Faouzi Atig and Ahmed Bouajjani. What’s decidable about weak memory models? Lecture Notes in Computer Science, 7211: 26–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28869-2_2/.

Barthe:2012:FVS

[222] Gilles Barthe, Delphine Demange, and David Pichardie. A formally verified SSA-based middle-end. Lecture Notes in Computer Science, 7211: 47–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28869-2_3/.

Burckhardt:2012:ECT

[223] Sebastian Burckhardt, Daan Leijen, and Manuel Fähndrich. Eventually consistent transactions. Lecture Notes in Computer Science, 7211: 67–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28869-2_4/.

Burckhardt:2012:CLC

[224] Sebastian Burckhardt and Alexey Gotsman. Concurrent library correctness on the TSO memory model. Lecture Notes in Computer Science, 7211:87–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28869-2_5/.

Chadha:2012:AVE

[225] Rohit Chadha, Ştefan Ciobâcă, and Steve Kremer. Automated verification of equivalence properties of cryptographic protocols. Lecture Notes in Computer Science, 7211:108–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28869-2_6/.

Chang:2012:CNL

[226] Stephen Chang and Matthias Felleisen. The call-by-need lambda calculus, revisited. Lecture Notes in Computer Science, 7211:128–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28869-2_7/.

Chen:2012:CST

[227] Taolue Chen, Chris Chilton, and Bengt Jonsson. A compositional specification theory for component behaviours. Lecture Notes in Computer Science, 7211:148–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28869-2_8/.

Cousot:2012:PAI

[228] Patrick Cousot and Michael Monerau. Probabilistic abstract interpretation. Lecture Notes in Computer Science, 7211:169–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28869-2_9/.

Denielou:2012:MST

[229] Pierre-Malo Deniélou and Nobuko Yoshida. Multiparty session types meet communicating automata. Lecture Notes in Computer Science, 7211: 194–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28869-2_10/.

Dimoulas:2012:CMB

[230] Christos Dimoulas and Sam Tobin-Hochstadt. Complete monitors for behavioral contracts. Lecture Notes in Computer Science, 7211:214– 233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28869-2_11/.

Downen:2012:SAD

[231] Paul Downen and Zena M. Ariola. A systematic approach to delimited control with multiple prompts. Lecture Notes in Computer Science, 7211: 234–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28869-2_12/.

Emoto:2012:GTA

[232] Kento Emoto, Sebastian Fischer, and Zhenjiang Hu. Generate, test, and aggregate. Lecture Notes in Computer Science, 7211:254–273, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28869-2_13/.

Fajstrup:2012:TSE

[233] Lisbeth Fajstrup, Éric Goubault, and Emmanuel Haucourt. Trace spaces: An efficient new technique for state-space reduction. Lecture Notes in Computer Science, 7211:274–294, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28869-2_14/.

Fehnker:2012:PAW

[234] Ansgar Fehnker, Rob van Glabbeek, and Peter Höfner. A process algebra for wireless mesh networks. Lecture Notes in Computer Science, 7211: 295–315, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28869-2_15/.

Habermaier:2012:CSE

[235] Axel Habermaier and Alexander Knapp. On the correctness of the SIMT execution model of GPUs. Lecture Notes in Computer Science, 7211: 316–335, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28869-2_16/.

Hawkins:2012:RAL

[236] Peter Hawkins, Alex Aiken, and Kathleen Fisher. Reasoning about lock placements. Lecture Notes in Computer Science, 7211:336–356, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28869-2_17/.

Inoue:2012:RAM

[237] Jun Inoue and Walid Taha. Reasoning about multi-stage programs. Lecture Notes in Computer Science, 7211:357–376, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28869-2_18/.

Jensen:2012:FSL

[238] Jonas Braband Jensen and Lars Birkedal. Fictional separation logic. Lecture Notes in Computer Science, 7211:377–396, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28869-2_19/.

Jourdan:2012:VLP

[239] Jacques-Henri Jourdan, François Pottier, and Xavier Leroy. Validating LR (1) parsers. Lecture Notes in Computer Science, 7211:397–416, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28869-2_20/.

Anonymous:2012:FMp

[240] Anonymous. Front matter. Lecture Notes in Computer Science, 7211: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28869-2/1.

vanderAalst:2012:DPD

[241] Wil M. P. van der Aalst. Distributed process discovery and conformance checking. Lecture Notes in Computer Science, 7212:1–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28872-2_1/.

DiRuscio:2012:MDT

[242] Davide Di Ruscio, Ivano Malavolta, Henry Muccini, and Patrizio Pelliccione. Model-driven techniques to enhance architectural languages interoperability. Lecture Notes in Computer Science, 7212:26– 42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28872-2_2/.

Bauer:2012:MSC

[243] Sebastian S. Bauer, Alexandre David, Rolf Hennicker, Kim Guldstrand Larsen, and Axel Legay. Moving from specifications to contracts in component-based design. Lecture Notes in Computer Science, 7212: 43–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28872-2_3/.

Bae:2012:ST

[244] Kyungmin Bae, Peter Csaba Ölveczky, José Meseguer, and Abdullah Al-Nayeem. The SynchAADL2Maude tool. Lecture Notes in Computer Science, 7212:59–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-28872-2_4/.

Fiadeiro:2012:CSC

[245] José Luiz Fiadeiro and Antónia Lopes. Consistency of service composition. Lecture Notes in Computer Science, 7212:63–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28872-2_5/.

Eckhardt:2012:SAU

[246] Jonas Eckhardt, Tobias Mühlbauer, Musab AlTurki, José Meseguer, and Martin Wirsing. Stable availability under denial of service attacks through formal patterns. Lecture Notes in Computer Science, 7212: 78–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28872-2_6/.

Naujokat:2012:LPP

[247] Stefan Naujokat, Anna-Lena Lamprecht, and Bernhard Steffen. Loose programming with PROPHETS. Lecture Notes in Computer Science, 7212:94–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28872-2_7/.

Kahlon:2012:SIR

[248] Vineet Kahlon. Schedule insensitivity reduction. Lecture Notes in Computer Science, 7212:99–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28872-2_8/.

Hatvani:2012:ATA

[249] Leo Hatvani, Paul Pettersson, and Cristina Seceleanu. Adaptive task automata: a framework for verifying adaptive embedded systems. Lecture Notes in Computer Science, 7212:115–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28872-2_9/.

Albert:2012:VRG

[250] Elvira Albert, Richard Bubel, Samir Genaim, Reiner Hähnle, and Guillermo Román-Díez. Verified resource guarantees for heap manipulating programs. Lecture Notes in Computer Science, 7212:130–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28872-2_10/.

Maggi:2012:ODS

[251] Fabrizio Maria Maggi, Marco Montali, and Wil M. P. van der Aalst. An operational decision support framework for monitoring business constraints. Lecture Notes in Computer Science, 7212:146–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28872-2_11/.

Diskin:2012:IQK

[252] Zinovy Diskin, Tom Maibaum, and Krzysztof Czarnecki. Intermodeling, queries, and Kleisli categories. Lecture Notes in Computer Science, 7212: 163–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28872-2_12/.

Hermann:2012:CMS

[253] Frank Hermann, Hartmut Ehrig, Claudia Ermel, and Fernando Orejas. Concurrent model synchronization with conflict resolution based on triple graph grammars. Lecture Notes in Computer Science, 7212: 178–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28872-2_13/.

Bradfield:2012:RCQ

[254] Julian Bradfield and Perdita Stevens. Recursive checkonly QVT-R transformations with general when and where clauses via the modal mu calculus. Lecture Notes in Computer Science, 7212:194–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28872-2_14/.

deMol:2012:GTJ

[255] Maarten de Mol, Arend Rensink, and James J. Hunt. Graph transforming Java data. Lecture Notes in Computer Science, 7212:209–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28872-2_15/.

Salay:2012:LIR

[256] Rick Salay, Michalis Famelis, and Marsha Chechik. Language independent refinement using partial modeling. Lecture Notes in Computer Science, 7212:224–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-28872-2_16/.

Bruni:2012:CFA

[257] Roberto Bruni, Andrea Corradini, Fabio Gadducci, Alberto Lluch Lafuente, and Andrea Vandin. A conceptual framework for adaptation. Lecture Notes in Computer Science, 7212:240–254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28872-2_17/.

Thum:2012:ADC

[258] Thomas Thüm, Ina Schaefer, Martin Kuhlemann, Sven Apel, and Gunter Saake. Applying design by contract to feature-oriented programming. Lecture Notes in Computer Science, 7212:255–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28872-2_18/.

Shi:2012:ITS

[259] Jiangfan Shi, Myra B. Cohen, and Matthew B. Dwyer. Integration testing of software product lines using compositional symbolic execution. Lecture Notes in Computer Science, 7212:270–284, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28872-2_19/.

Rubin:2012:CRP

[260] Julia Rubin and Marsha Chechik. Combining related products into product lines. Lecture Notes in Computer Science, 7212:285–300, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28872-2_20/.

Anonymous:2012:FMq

[261] Anonymous. Front matter. Lecture Notes in Computer Science, 7212: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28872-2/1.

Cali:2012:QUC

[262] Andrea Calì, Georg Gottlob, and Giorgio Orsi. Querying UML class diagrams. Lecture Notes in Computer Science, 7213:1–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28729-9_1/.

Winskel:2012:BCG

[263] Glynn Winskel. Bicategories of concurrent games. Lecture Notes in Computer Science, 7213:26–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28729-9_2/.

Atkey:2012:FIM

[264] Robert Atkey, Neil Ghani, and Bart Jacobs. Fibrational induction meets effects. Lecture Notes in Computer Science, 7213:42–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28729-9_3/.

Adamek:2012:CPM

[265] Jiří Adámek and Filippo Bonchi. A coalgebraic perspective on minimization and determinization. Lecture Notes in Computer Science, 7213: 58–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28729-9_4/.

Ahman:2012:WCC

[266] Danel Ahman, James Chapman, and Tarmo Uustalu. When is a container a comonad? Lecture Notes in Computer Science, 7213:74–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28729-9_5/.

Adamek:2012:WPC

[267] Jiří Adámek, Stefan Milius, and Lawrence S. Moss. Well-pointed coalgebras (extended abstract). Lecture Notes in Computer Science, 7213: 89–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28729-9_6/.

Bove:2012:CIA

[268] Ana Bove and Peter Dybjer. Combining interactive and automatic reasoning in first order theories of functional programs. Lecture Notes in Computer Science, 7213:104–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28729-9_7/.

Biernacki:2012:ABD

[269] Dariusz Biernacki and Sergueï Lenglet. Applicative bisimulations for delimited-control operators. Lecture Notes in Computer Science, 7213: 119–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28729-9_8/.

Preugschat:2012:ECS

[270] Sebastian Preugschat and Thomas Wilke. Effective characterizations of simple fragments of temporal logic using prophetic automata. Lecture Notes in Computer Science, 7213:135–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28729-9_9/.

Breuers:2012:IRB

[271] Stefan Breuers and Christof Löding. Improved Ramsey-based Büchi complementation. Lecture Notes in Computer Science, 7213:150–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28729-9_10/.

Seidl:2012:ECP

[272] Helmut Seidl and Andreas Reuß. Extending ⟨_{1}-clauses with path disequalities. Lecture Notes in Computer Science, 7213:165–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28729-9_11/.

Jagadeesan:2012:BRA

[273] Radha Jagadeesan, Gustavo Petri, and James Riely. Brookes is relaxed, almost! Lecture Notes in Computer Science, 7213:180–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28729-9_12/.

Bernardo:2012:RTT

[274] Marco Bernardo and Rocco De Nicola. Revisiting trace and testing equivalences for nondeterministic and probabilistic processes. Lecture Notes in Computer Science, 7213:195–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28729-9_13/.

Peters:2012:IGE

[275] Kirstin Peters. Is it a “good” encoding of mixed choice? Lecture Notes in Computer Science, 7213:210–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28729-9_14/.

Crafa:2012:ESS

[276] Silvia Crafa and Daniele Varacca. Event structure semantics of parallel extrusion in the pi-calculus. Lecture Notes in Computer Science, 7213: 225–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28729-9_15/.

Gorin:2012:NES

[277] Daniel Gorín and Lutz Schröder. Narcissists are easy, stepmothers are hard. Lecture Notes in Computer Science, 7213:240–254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28729-9_16/.

Kurz:2012:NRL

[278] Alexander Kurz and Tomoyuki Suzuki. On nominal regular languages with binders. Lecture Notes in Computer Science, 7213:255–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28729-9_17/.

Chatterjee:2012:RSE

[279] Krishnendu Chatterjee. Robustness of structurally equivalent concurrent parity games. Lecture Notes in Computer Science, 7213:270–285, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28729-9_18/.

Brihaye:2012:SPE

[280] Thomas Brihaye and Véronique Bruyère. Subgame perfection for equilibria in quantitative reachability games. Lecture Notes in Computer Science, 7213:286–300, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28729-9_19/.

Bouyer:2012:CGO

[281] Patricia Bouyer and Romain Brenguier. Concurrent games with ordered objectives. Lecture Notes in Computer Science, 7213:301–315, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28729-9_20/.

Anonymous:2012:FMr

[282] Anonymous. Front matter. Lecture Notes in Computer Science, 7213: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28729-9/1.

Hermanns:2012:QMD

[283] Holger Hermanns. Quantitative models for a not so dumb grid. Lecture Notes in Computer Science, 7214:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-28756-5_1.

Zaeem:2012:HAD

[284] Razieh Nokhbeh Zaeem, Divya Gopinath, Sarfraz Khurshid, and Kathryn S. McKinley. History-aware data structure repair using SAT. Lecture Notes in Computer Science, 7214:2–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28756-5_2/.

Hardin:2012:GLV

[285] David Hardin, Konrad Slind, Michael Whalen, and Tuan-Hung Pham. The guardol language and verification system. Lecture Notes in Computer Science, 7214:18–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28756-5_3/.

Cox:2012:BTP

[286] Arlen Cox, Sriram Sankaranarayanan, and Bor-Yuh Evan Chang. A bit too precise? Bounded verification of quantized digital filters. Lecture Notes in Computer Science, 7214:33–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28756-5_4/.

DSilva:2012:NBA

[287] Vijay D’Silva, Leopold Haller, Daniel Kroening, and Michael Tautschnig. Numeric bounds analysis with conflict-driven learning. Lecture Notes in Computer Science, 7214:48–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28756-5_5/.

Friedmann:2012:RBA

[288] Oliver Friedmann and Martin Lange. Ramsey-based analysis of parity automata. Lecture Notes in Computer Science, 7214:64–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28756-5_6/.

Lengal:2012:VLE

[289] Ondřej Lengál, Jiří Šimáček, and Tomáš Vojnar. VATA: a library for efficient manipulation of non-deterministic tree automata. Lecture Notes in Computer Science, 7214:79–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28756-5_7/.

Babiak:2012:LBA

[290] Tomáš Babiak, Mojmír Křetínský, Vojtěch Řehák, and Jan Strejček. LTL to Büchi automata translation: Fast and more deterministic. Lecture Notes in Computer Science, 7214:95–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28756-5_8/.

Song:2012:PMC

[291] Fu Song and Tayssir Touili. Pushdown model checking for malware detection. Lecture Notes in Computer Science, 7214:110–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28756-5_9/.

Hamlen:2012:AOR

[292] Kevin W. Hamlen, Micah M. Jones, and Meera Sridhar. Aspect-oriented runtime monitor certification. Lecture Notes in Computer Science, 7214: 126–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28756-5_10/.

Lang:2012:PMC

[293] Frédéric Lang and Radu Mateescu. Partial model checking using networks of labelled transition systems and Boolean equation systems. Lecture Notes in Computer Science, 7214:141–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28756-5_11/.

Albarghouthi:2012:UAA

[294] Aws Albarghouthi, Arie Gurfinkel, and Marsha Chechik. From underapproximations to over-approximations and back. Lecture Notes in Computer Science, 7214:157–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28756-5_12/.

Fehnker:2012:AAA

[295] Ansgar Fehnker, Rob van Glabbeek, Peter Höfner, Annabelle McIver, and Marius Portmann. Automated analysis of AODV using UPPAAL. Lecture Notes in Computer Science, 7214:173–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28756-5_13/.

Jiang:2012:MVD

[296] Zhihao Jiang, Miroslav Pajic, Salar Moarref, Rajeev Alur, and Rahul Mangharam. Modeling and verification of a dual chamber implantable pacemaker. Lecture Notes in Computer Science, 7214:188– 203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28756-5_14/.

Abdulla:2012:CEG

[297] Parosh Aziz Abdulla, Mohamed Faouzi Atig, Yu-Fang Chen, Carl Leonardsson, and Ahmed Rezine. Counter-example guided fence insertion under TSO. Lecture Notes in Computer Science, 7214:204–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28756-5_15/.

Jin:2012:JMM

[298] Huafeng Jin, Tuba Yavuz-Kahveci, and Beverly A. Sanders. Java memory model-aware model checking. Lecture Notes in Computer Science, 7214: 220–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28756-5_16/.

Popeea:2012:CTP

[299] Corneliu Popeea and Andrey Rybalchenko. Compositional termination proofs for multi-threaded programs. Lecture Notes in Computer Science, 7214:237–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-28756-5_17/.

Bozga:2012:DCT

[300] Marius Bozga, Radu Iosif, and Filip Konečný. Deciding conditional termination. Lecture Notes in Computer Science, 7214:252–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28756-5_18/.

Armando:2012:APA

[301] Alessandro Armando, Wihem Arsac, Tigran Avanesov, Michele Barletta, and Alberto Calvi. The AVANTSSAR platform for the automated validation of trust and security of service-oriented architectures. Lecture Notes in Computer Science, 7214:267–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28756-5_19/.

Wang:2012:RBF

[302] Anduo Wang, Carolyn Talcott, Alexander J. T. Gurney, Boon Thau Loo, and Andre Scedrov. Reduction-based formal analysis of BGP instances. Lecture Notes in Computer Science, 7214:283–298, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28756-5_20/.

Anonymous:2012:FMs

[303] Anonymous. Front matter. Lecture Notes in Computer Science, 7214: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28756-5/1.

Dwork:2012:DPP

[304] Cynthia Dwork. Differential privacy and the power of (formalizing) negative thinking. Lecture Notes in Computer Science, 7215:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-28641-4_1.

Blanchet:2012:SPV

[305] Bruno Blanchet. Security protocol verification: Symbolic and computational models. Lecture Notes in Computer Science, 7215:3–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28641-4_2/.

Cortier:2012:ARP

[306] Véronique Cortier, Jan Degrieck, and Stéphanie Delaune. Analysing routing protocols: Four nodes topologies are sufficient. Lecture Notes in Computer Science, 7215:30–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28641-4_3/.

Franklin:2012:PVA

[307] Jason Franklin, Sagar Chaki, and Anupam Datta. Parametric verification of address space separation. Lecture Notes in Computer Science, 7215:51–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28641-4_4/.

Paiola:2012:VSP

[308] Miriam Paiola and Bruno Blanchet. Verification of security protocols with lists: From length one to unbounded length. Lecture Notes in Computer Science, 7215:69–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28641-4_5/.

Arapinis:2012:PSC

[309] Myrto Arapinis, Sergiu Bursuc, and Mark Ryan. Privacy supporting cloud computing: ConfiChair, a case study. Lecture Notes in Computer Science, 7215:89–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28641-4_6/.

Cortier:2012:FAN

[310] Véronique Cortier and Cyrille Wiedling. A formal analysis of the Norwegian E-voting protocol. Lecture Notes in Computer Science, 7215: 109–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28641-4_7/.

Basin:2012:PRI

[311] David Basin, Cas Cremers, and Simon Meier. Provably repairing the ISO/ IEC 9798 standard for entity authentication. Lecture Notes in Computer Science, 7215:129–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28641-4_8/.

Comon-Lundh:2012:SPD

[312] Hubert Comon-Lundh, Véronique Cortier, and Guillaume Scerri. Security proof with dishonest keys. Lecture Notes in Computer Science, 7215: 149–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28641-4_9/.

Arapinis:2012:RET

[313] Myrto Arapinis, Sergiu Bursuc, and Mark D. Ryan. Reduction of equational theories for verification of trace equivalence: Re-encryption, associativity and commutativity. Lecture Notes in Computer Science, 7215: 169–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28641-4_10/.

Bana:2012:TUS

[314] Gergei Bana and Hubert Comon-Lundh. Towards unconditional soundness: Computationally complete symbolic attacker. Lecture Notes in Computer Science, 7215:189–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28641-4_11/.

Barthe:2012:VIH

[315] Gilles Barthe, Benjamin Grégoire, and Sylvain Heraud. Verified indifferentiable hashing into elliptic curves. Lecture Notes in Computer Science, 7215:209–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-28641-4_12/.

Datta:2012:PAL

[316] Anupam Datta, Divya Sharma, and Arunesh Sinha. Provable deanonymization of large datasets with sparse dimensions. Lecture Notes in Computer Science, 7215:229–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28641-4_13/.

Yen:2012:RBM

[317] Ting-Fang Yen and Michael K. Reiter. Revisiting botnet models and their implications for takedown strategies. Lecture Notes in Computer Science, 7215:249–268, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28641-4_14/.

Yang:2012:GTA

[318] Mu Yang, Vladimiro Sassone, and Sardaouna Hamadou. A game-theoretic analysis of cooperation in anonymity networks. Lecture Notes in Computer Science, 7215:269–289, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28641-4_15/.

Best:2012:DSD

[319] Eike Best and Philippe Darondeau. Deciding selective declassification of Petri nets. Lecture Notes in Computer Science, 7215:290–308, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28641-4_16/.

Basin:2012:ESP

[320] David Basin, Vincent Jugé, and Felix Klaedtke. Enforceable security policies revisited. Lecture Notes in Computer Science, 7215:309–328, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28641-4_17/.

Milushev:2012:TIH

[321] Dimiter Milushev and Dave Clarke. Towards incrementalization of holistic hyperproperties. Lecture Notes in Computer Science, 7215:329–348, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28641-4_18/.

Centenaro:2012:TBA

[322] Matteo Centenaro, Riccardo Focardi, and Flaminia L. Luccio. Type-based analysis of PKCS#11 key management. Lecture Notes in Computer Science, 7215:349–368, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28641-4_19/.

Amtoft:2012:CIM

[323] Torben Amtoft, Josiah Dodds, Zhi Zhang, and Andrew Appel. A certificate infrastructure for machine-checked proofs of conditional information flow. Lecture Notes in Computer Science, 7215:369–389, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28641-4_20/.

Anonymous:2012:FMt

[324] Anonymous. Front matter. Lecture Notes in Computer Science, 7215: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28641-4/1.

Hadjiantonis:2012:ITE

[325] Antonis M. Hadjiantonis. Introduction to telecommunication economics. Lecture Notes in Computer Science, 7216:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-30382-1_1.

Mendonca:2012:IER

[326] Sandro Mendonça. Introduction to evolution and regulation of communication ecosystems. Lecture Notes in Computer Science, 7216:3– 4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-30382-1_2.

Ramos:2012:ERM

[327] Sergio Ramos, José Luis Gómez-Barroso, and Claudio Feijóo. Evolution and regulation of mobile ecosystems: European information society policies for the mobile search domain. Lecture Notes in Computer Science, 7216:5–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/10.1007/
978-3-642-30382-1_3.

Falch:2012:IRM

[328] Morten Falch. International roaming of mobile services: The need for regulation. Lecture Notes in Computer Science, 7216:14–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30382-1_4.

Kosa:2012:MRF

[329] Zsuzsanna Kósa. Mobile regulation for the future. Lecture Notes in Computer Science, 7216:22–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-30382-1_5.

Mendonca:2012:TTI

[330] Sandro Mendonça. Trademarks as a telecommunications indicator for industrial analysis and policy. Lecture Notes in Computer Science, 7216:

33–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-30382-1_6.

Hadjiantonis:2012:ISP

[331] Antonis M. Hadjiantonis. Introduction to social and policy implications of communication technologies. Lecture Notes in Computer Science, 7216: 42–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-30382-1_7.

Pau:2012:MCN

[332] Louis-François Pau, Pedro Puga, Hong Chen, and Zviad Kirtava. Mobile communications of the needy and poor: Affordability indicators, European data, social tariffs. Lecture Notes in Computer Science, 7216: 44–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-30382-1_8.

Vannieuwenborg:2012:IMS

[333] Frederic Vannieuwenborg, Zviad Kirtava, Lambros Lambrinos, and Jan Van Ooteghem. Implications of mHealth service deployments: a comparison between dissimilar European countries. Lecture Notes in Computer Science, 7216:56–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/pdf/
10.1007/978-3-642-30382-1_9.

Shwarts-Asher:2012:SCB

[334] Daphna Shwarts-Asher. Social communication behaviors of virtual leaders. Lecture Notes in Computer Science, 7216:67–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/978-3-642-30382-1_10.

Sapio:2012:PID

[335] Bartolomeo Sapio. Policy implications of digital television. Lecture Notes in Computer Science, 7216:78–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-30382-1_11.

Tuffin:2012:IEG

[336] Bruno Tuffin. Introduction to economics and governance of future networks. Lecture Notes in Computer Science, 7216:89–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30382-1_12.

Gomez-Barroso:2012:PID

[337] José Luis Gómez-Barroso and Claudio Feijóo. Public intervention in the deployment of NGNs. Lecture Notes in Computer Science, 7216:91– 99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-30382-1_13.

Williams:2012:PPP

[338] Idongesit Williams and Morten Falch. Public private partnerships and next generation networks. Lecture Notes in Computer Science, 7216:100– 107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-30382-1_14.

Maille:2012:IGE

[339] Patrick Maillé, Peter Reichl, and Bruno Tuffin. Internet governance and economics of network neutrality. Lecture Notes in Computer Science, 7216: 108–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/10.1007/
978-3-642-30382-1_15.

Pau:2012:GNG

[340] Louis-François Pau. Green networks and green tariffs as driven by user service demand. Lecture Notes in Computer Science, 7216:117– 125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-30382-1_16.

Falch:2012:EII

[341] Morten Falch. Environmental impact of ICT on the transport sector. Lecture Notes in Computer Science, 7216:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-642-30382-1_17.

Verbrugge:2012:CEN

[342] Sofie Verbrugge, Jan Van Ooteghem, Koen Casier, Marlies Van der Wee, and Mathieu Tahon. Cost-efficient NGN rollout. Lecture Notes in Computer Science, 7216:138–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-30382-1_18.

Frank:2012:SBT

[343] Lauri Frank, Eetu Luoma, Oleksiy Mazhelis, Mirja Pulkkinen, and Pasi Tyrväinen. Software business in the telecommunications sector. Lecture Notes in Computer Science, 7216:148–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-30382-1_19.

Starc:2012:IFN

[344] Iztok Starc. Introduction to future networks management architectures and mechanisms. Lecture Notes in Computer Science, 7216:156–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/978-
3-642-30382-1_20.

Anonymous:2012:FMu

[345] Anonymous. Front matter. Lecture Notes in Computer Science, 7216: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30382-1/1.

Anonymous:2012:FMv

[346] Anonymous. Front matter. Lecture Notes in Computer Science, 7217: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31915-0/1/1.

Khan:2012:LFR

[347] Shakil M. Khan and Yves Lespérance. Logical foundations for a rational BDI agent programming language (extended version). Lecture Notes in Computer Science, 7217:3–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31915-0_1/.

Telang:2012:RGC

[348] Pankaj R. Telang, Munindar P. Singh, and Neil Yorke-Smith. Relating goal and commitment semantics. Lecture Notes in Computer Science, 7217:22–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31915-0_2/.

Anonymous:2012:FMw

[349] Anonymous. Front matter. Lecture Notes in Computer Science, 7217: 39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31915-0/2/1.

Toledo:2012:DKM

[350] Carlos M. Toledo, Rafael H. Bordini, Omar Chiotti, and María R. Galli. Developing a knowledge management multi-agent system using JaCaMo. Lecture Notes in Computer Science, 7217:41–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31915-0_3/.

Pibil:2012:NPA

[351] Radek Píbil, Peter Novák, Cyril Brom, and Jakub Gemrot. Notes on pragmatic agent-programming with jason. Lecture Notes in Computer Science, 7217:58–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31915-0_4/.

Ranathunga:2012:IEM

[352] Surangika Ranathunga, Stephen Cranefield, and Martin Purvis. Integrating expectation monitoring into BDI agents. Lecture Notes in Computer Science, 7217:74–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31915-0_5/.

Anonymous:2012:FMx

[353] Anonymous. Front matter. Lecture Notes in Computer Science, 7217: 93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31915-0/3/1.

Koster:2012:AMC

[354] Michael Köster and Peter Lohmann. Abstraction for model checking modular interpreted systems over ATL. Lecture Notes in Computer Science, 7217:95–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31915-0_6/.

Mohammed:2012:MQQ

[355] Ammar Mohammed and Ulrich Furbach. MAS: Qualitative and quantitative reasoning. Lecture Notes in Computer Science, 7217:114–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31915-0_7/.

Jongmans:2012:SSR

[356] Sung-Shik T. Q. Jongmans, Koen V. Hindriks, and M. Birna van Riemsdijk. State space reduction for model checking agent programs. Lecture Notes in Computer Science, 7217:133–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31915-0_8/.

Anonymous:2012:FMy

[357] Anonymous. Front matter. Lecture Notes in Computer Science, 7217: 153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/bfm:978-
3-642-31915-0/4/1.

Behrens:2012:MAP

[358] Tristan Behrens, Michael Köster, Federico Schlesinger, Jürgen Dix, and Jomi F. Hübner. The multi-agent programming contest 2011: a Résumé. Lecture Notes in Computer Science, 7217:155–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-31915-0_9/.

Dekker:2012:HAT

[359] Marc Dekker, Pieter Hameete, Michiel Hegemans, Sebastiaan Leysen, and Joris van den Oever. HactarV2: An agent team strategy based on implicit coordination. Lecture Notes in Computer Science, 7217: 173–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31915-0_10/.

Ettienne:2012:IMA

[360] Mikko Berggren Ettienne, Steen Vester, and Jørgen Villadsen. Implementing a multi-agent system in Python with an auction-based agreement approach. Lecture Notes in Computer Science, 7217:185–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31915-0_11/.

Carr:2012:BS

[361] Dominic Carr, Sean Russell, Balazs Pete, G. M. P. O’Hare, and Rem W. Collier. Bogtrotters in space. Lecture Notes in Computer Science, 7217: 197–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31915-0_12/.

Mirzayi:2012:GDA

[362] Sahar Mirzayi, Vahid Nateghi, and Fatemeh Eskandari. A Gaia-driven approach for competitive multi-agent systems. Lecture Notes in Computer Science, 7217:208–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31915-0_13/.

Anonymous:2012:BMd

[363] Anonymous. Back matter. Lecture Notes in Computer Science, 7217: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31915-0/1.

Anonymous:2012:FMz

[364] Anonymous. Front matter. Lecture Notes in Computer Science, 7217: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31915-0/1.

Schlenker:2012:FII

[365] Philippe Schlenker and Jonathan Lamberton. Formal indices and iconicity in ASL. Lecture Notes in Computer Science, 7218:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31482-7_1/.

Yalcin:2012:CP

[366] Seth Yalcin. Context probabilism. Lecture Notes in Computer Science, 7218:12–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31482-7_2/.

Aher:2012:FCD

[367] Martin Aher. Free choice in Deontic Inquisitive Semantics (DIS). Lecture Notes in Computer Science, 7218:22–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31482-7_3/.

Cooper:2012:NIA

[368] Robin Cooper and Jonathan Ginzburg. Negative inquisitiveness and alternatives-based negation. Lecture Notes in Computer Science, 7218: 32–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31482-7_4/.

Onea:2012:WQC

[369] Edgar Onea and Markus Steinbach. Where question, conditionals and topics converge. Lecture Notes in Computer Science, 7218:42–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31482-7_5/.

Uegaki:2012:IKA

[370] Wataru Uegaki. Inquisitive knowledge attribution and the gettier problem. Lecture Notes in Computer Science, 7218:52–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31482-7_6/.

Barbera:2012:WWS

[371] Gemma Barberà. When wide Scope is not enough: Scope and topicality of discourse referents. Lecture Notes in Computer Science, 7218: 62–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31482-7_7/.

Davidson:2012:WDL

[372] Kathryn Davidson. When disjunction looks like conjunction: Pragmatic consequences in ASL. Lecture Notes in Computer Science, 7218: 72–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31482-7_8/.

Quer:2012:QSA

[373] Josep Quer. Quantificational strategies across language modalities. Lecture Notes in Computer Science, 7218:82–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31482-7_9/.

Wilbur:2012:DMI

[374] Ronnie B. Wilbur, Evie Malaia, and Robin A. Shay. Degree modification and intensification in American sign language adjectives. Lecture Notes in Computer Science, 7218:92–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31482-7_10/.

Abrusan:2012:EKF

[375] Márta Abrusán and Kriszta Szendröi. Experimenting with the King of France. Lecture Notes in Computer Science, 7218:102–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31482-7_11/.

Boleda:2012:ASV

[376] Gemma Boleda, Stefan Evert, Berit Gehrke, and Louise McNally. Adjectives as saturators vs. modifiers: Statistical evidence. Lecture Notes in Computer Science, 7218:112–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31482-7_12/.

Brasoveanu:2012:LSI

[377] Adrian Brasoveanu and Jakub Dotlačil. Licensing sentence-internal readings in English. Lecture Notes in Computer Science, 7218:122–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31482-7_13/.

Bylinina:2012:EAS

[378] Lisa Bylinina and Stas Zadorozhny. Evaluative adjectives, scale structure, and ways of being polite. Lecture Notes in Computer Science, 7218: 133–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31482-7_14/.

Chemla:2012:PFC

[379] Emmanuel Chemla and Lewis Bott. Processing: Free choice at no cost. Lecture Notes in Computer Science, 7218:143–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31482-7_15/.

Djalali:2012:CEP

[380] Alex Djalali, Sven Lauer, and Christopher Potts. Corpus evidence for preference-driven interpretation. Lecture Notes in Computer Science, 7218:150–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31482-7_16/.

Naumann:2012:REE

[381] Ralf Naumann. Relating ERP-effects to theories of belief update and combining systems. Lecture Notes in Computer Science, 7218:160–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31482-7_17/.

Panzeri:2012:CCT

[382] Francesca Panzeri and Francesca Foppolo. Can children tell us something about the semantics of adjectives? Lecture Notes in Computer Science, 7218:170–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31482-7_18/.

Rado:2012:URS

[383] Janina Radó and Oliver Bott. Underspecified representations of Scope ambiguity? Lecture Notes in Computer Science, 7218:180–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31482-7_19/.

Renans:2012:PBN

[384] Agata Maria Renans. Projective behaviour of Nur — quantitative experimental research. Lecture Notes in Computer Science, 7218:190–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31482-7_20/.

Anonymous:2012:FMba

[385] Anonymous. Front matter. Lecture Notes in Computer Science, 7218: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31482-7/1.

Kao:2012:ABC

[386] Fu-Chien Kao, Han-Chien Hsieh, and Wei-Te Li. Analysis of brainwave characteristics for playing heterogeneous computer games. Lecture Notes in Computer Science, 7220:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31439-1_1/.

Zhang:2012:ECS

[387] Li Zhang. Exploitation in context-sensitive affect sensing from improvisational interaction. Lecture Notes in Computer Science, 7220: 12–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31439-1_2/.

Park:2012:HDM

[388] Jonghee Park and Woontack Woo. Hybrid document matching method for page identification of Digilog books. Lecture Notes in Computer Science, 7220:24–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31439-1_3/.

Shiratuddin:2012:GDC

[389] Mohd Fairuz Shiratuddin and Kok Wai Wong. Game design considerations when using non-touch based natural user interface. Lecture Notes in Computer Science, 7220:35–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31439-1_4/.

Wang:2012:ESG

[390] Li-Chun Wang and Ming-Puu Chen. Effects of the sequence of game-play and game-design on novices’ motivation, flow, and performance. Lecture Notes in Computer Science, 7220:46–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31439-1_5/.

Zhang:2012:BTO

[391] Sujing Zhang and Feng Li. Behavioral traits of the online parent-child game players: a case study and its inspirations. Lecture Notes in Computer Science, 7220:56–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31439-1_6/.

Berger:2012:TOS

[392] Florian Berger and Wolfgang Müller. Towards an open source game engine for teaching and research. Lecture Notes in Computer Science, 7220:69–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31439-1_7/.

Chang:2012:DMD

[393] Long-Chyr Chang and Heien-Kun Chiang. Designing a mixed digital signage and multi-touch interaction for social learning. Lecture Notes in Computer Science, 7220:77–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31439-1_8/.

Chiang:2012:ODC

[394] I-Tsun Chiang. Old dogs can learn new tricks: Exploring effective strategies to facilitate somatosensory video games for institutionalized older veterans. Lecture Notes in Computer Science, 7220:88–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31439-1_9/.

Yeh:2012:EMS

[395] Shih-Ching Yeh, Wu-Yuin Hwang, Jing-Liang Wang, and Yuin-Ren Chen. Effects of multi-symbols on enhancing virtual reality based collaborative task. Lecture Notes in Computer Science, 7220:101–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31439-1_10/.

Liu:2012:PFA

[396] Sheng Liu, Ting Fang, Shengyong Chen, Hanyang Tong, Changchun Yuan, and Zichen Chen. Particle filter with affine transformation for multiple key points tracking. Lecture Notes in Computer Science, 7220: 112–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31439-1_11/.

Zeng:2012:MSP

[397] Liang Zeng, Bo Wu, Jiangfan Ning, Jiawen Ma, and Sikun Li. Meshless simulation of plastic deformation and crack. Lecture Notes in Computer Science, 7220:127–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31439-1_12/.

Luo:2012:LVT

[398] Yong Luo, Yan Zhao, Lei Cheng, Jianxin Wang, and Xuchong Liu. Lossless visible three-dimensional watermark of digital elevation model data. Lecture Notes in Computer Science, 7220:138–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31439-1_13/.

Feng:2012:AHT

[399] Zhiquan Feng, Bo Yang, Yi Li, Haokui Tang, Yanwei Zheng, Minming Zhang, and Zhigeng Pan. 3D articulated hand tracking based on behavioral model. Lecture Notes in Computer Science, 7220:148– 165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31439-1_14/.

Lin:2012:SPP

[400] Liping Lin, Yongtian Wang, Katsuhito Akahane, and Makoto Sato. SPIDAR-pen: a 2T1R pen-based interface with co-located haptic-visual display. Lecture Notes in Computer Science, 7220:166–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31439-1_15/.

Zhang:2012:GIF

[401] Liliang Zhang and Zhigeng Pan. Generation of IFS fractal images based on hidden Markov model. Lecture Notes in Computer Science, 7220: 178–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31439-1_16/.

Wang:2012:HPH

[402] Yong Wang and Liangliang Hu. High performance hybrid streaming scheme and VCR support. Lecture Notes in Computer Science, 7220: 188–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31439-1_17/.

Huang:2012:MIS

[403] Xianying Huang and Weiwei Chen. A modular image search engine based on key words and color features. Lecture Notes in Computer Science, 7220: 200–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31439-1_18/.

Ji:2012:CDR

[404] Gang Ji, Yehua Zhu, and Yongzhi Zhang. The corroded defect rating system of coating material based on computer vision. Lecture Notes in Computer Science, 7220:210–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31439-1_19/.

Zhang:2012:SRA

[405] Yang Zhang and Chuanyun Xu. Service replacement algorithm using complex network. Lecture Notes in Computer Science, 7220:221–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31439-1_20/.

Anonymous:2012:FMbb

[406] Anonymous. Front matter. Lecture Notes in Computer Science, 7220: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31439-1/1.

Feuerlicht:2012:SIW

[407] George Feuerlicht, Howard Foster, Winfried Lamersdorf, and Guadalupe Ortiz. Seventh International Workshop on Engineering Service-Oriented Applications (WESOA 2011). Lecture Notes in Computer Science, 7221: 1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-31875-7_1.

Foster:2012:TCS

[408] Howard Foster and George Spanoudakis. Taming the cloud: Safety, certification and compliance for software services. Lecture Notes in Computer Science, 7221:3–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31875-7_2/.

Morrison:2012:SAB

[409] Evan D. Morrison, Aditya K. Ghose, Hoa K. Dam, Kerry G. Hinge, and Konstantin Hoesch-Klohe. Strategic alignment of business processes. Lecture Notes in Computer Science, 7221:9–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31875-7_3/.

Fernandez:2012:DWC

[410] Héctor Fernández, Cédric Tedeschi, and Thierry Priol. Decentralized workflow coordination through molecular composition. Lecture Notes in Computer Science, 7221:22–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31875-7_4/.

Meilander:2012:ULM

[411] Dominik Meiländer, Antonio Bucchiarone, Cinzia Cappiello, and Elisabetta Di Nitto. Using a lifecycle model for developing and executing realtime online applications on clouds. Lecture Notes in Computer Science, 7221:33–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31875-7_5/.

Lago:2012:PAA

[412] Patricia Lago and Maryam Razavian. A pragmatic approach for analysis and design of service inventories. Lecture Notes in Computer Science, 7221:44–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31875-7_6/.

Lohmann:2012:ACM

[413] Niels Lohmann and Martin Nyolt. Artifact-centric modeling using BPMN. Lecture Notes in Computer Science, 7221:54–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31875-7_7/.

Barkhordarian:2012:MBP

[414] Angineh Barkhordarian, Frederik Demuth, Kristof Hamann, Minh Hoang, and Sonja Weichler. Migratability of BPMN 2.0 process instances. Lecture Notes in Computer Science, 7221:66–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31875-7_8/.

Holotescu:2012:ALS

[415] Casandra Holotescu. Asynchronous learning for service composition. Lecture Notes in Computer Science, 7221:76–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31875-7_9/.

Bartalos:2012:EEA

[416] Peter Bartalos and M. Brian Blake. Engineering energy-aware Web services toward dynamically-green computing. Lecture Notes in Computer Science, 7221:87–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31875-7_10/.

DePaoli:2012:IFW

[417] Flavio De Paoli, Ioan Toma, Carlos Pedrinaci, and Marcel Tilly. Introduction to the Fifth Workshop on Non-Functional Properties and Service Level Agreements Management in Service-Oriented Computing (NFPSLAM-SOC 2011). Lecture Notes in Computer Science, 7221: 97–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31875-7_11/.

Wagner:2012:AQA

[418] Florian Wagner, Fuyuki Ishikawa, and Shinichi Honiden. Applying QoSaware service selection on functionally diverse services. Lecture Notes in Computer Science, 7221:100–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31875-7_12/.

Brahim:2012:SMW

[419] Monia Ben Brahim, Tarak Chaari, Maher Ben Jemaa, and Mohamed Jmaiel. Semantic matching of WS-SecurityPolicy assertions. Lecture Notes in Computer Science, 7221:114–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31875-7_13/.

Klatt:2012:QPS

[420] Benjamin Klatt, Franz Brosch, Zoya Durdik, and Christoph Rathfelder. Quality prediction in service composition frameworks. Lecture Notes in Computer Science, 7221:131–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31875-7_14/.

Zeginis:2012:EEB

[421] Chrysostomos Zeginis, Konstantina Konsolaki, Kyriakos Kritikos, and Dimitris Plexousakis. ECMAF: An event-based cross-layer service monitoring and adaptation framework. Lecture Notes in Computer Science, 7221:147–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31875-7_15/.

Sheng:2012:PSP

[422] Quan Z. Sheng, Cesare Pautasso, and Sonia Ben Mokhtar. Phd symposium preface. Lecture Notes in Computer Science, 7221:162–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-31875-7_16.

Palacios:2012:DSA

[423] Marcos Palacios. Defining an SLA-aware method to test serviceoriented systems. Lecture Notes in Computer Science, 7221:164–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31875-7_17/.

Stocker:2012:DFO

[424] Thomas Stocker. Data flow-oriented process mining to support security audits. Lecture Notes in Computer Science, 7221:171–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31875-7_18/.

Neto:2012:AUW

[425] Waldemar Ferreira Neto. Adaptable UI for Web service composition: a model-driven approach. Lecture Notes in Computer Science, 7221: 177–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31875-7_19/.

Ye:2012:EMB

[426] Zhen Ye. Economic model based cloud service composition. Lecture Notes in Computer Science, 7221:183–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31875-7_20/.

Anonymous:2012:FMbc

[427] Anonymous. Front matter. Lecture Notes in Computer Science, 7221: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31875-7/1.

Winter:2012:FSD

[428] Johannes Winter, Paul Wiegele, Martin Pirker, and Ronald Tögl. A flexible software development and emulation framework for ARM Trust-Zone. Lecture Notes in Computer Science, 7222:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32298-3_1/.

Chen:2012:BGP

[429] Chunhua Chen, Chris J. Mitchell, and Shaohua Tang. Building general purpose security services on trusted computing. Lecture Notes in Computer Science, 7222:16–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32298-3_2/.

Kounga:2012:ESP

[430] Gina Kounga and Liqun Chen. Enforcing sticky policies with TPM and virtualization. Lecture Notes in Computer Science, 7222:32–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32298-3_3/.

Ekberg:2012:MTT

[431] Jan-Erik Ekberg and Sandeep Tamrakar. Mass transit ticketing with NFC mobile phones. Lecture Notes in Computer Science, 7222:48–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32298-3_4/.

Derler:2012:ATN

[432] David Derler, Klaus Potzmader, Johannes Winter, and Kurt Dietrich. Anonymous ticketing for NFC-enabled mobile phones. Lecture Notes in Computer Science, 7222:66–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32298-3_5/.

Yue:2012:SIC

[433] Qinggang Yue, Feng Liu, and Rui Xue. Some improvements to the costbased framework for analyzing denial of service attacks. Lecture Notes in Computer Science, 7222:84–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32298-3_6/.

Li:2012:FDM

[434] Wei Li, Dawu Gu, Zhiqiang Liu, Ya Liu, and Xiaohu Huang. Fault detection of the MacGuffin Cipher against differential fault attack. Lecture Notes in Computer Science, 7222:102–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32298-3_7/.

Zhang:2012:CSSa

[435] Zijian Zhang, Liehuang Zhu, Feng Wang, Lejian Liao, Cong Guo, and Hongyuan Wang. Computationally sound symbolic analysis of EAP-TNC protocol. Lecture Notes in Computer Science, 7222:113–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32298-3_8/.

Jamthagen:2012:TRD

[436] Christopher Jämthagen, Martin Hell, and Ben Smeets. A technique for remote detection of certain virtual machine monitors. Lecture Notes in Computer Science, 7222:129–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32298-3_9/.

Luo:2012:IDE

[437] Yiyuan Luo, Xuejia Lai, and Zheng Gong. Indifferentiability of domain extension modes for hash functions. Lecture Notes in Computer Science, 7222:138–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32298-3_10/.

Halunen:2012:MGB

[438] Kimmo Halunen. Multicollisions and graph-based hash functions. Lecture Notes in Computer Science, 7222:156–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32298-3_11/.

Peng:2012:GFE

[439] Kun Peng. A general, flexible and efficient proof of inclusion and exclusion. Lecture Notes in Computer Science, 7222:168–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32298-3_12/.

Gligoroski:2012:MS

[440] Danilo Gligoroski, Rune Steinsmo Ødegård, Rune Erlend Jensen, and Ludovic Perret. MQQ-SIG. Lecture Notes in Computer Science, 7222: 184–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32298-3_13/.

Matsuo:2012:MAK

[441] Shin’ichiro Matsuo, Daisuke Moriyama, and Moti Yung. Multifactor authenticated key renewal. Lecture Notes in Computer Science, 7222: 204–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32298-3_14/.

Kutylowski:2012:RIS

[442] Mirosław Kutyłowski, Łukasz Krzywiecki, Przemysław Kubiak, and Michał Koza. Restricted identification scheme and Diffie–Hellman linking problem. Lecture Notes in Computer Science, 7222:221–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32298-3_15/.

Hao:2012:MSG

[443] Dong Hao, Avishek Adhikari, and Kouichi Sakurai. Mixed-strategy game based trust management for clustered wireless sensor networks. Lecture Notes in Computer Science, 7222:239–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32298-3_16/.

Bouabana-Tebibel:2012:HCB

[444] Thouraya Bouabana-Tebibel. Hash chains at the basis of a secure reactive routing protocol. Lecture Notes in Computer Science, 7222: 258–270, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32298-3_17/.

Koeberl:2012:EPD

[445] Patrick Koeberl, Jiangtao Li, Roel Maes, Anand Rajan, Claire Vishik, and Marcin Wójcik. Evaluation of a PUF device authentication scheme on a discrete 0.13um SRAM. Lecture Notes in Computer Science, 7222: 271–288, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32298-3_18/.

Cheng:2012:PAI

[446] Pengqi Cheng, Yan Gu, Zihong Lv, Jianfei Wang, Wenlei Zhu, Zhen Chen, and Jiwei Huang. A performance analysis of identity-based encryption schemes. Lecture Notes in Computer Science, 7222:289–303, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32298-3_19/.

Brickell:2012:CDS

[447] Ernie Brickell, Liqun Chen, and Jiangtao Li. A (corrected) DAA scheme using batch proof and verification. Lecture Notes in Computer Science, 7222:304–337, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32298-3_20/.

Anonymous:2012:BMe

[448] Anonymous. Back matter. Lecture Notes in Computer Science, 7222: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32298-3/1.

Anonymous:2012:FMbd

[449] Anonymous. Front matter. Lecture Notes in Computer Science, 7222: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32298-3/1.

Turner:2012:UAE

[450] Alexander P. Turner, Michael A. Lones, Luis A. Fuente, Susan Stepney, and Leo S. Caves. Using artificial epigenetic regulatory networks to control complex tasks within chaotic systems. Lecture Notes in Computer Science, 7223:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28792-3_1/.

Emmrich:2012:GRN

[451] Peter Martin Ferdinand Emmrich, Hannah Elizabeth Roberts, and Vera Pancaldi. A gene regulatory network simulation of heterosis. Lecture Notes in Computer Science, 7223:12–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28792-3_2/.

Jamshidi:2012:CDP

[452] Shahrad Jamshidi, Heike Siebert, and Alexander Bockmayr. Comparing discrete and piecewise affine differential equation models of gene regulatory networks. Lecture Notes in Computer Science, 7223:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28792-3_3/.

Corblin:2012:AIR

[453] Fabien Corblin, Eric Fanchon, Laurent Trilling, Claudine Chaouiya, and Denis Thieffry. Automatic inference of regulatory and dynamical properties from incomplete gene interaction and expression data. Lecture Notes in Computer Science, 7223:25–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28792-3_4/.

Djordjevic:2012:CTP

[454] Marko Djordjevic, Konstantin Severinov, and Magdalena Djordjevic. CRISPR transcript processing: An unusual mechanism for rapid production of desired molecules. Lecture Notes in Computer Science, 7223: 31–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28792-3_5/.

Zabet:2012:CCM

[455] Nicolae Radu Zabet and Boris Adryan. A comprehensive computational model to simulate transcription factor binding in prokaryotes. Lecture Notes in Computer Science, 7223:35–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28792-3_6/.

Fuente:2012:EAS

[456] Luis A. Fuente, Michael A. Lones, Alexander P. Turner, Susan Stepney, and Leo S. Caves. Evolved artificial signalling networks for the control of a conservative complex dynamical system. Lecture Notes in Computer Science, 7223:38–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28792-3_7/.

Care:2012:EMR

[457] Bertrand R. Caré and Hédi A. Soula. The effect of membrane receptor clustering on spatio-temporal cell signalling dynamics. Lecture Notes in Computer Science, 7223:50–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28792-3_8/.

Colinge:2012:SBA

[458] Jacques Colinge, Uwe Rix, Keiryn L. Bennett, and Giulio Superti-Furga. Systems biology analysis of kinase inhibitor protein target profiles in leukemia treatments. Lecture Notes in Computer Science, 7223: 62–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28792-3_9/.

Seaton:2012:MIE

[459] Daniel D. Seaton and J. Krishnan. Multispecific interactions in enzymatic signalling cascades. Lecture Notes in Computer Science, 7223: 67–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28792-3_10/.

Sengupta:2012:RPC

[460] Dhriti Sengupta and Sudip Kundu. Role of physico-chemical properties of amino acids in protein’s structural organization: a network perspective. Lecture Notes in Computer Science, 7223:74–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28792-3_11/.

Feher:2012:TSA

[461] Kristen Feher, Kathrin Jürchott, and Joachim Selbig. Tailored strategies for the analysis of metabolomic data. Lecture Notes in Computer Science, 7223:82–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28792-3_12/.

Klaus:2012:ELS

[462] Gordon Klaus, Kyrre Glette, and Mats Høvin. Evolving locomotion for a simulated 12-DOF quadruped robot. Lecture Notes in Computer Science, 7223:90–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28792-3_13/.

Beyer:2012:PMS

[463] Antje Beyer, Ralf Eberhard, Nir Piterman, Michael O. Hengartner, and Alex Hajnal. Predictive modelling of stem cell differentiation and apoptosis in C. elegans. Lecture Notes in Computer Science, 7223: 99–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28792-3_14/.

Flann:2012:CSD

[464] Nicholas S. Flann, Hamid Mohamadlou, and Gregory J. Podgorski. Criticality of spatiotemporal dynamics in contact mediated pattern formation. Lecture Notes in Computer Science, 7223:105–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28792-3_15/.

MacGregor:2012:VSA

[465] Duncan J. MacGregor, Tom F. Clayton, and Gareth Leng. The Vasopressin system — asynchronous burst firing as a signal encoding mechanism. Lecture Notes in Computer Science, 7223:117–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28792-3_16/.

Pinto:2012:ECC

[466] Thiago M. Pinto, Maria J. Schilstra, and Volker Steuber. The effective calcium/calmodulin concentration determines the sensitivity of CaMKII to the frequency of calcium oscillations. Lecture Notes in Computer Science, 7223:131–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28792-3_17/.

Metaxas:2012:EDT

[467] Alex Metaxas, Reinoud Maex, Volker Steuber, Rod Adams, and Neil Davey. The effect of different types of synaptic plasticity on the performance of associative memory networks with excitatory and inhibitory sub-populations. Lecture Notes in Computer Science, 7223:136–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28792-3_18/.

Stovold:2012:SNR

[468] James Stovold and Simon O’Keefe. Simulating neurons in reactiondiffusion chemistry. Lecture Notes in Computer Science, 7223:143–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28792-3_19/.

Greaves:2012:EES

[469] Richard B. Greaves, Mark Read, Jon Timmis, Paul S. Andrews, and Vipin Kumar. Extending an established simulation: Exploration of the possible effects using a case study in experimental autoimmune encephalomyelitis. Lecture Notes in Computer Science, 7223:150–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28792-3_20/.

Anonymous:2012:FMbe

[470] Anonymous. Front matter. Lecture Notes in Computer Science, 7223: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28792-3/1.

Hollink:2012:EQM

[471] Vera Hollink, Jiyin He, and Arjen de Vries. Explaining query modifications. Lecture Notes in Computer Science, 7224:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28997-2_1/.

Bortnikov:2012:MTQ

[472] Edward Bortnikov, Pinar Donmez, Amit Kagian, and Ronny Lempel. Modeling transactional queries via templates. Lecture Notes in Computer Science, 7224:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28997-2_2/.

Harvey:2012:EQP

[473] Morgan Harvey and David Elsweiler. Exploring query patterns in email search. Lecture Notes in Computer Science, 7224:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28997-2_3/.

Diriye:2012:ISS

[474] Abdigani Diriye, Giridhar Kumaran, and Jeff Huang. Interactive search support for difficult Web queries. Lecture Notes in Computer Science, 7224:37–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28997-2_4/.

Gaugaz:2012:PFI

[475] Julien Gaugaz, Patrick Siehndel, Gianluca Demartini, Tereza Iofciu, and Mihai Georgescu. Predicting the future impact of news events. Lecture Notes in Computer Science, 7224:50–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28997-2_5/.

Polajnar:2012:DNF

[476] Tamara Polajnar, Richard Glassey, and Leif Azzopardi. Detection of news feeds items appropriate for children. Lecture Notes in Computer Science, 7224:63–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28997-2_6/.

Harvey:2012:CTT

[477] Morgan Harvey, Mark Carman, and David Elsweiler. Comparing tweets and tags for URLs. Lecture Notes in Computer Science, 7224:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28997-2_7/.

Hauff:2012:GLE

[478] Claudia Hauff and Geert-Jan Houben. Geo-location estimation of Flickr images: Social Web based enrichment. Lecture Notes in Computer Science, 7224:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28997-2_8/.

Kim:2012:FRM

[479] Jin Young Kim and W. Bruce Croft. A field relevance model for structured document retrieval. Lecture Notes in Computer Science, 7224: 97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28997-2_9/.

Wu:2012:RBT

[480] Hao Wu and Hui Fang. Relation based term weighting regularization. Lecture Notes in Computer Science, 7224:109–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28997-2_10/.

Yan:2012:NAA

[481] Zhenlei Yan and Jie Zhou. A new approach to answerer recommendation in community question answering services. Lecture Notes in Computer Science, 7224:121–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28997-2_11/.

Neumayer:2012:MEA

[482] Robert Neumayer, Krisztian Balog, and Kjetil Nørvåg. On the modeling of entities for ad-hoc entity search in the Web of data. Lecture Notes in Computer Science, 7224:133–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28997-2_12/.

Berendsen:2012:RDW

[483] Richard Berendsen, Bogomil Kovachev, Evangelia-Paraskevi Nastou, and Maarten de Rijke. Result disambiguation in Web people search. Lecture Notes in Computer Science, 7224:146–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28997-2_13/.

Robertson:2012:SAP

[484] Stephen Robertson. On smoothing average precision. Lecture Notes in Computer Science, 7224:158–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28997-2_14/.

Eskevich:2012:NMM

[485] Maria Eskevich, Walid Magdy, and Gareth J. F. Jones. New metrics for meaningful evaluation of informally structured speech retrieval. Lecture Notes in Computer Science, 7224:170–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28997-2_15/.

Hosseini:2012:ALM

[486] Mehdi Hosseini, Ingemar J. Cox, Nataša Milić-Frayling, Gabriella Kazai, and Vishwa Vinay. On aggregating labels from multiple crowd workers to infer relevance of documents. Lecture Notes in Computer Science, 7224: 182–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28997-2_16/.

Lucchese:2012:HRW

[487] Claudio Lucchese, Raffaele Perego, Fabrizio Silvestri, and Hossein Vahabi. How random walks can help tourism. Lecture Notes in Computer Science, 7224:195–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-28997-2_17/.

Nawab:2012:RCP

[488] Rao Muhammad Adeel Nawab, Mark Stevenson, and Paul Clough. Retrieving candidate plagiarised documents using query expansion. Lecture Notes in Computer Science, 7224:207–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28997-2_18/.

Sondhi:2012:RPW

[489] Parikshit Sondhi, V. G. Vinod Vydiswaran, and ChengXiang Zhai. Reliability prediction of Webpages in the medical domain. Lecture Notes in Computer Science, 7224:219–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28997-2_19/.

Tam:2012:AFE

[490] Tony Tam, Artur Ferreira, and André Lourenço. Automatic foldering of email messages:A combination approach. Lecture Notes in Computer Science, 7224:232–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28997-2_20/.

Anonymous:2012:FMbf

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28997-2/1.

Gallagher:2012:ALP

[492] John P. Gallagher. Analysis of logic programs using regular tree languages. Lecture Notes in Computer Science, 7225:1–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32211-2_1/.

Henglein:2012:DSC

[493] Fritz Henglein. Dynamic symbolic computation for domain-specific language implementation. Lecture Notes in Computer Science, 7225:4–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32211-2_2/.

Albert:2012:RDC

[494] Elvira Albert and Miguel Gómez-Zamalloa. Resource-driven CLP-based test case generation. Lecture Notes in Computer Science, 7225:25–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32211-2_3/.

Almendros-Jimenez:2012:DEX

[495] Jesús M. Almendros-Jiménez and Rafael Caballero. A declarative embedding of XQuery in a functional-logic language. Lecture Notes in Computer Science, 7225:42–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32211-2_4/.

Cabalar:2012:ABC

[496] Pedro Cabalar and Stéphane Demri. Automata-based computation of temporal equilibrium models. Lecture Notes in Computer Science, 7225: 57–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32211-2_5/.

Caballero:2012:SQM

[497] Rafael Caballero and Adrián Riesco. Simplifying questions in Maude declarative debugger by transforming proof trees. Lecture Notes in Computer Science, 7225:73–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32211-2_6/.

Dandois:2012:CLP

[498] Céline Dandois and Wim Vanhoof. Clones in logic programs and how to detect them. Lecture Notes in Computer Science, 7225:90–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32211-2_7/.

Fioravanti:2012:URR

[499] Fabio Fioravanti and Alberto Pettorossi. Using real relaxations during program specialization. Lecture Notes in Computer Science, 7225: 106–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32211-2_8/.

Hidaka:2012:MDO

[500] Soichiro Hidaka and Zhenjiang Hu. Marker-directed optimization of Un-CAL graph transformations. Lecture Notes in Computer Science, 7225: 123–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32211-2_9/.

Morales:2012:MEM

[501] José F. Morales and Manuel V. Hermenegildo. Modular extensions for modular (logic) languages. Lecture Notes in Computer Science, 7225: 139–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32211-2_10/.

Moura:2012:MPS

[502] Paulo Moura. Meta-predicate semantics. Lecture Notes in Computer Science, 7225:155–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32211-2_11/.

Fernandez:2012:SLG

[503] Maribel Fernández and Hélène Kirchner. A strategy language for graph rewriting. Lecture Notes in Computer Science, 7225:173–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32211-2_12/.

Pilozzi:2012:ITA

[504] Paolo Pilozzi and Danny De Schreye. Improved termination analysis of CHR using self-sustainability analysis. Lecture Notes in Computer Science, 7225:189–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32211-2_13/.

Seki:2012:PPC

[505] Hirohisa Seki. Proving properties of co-logic programs by unfold/fold transformations. Lecture Notes in Computer Science, 7225:205–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32211-2_14/.

Sneyers:2012:PTC

[506] Jon Sneyers and Danny De Schreye. Probabilistic termination of CHRiSM programs. Lecture Notes in Computer Science, 7225:221–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32211-2_15/.

Stroder:2012:LOS

[507] Thomas Ströder and Fabian Emmes. A linear operational semantics for termination and complexity analysis of ISO Prolog. Lecture Notes in Computer Science, 7225:237–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32211-2_16/.

Anonymous:2012:BMf

[508] Anonymous. Back matter. Lecture Notes in Computer Science, 7225: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32211-2/1.

Anonymous:2012:FMbg

[509] Anonymous. Front matter. Lecture Notes in Computer Science, 7225: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32211-2/1.

Tinelli:2012:SBM

[510] Cesare Tinelli. SMT-based model checking. Lecture Notes in Computer Science, 7226:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-28891-3_1.

Appel:2012:VST

[511] Andrew W. Appel. Verified software toolchain. Lecture Notes in Computer Science, 7226:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-28891-3_2.

Cousot:2012:FVA

[512] Patrick Cousot. Formal verification by abstract interpretation. Lecture Notes in Computer Science, 7226:3–7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28891-3_3/.

Guck:2012:QTA

[513] Dennis Guck, Tingting Han, Joost-Pieter Katoen, and Martin R. Neuhäußer. Quantitative timed analysis of interactive Markov chains. Lecture Notes in Computer Science, 7226:8–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28891-3_4/.

Ferrari:2012:LLA

[514] Alessio Ferrari, Alessandro Fantechi, and Stefania Gnesi. Lessons learnt from the adoption of formal model-based development. Lecture Notes in Computer Science, 7226:24–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28891-3_5/.

Zurowska:2012:SEC

[515] Karolina Zurowska and Juergen Dingel. Symbolic execution of communicating and hierarchically composed UML-RT state machines. Lecture Notes in Computer Science, 7226:39–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28891-3_6/.

Brauer:2012:IDC

[516] Jörg Brauer and Axel Simon. Inferring definite counterexamples through under-approximation. Lecture Notes in Computer Science, 7226:54–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28891-3_7/.

Gore:2012:MTS

[517] Ross Gore and Paul F. Reynolds, Jr. Modifying test suite composition to enable effective predicate-level statistical debugging. Lecture Notes in Computer Science, 7226:70–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28891-3_8/.

Brisebarre:2012:RPA

[518] Nicolas Brisebarre, Mioara Joldeş, Érik Martin-Dorel, and Micaela Mayero. Rigorous polynomial approximation using Taylor models in Coq. Lecture Notes in Computer Science, 7226:85–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28891-3_9/.

Andre:2012:EIM

[519] Étienne André, Laurent Fribourg, and Romain Soulat. Enhancing the inverse method with state merging. Lecture Notes in Computer Science, 7226:100–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-28891-3_10/.

Herz:2012:CMC

[520] Alexander Herz and Kalmer Apinis. Class-modular, class-escape and points-to analysis for object-oriented languages. Lecture Notes in Computer Science, 7226:106–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28891-3_11/.

Cuoq:2012:TSA

[521] Pascal Cuoq, Benjamin Monate, Anne Pacalet, and Virgile Prevosto. Testing static analyzers with randomly generated programs. Lecture Notes in Computer Science, 7226:120–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28891-3_12/.

Cofer:2012:CVA

[522] Darren Cofer, Andrew Gacek, Steven Miller, and Michael W. Whalen. Compositional verification of architectural models. Lecture Notes in Computer Science, 7226:126–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28891-3_13/.

Ayoub:2012:SCP

[523] Anaheed Ayoub, BaekGyu Kim, Insup Lee, and Oleg Sokolsky. A safety case pattern for model-based development approach. Lecture Notes in Computer Science, 7226:141–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28891-3_14/.

Herencia-Zapana:2012:PLA

[524] Heber Herencia-Zapana, Romain Jobredeaux, and Sam Owre. PVS linear algebra libraries for verification of control software algorithms in C/ ACSL. Lecture Notes in Computer Science, 7226:147–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28891-3_15/.

Li:2012:TAL

[525] Wenbin Li, Jane Huffman Hayes, and Mirosław Truszczyński. Temporal action language (TAL): a controlled language for consistency checking of natural language temporal requirements. Lecture Notes in Computer Science, 7226:162–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28891-3_16/.

Muller:2012:SSV

[526] Norbert Th. Müller and Christian Uhrhan. Some steps into verification of exact real arithmetic. Lecture Notes in Computer Science, 7226: 168–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28891-3_17/.

Bauer:2012:RVM

[527] Andreas Bauer, Jan-Christoph Küster, and Gil Vegliach. Runtime verification meets Android security. Lecture Notes in Computer Science, 7226: 174–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28891-3_18/.

Liu:2012:SPR

[528] Xinxin Liu and Bingtian Xue. Specification in PDL with recursion. Lecture Notes in Computer Science, 7226:181–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28891-3_19/.

Tagore:2012:APT

[529] Aditi Tagore, Diego Zaccai, and Bruce W. Weide. Automatically proving thousands of verification conditions using an SMT solver: An empirical study. Lecture Notes in Computer Science, 7226:195–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28891-3_20/.

Anonymous:2012:FMbh

[530] Anonymous. Front matter. Lecture Notes in Computer Science, 7226: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28891-3/1.

Lakkaraju:2012:CUC

[531] Kiran Lakkaraju, Samarth Swarup, and Les Gasser. Consensus under constraints: Modeling the great English vowel shift. Lecture Notes in Computer Science, 7227:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29047-3_1/.

Chuang:2012:SIS

[532] Katherine Y. Chuang and Christopher C. Yang. A study of informational support exchanges in MedHelp alcoholism community. Lecture Notes in Computer Science, 7227:9–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29047-3_2/.

Abbasi:2012:RWB

[533] Mohammad-Ali Abbasi, Sun-Ki Chai, and Huan Liu. Real-world behavior analysis through a social media lens. Lecture Notes in Computer Science, 7227:18–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29047-3_3/.

McCubbins:2012:MGT

[534] Mathew D. McCubbins and Mark Turner. The mythology of game theory. Lecture Notes in Computer Science, 7227:27–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-29047-3_4/.

Paulhamus:2012:CCD

[535] Barton Paulhamus and Alison Ebaugh. Crowdsourced cyber defense: Lessons from a large-scale, game-based approach to threat identification on a live network. Lecture Notes in Computer Science, 7227: 35–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29047-3_5/.

Altshuler:2012:HMM

[536] Yaniv Altshuler, Michael Fire, and Nadav Aharony. How many makes a crowd? On the evolution of learning as a factor of community coverage. Lecture Notes in Computer Science, 7227:43–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29047-3_6/.

Agrawal:2012:CCT

[537] Kalin Agrawal and William H. Batchelder. Cultural consensus theory: Aggregating signed graphs under a balance constraint. Lecture Notes in Computer Science, 7227:53–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29047-3_7/.

Melamed:2012:UOS

[538] David Melamed, Eric Schoon, and Ronald Breiger. Using organizational similarity to identify statistical interactions for improving situational awareness of CBRN activities. Lecture Notes in Computer Science, 7227:61–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29047-3_8/.

Moore:2012:ODG

[539] Thomas W. Moore and Patrick D. Finley. Opinion dynamics in gendered social networks: An examination of female engagement teams in Afghanistan. Lecture Notes in Computer Science, 7227:69–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29047-3_9/.

Chakraborty:2012:SCE

[540] Subhadeep Chakraborty and Matthew M. Mench. Socio-cultural evolution of opinion dynamics in networked societies. Lecture Notes in Computer Science, 7227:78–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29047-3_10/.

Zhang:2012:CSSb

[541] Mi Zhang, Christopher C. Yang, and Jiexun Li. A comparative study of smoking cessation intervention programs on social media. Lecture Notes in Computer Science, 7227:87–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29047-3_11/.

Altshuler:2012:TPU

[542] Yaniv Altshuler and Wei Pan. Trends prediction using social diffusion models. Lecture Notes in Computer Science, 7227:97–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29047-3_12/.

Coscia:2012:TDG

[543] Michele Coscia and Fosca Giannotti. Towards democratic group detection in complex networks. Lecture Notes in Computer Science, 7227: 105–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29047-3_13/.

Tuli:2012:ADM

[544] Gaurav Tuli, Madhav Marathe, and S. S. Ravi. Addiction dynamics may explain the slow decline of smoking prevalence. Lecture Notes in Computer Science, 7227:114–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29047-3_14/.

Ahmad:2012:LAT

[545] Muhammad Aurangzeb Ahmad and Zoheb Borbora. Love all, trust a few: Link prediction for trust and psycho-social factors in MMOs. Lecture Notes in Computer Science, 7227:123–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29047-3_15/.

Ohara:2012:EID

[546] Kouzou Ohara, Kazumi Saito, and Masahiro Kimura. Effect of in/ out-degree correlation on influence degree of two contrasting information diffusion models. Lecture Notes in Computer Science, 7227:131– 138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29047-3_16/.

Mancilla-Caceres:2012:CLA

[547] Juan F. Mancilla-Caceres, Wen Pu, and Eyal Amir. A computer-in-the-loop approach for detecting bullies in the classroom. Lecture Notes in Computer Science, 7227:139–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29047-3_17/.

Turnley:2012:BVA

[548] Jessica Glicken Turnley and Peter A. Chew. Beyond validation: Alternative uses and associated assessments of goodness for computational social models. Lecture Notes in Computer Science, 7227:147–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29047-3_18/.

Chen:2012:PRL

[549] Hung-Hsuan Chen and Liang Gou. Predicting recent links in FOAF networks. Lecture Notes in Computer Science, 7227:156–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29047-3_19/.

Fink:2012:MTD

[550] Clayton Fink, Jonathon Kopecky, and Nathan Bos. Mapping the Twitterverse in the developing world: An analysis of social media use in Nigeria. Lecture Notes in Computer Science, 7227:164–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29047-3_20/.

Anonymous:2012:FMbi

[551] Anonymous. Front matter. Lecture Notes in Computer Science, 7227: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29047-3/1.

Al-Assam:2012:ERE

[552] Hisham Al-Assam, Ali Abboud, and Harin Sellahewa. Exploiting relative entropy and quality analysis in cumulative partial biometric fusion. Lecture Notes in Computer Science, 7228:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-31971-6_1/.

Makrushin:2012:IRB

[553] Andrey Makrushin and Tobias Scheidat. Improving reliability of biometric hash generation through the selection of dynamic handwriting features. Lecture Notes in Computer Science, 7228:19–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31971-6_2/.

Gloe:2012:FBF

[554] Thomas Gloe. Feature-based forensic camera model identification. Lecture Notes in Computer Science, 7228:42–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31971-6_3/.

Zmudzinski:2012:WEU

[555] Sascha Zmudzinski and Martin Steinebach. Watermark embedding using audio fingerprinting. Lecture Notes in Computer Science, 7228: 63–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31971-6_4/.

Kraetzer:2012:PCS

[556] Christian Kraetzer and Jana Dittmann. Plausibility considerations on steganalysis as a security mechanism — discussions on the example of audio steganalysis. Lecture Notes in Computer Science, 7228:80–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31971-6_5/.

Anonymous:2012:BMg

[557] Anonymous. Back matter. Lecture Notes in Computer Science, 7228: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31971-6/1.

Anonymous:2012:FMbj

[558] Anonymous. Front matter. Lecture Notes in Computer Science, 7228: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31971-6/1.

Andreka:2012:RKA

[559] Hajnal Andréka, Szabolcs Mikulás, and István Németi. Residuated Kleene algebras. Lecture Notes in Computer Science, 7230:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29485-3_1/.

Bonchi:2012:BAC

[560] Filippo Bonchi, Marcello M. Bonsangue, Jan J. M. M. Rutten, and Alexandra Silva. Brzozowski’s algorithm (co)algebraically. Lecture Notes in Computer Science, 7230:12–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29485-3_2/.

Carmosino:2012:EDC

[561] Marco Carmosino, Neil Immerman, and Charles Jordan. Experimental descriptive complexity. Lecture Notes in Computer Science, 7230: 24–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29485-3_3/.

Chen:2012:MQC

[562] Hubie Chen. Meditations on quantified constraint satisfaction. Lecture Notes in Computer Science, 7230:35–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29485-3_4/.

Donald:2012:CSR

[563] Bruce R. Donald. The compass that steered robotics. Lecture Notes in Computer Science, 7230:50–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29485-3_5/.

Glew:2012:SBQ

[564] Neal Glew. Subtyping for F -bounded quantifiers and equirecursive types. Lecture Notes in Computer Science, 7230:66–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29485-3_6/.

Gorecki:2012:IES

[565] Paweł Górecki and Jerzy Tiuryn. Inferring evolutionary scenarios in the duplication, loss and horizontal gene transfer model. Lecture Notes in Computer Science, 7230:83–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29485-3_7/.

Jeannin:2012:CCS

[566] Jean-Baptiste Jeannin. Capsules and closures: a small-step approach. Lecture Notes in Computer Science, 7230:106–123, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29485-3_8/.

Kreitz:2012:NLF

[567] Christoph Kreitz. Nuprl as logical framework for automating proofs in category theory. Lecture Notes in Computer Science, 7230:124–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29485-3_9/.

Kupke:2012:FCA

[568] Clemens Kupke and Jan J. M. M. Rutten. On the final coalgebra of automatic sequences. Lecture Notes in Computer Science, 7230:149– 164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29485-3_10/.

Michalewski:2012:TCR

[569] Henryk Michalewski and Damian Niwiński. On topological completeness of regular tree languages. Lecture Notes in Computer Science, 7230: 165–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29485-3_11/.

Moss:2012:CLS

[570] Lawrence S. Moss, Erik Wennstrom, and Glen T. Whitney. A complete logical system for the equality of recursive terms for sets. Lecture Notes in Computer Science, 7230:180–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29485-3_12/.

Palsberg:2012:ONC

[571] Jens Palsberg. Overloading is NP-Complete. Lecture Notes in Computer Science, 7230:204–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29485-3_13/.

Knight:2012:CEL

[572] Sophia Knight, Radu Mardare, and Prakash Panangaden. Combining epistemic logic and Hennessy–Milner logic. Lecture Notes in Computer Science, 7230:219–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29485-3_14/.

Parikh:2012:CUG

[573] Rohit Parikh, Çağıl Taşdemir, and Andreas Witzel. Choice and uncertainty in games. Lecture Notes in Computer Science, 7230:244–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29485-3_15/.

Rehof:2012:CIE

[574] Jakob Rehof and Paweł Urzyczyn. The complexity of inhabitation with explicit intersection. Lecture Notes in Computer Science, 7230: 256–270, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29485-3_16/.

Salomaa:2012:SSD

[575] Arto Salomaa. On state sequences defined by reaction systems. Lecture Notes in Computer Science, 7230:271–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29485-3_17/.

Sharp:2012:DC

[576] Alexa Sharp. On distance coloring. Lecture Notes in Computer Science, 7230:283–297, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29485-3_18/.

Winskel:2012:WLD

[577] Glynn Winskel. Winning, losing and drawing in concurrent games with perfect or imperfect information. Lecture Notes in Computer Science, 7230:298–317, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29485-3_19/.

Aboul-Hosn:2012:RTD

[578] Kamal Aboul-Hosn. Reflections on a time with Dexter Kozen. Lecture Notes in Computer Science, 7230:318–321, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29485-3_20/.

Anonymous:2012:FMbk

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29485-3/1.

Goldschmidt:2012:TSI

[580] Leonard Goldschmidt. Telehealth strategies and information technology transform patient care within the U.S. Department of Veterans Affairs. Lecture Notes in Computer Science, 7231:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-29361-0_1.

Steyn:2012:CDK

[581] Michael Steyn. From clinical data to knowledge mining for decision making within hospitals. Lecture Notes in Computer Science, 7231:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-29361-0_2.

Yu:2012:BSE

[582] Xiaodan Yu, Dawn Owens, and Deepak Khazanchi. Building socioemotional environments in metaverses for virtual teams in healthcare: a conceptual exploration. Lecture Notes in Computer Science, 7231: 4–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29361-0_3/.

Loper:2012:IHR

[583] Dortje Löper, Meike Klettke, Ilvio Bruder, and Andreas Heuer. Integrating healthcare-related information using the entity-attribute-value storage model. Lecture Notes in Computer Science, 7231:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29361-0_4/.

Klawonn:2012:MPP

[584] Frank Klawonn, Katja Crull, Akiko Kukita, and Frank Pessler. Median Polish with power transformations as an alternative for the analysis of contingency tables with patient data. Lecture Notes in Computer Science, 7231:25–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29361-0_5/.

Henriques:2012:IAH

[585] Rui Henriques and Cláudia Antunes. An integrated approach for healthcare planning over multi-dimensional data using long-term prediction. Lecture Notes in Computer Science, 7231:36–48, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29361-0_6/.

Zhang:2012:NAR

[586] Hao Lan Zhang, Weitao Jiang, Huiqin Wu, and Libing Shu. A novel automated recognition system based on medical machining CAD models. Lecture Notes in Computer Science, 7231:49–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-29361-0_7/.

Hansen:2012:COH

[587] David Hansen. Challenges and opportunities for health information systems research. Lecture Notes in Computer Science, 7231:60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-29361-0_8.

Chiang:2012:HAG

[588] Stanley Chiang. E-health in Australia — a general practitioner’s perspective. Lecture Notes in Computer Science, 7231:61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-29361-0_9.

Young:2012:MDP

[589] Terry Young. Medical devices: Panel discussion. Lecture Notes in Computer Science, 7231:62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29361-0_10.

Chen:2012:CIS

[590] Guiqiang Chen, Jianjun Liu, and Liqin Wang. Color image sharing method based on Lagrange’s interpolating polynomial. Lecture Notes in Computer Science, 7231:63–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29361-0_11/.

Larios:2012:ACM

[591] E. Larios, Y. Zhang, K. Yan, Z. Di, S. LeDévédec, F. Groffen, and F. J. Verbeek. Automation in cytomics: a modern RDBMS based platform for image analysis and management in high-throughput screening experiments. Lecture Notes in Computer Science, 7231:76–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29361-0_12/.

Salazar-Gonzalez:2012:MRR

[592] Ana Salazar-Gonzalez, Yongmin Li, and Djibril Kaba. MRF reconstruction of retinal images for the optic disc segmentation. Lecture Notes in Computer Science, 7231:88–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29361-0_13/.

Zhao:2012:NCS

[593] Jie Zhao, Li Zhang, Wei Zheng, Hua Tian, Dong mei Hao, and Song hong Wu. Normalized cut segmentation of thyroid tumor image based on fractional derivatives. Lecture Notes in Computer Science, 7231: 100–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29361-0_14/.

Zeng:2012:CNN

[594] Nianyin Zeng, Zidong Wang, Yurong Li, and Min Du. Cellular neural networks for gold immunochromatographic strip image segmentation. Lecture Notes in Computer Science, 7231:110–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29361-0_15/.

Wei:2012:NCM

[595] Dan Wei, Weiwei Zhuang, Qingshan Jiang, and Yanjie Wei. A new classification method for human gene splice site prediction. Lecture Notes in Computer Science, 7231:121–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29361-0_16/.

He:2012:ARA

[596] Jing He, Yanchun Zhang, Guangyan Huang, Yefei Xin, Xiaohui Liu, and Hao Lan Zhang. An association rule analysis framework for complex physiological and genetic data. Lecture Notes in Computer Science, 7231: 131–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29361-0_17/.

Kato:2012:ANS

[597] Jien Kato and Yu Wang. Analysis of nursery school observations for understanding children’s behavior. Lecture Notes in Computer Science, 7231:143–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29361-0_18/.

Ji:2012:EOS

[598] Xiang Ji, Soon Ae Chun, and James Geller. Epidemic outbreak and spread detection system based on Twitter data. Lecture Notes in Computer Science, 7231:152–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29361-0_19/.

Okumura:2012:LAE

[599] Takashi Okumura and Yuka Tateisi. A lightweight approach for extracting disease-symptom relation with MetaMap toward automated generation of disease knowledge base. Lecture Notes in Computer Science, 7231: 164–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29361-0_20/.

Anonymous:2012:BMh

[600] Anonymous. Back matter. Lecture Notes in Computer Science, 7231: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29361-0/1.

Anonymous:2012:FMbl

[601] Anonymous. Front matter. Lecture Notes in Computer Science, 7231: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29361-0/1.

Guo:2012:PCS

[602] Fuchun Guo, Yi Mu, Willy Susilo, and Vijay Varadharajan. A precomputable signature scheme with efficient verification for RFID. Lecture Notes in Computer Science, 7232:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29101-2_1/.

Samelin:2012:RSI

[603] Kai Samelin, Henrich C. Pöhls, Arne Bilzhause, Joachim Posegga, and Hermann de Meer. Redactable signatures for independent removal of structure and content. Lecture Notes in Computer Science, 7232:17– 33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29101-2_2/.

Lu:2012:IEC

[604] Xianhui Lu, Bao Li, Qixiang Mei, and Yamin Liu. Improved efficiency of chosen ciphertext secure encryption from factoring. Lecture Notes in Computer Science, 7232:34–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29101-2_3/.

Gao:2012:DES

[605] Chong zhi Gao, Dongqing Xie, and Baodian Wei. Deniable encryptions secure against adaptive chosen ciphertext attack. Lecture Notes in Computer Science, 7232:46–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29101-2_4/.

Comon-Lundh:2012:CSI

[606] Hubert Comon-Lundh, Masami Hagiya, Yusuke Kawamoto, and Hideki Sakurada. Computational soundness of indistinguishability properties without computable parsing. Lecture Notes in Computer Science, 7232: 63–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29101-2_5/.

Bai:2012:NID

[607] Dongxia Bai and Leibo Li. New impossible differential attacks on Camellia. Lecture Notes in Computer Science, 7232:80–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-29101-2_6/.

Liu:2012:IDA

[608] Ya Liu, Dawu Gu, Zhiqiang Liu, and Wei Li. Impossible differential attacks on reduced-round LBlock. Lecture Notes in Computer Science, 7232: 97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29101-2_7/.

Koyama:2012:NTD

[609] Takuma Koyama, Lei Wang, Yu Sasaki, Kazuo Sakiyama, and Kazuo Ohta. New truncated differential cryptanalysis on 3D block cipher. Lecture Notes in Computer Science, 7232:109–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29101-2_8/.

Corniaux:2012:DOT

[610] Christian L. F. Corniaux and Hossein Ghodosi. T -out-of-n distributed oblivious transfer protocols in non-adaptive and adaptive settings. Lecture Notes in Computer Science, 7232:126–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29101-2_9/.

K:2012:CBO

[611] Preetha Mathew K., Sachin Vasant, Sridhar Venkatesan, and C. Pandu Rangan. A code-based 1-out-of-N oblivious transfer based on McEliece assumptions. Lecture Notes in Computer Science, 7232:144–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29101-2_10/.

Wang:2012:TFG

[612] Lei Wang, Ji Xiang, Jiwu Jing, and Lingchen Zhang. Towards fine-grained access control on browser extensions. Lecture Notes in Computer Science, 7232:158–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29101-2_11/.

Chow:2012:ESC

[613] Yang-Wai Chow and Willy Susilo. Enhanced STE3D-CAP: a novel 3D CAPTCHA family. Lecture Notes in Computer Science, 7232:170–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29101-2_12/.

Filardo:2012:HEV

[614] Nathaniel Wesley Filardo and Giuseppe Ateniese. High-entropy visual identification for touch screen devices. Lecture Notes in Computer Science, 7232:182–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29101-2_13/.

Wen:2012:FSA

[615] Chuah Chai Wen, Edward Dawson, Juan Manuel González Nieto, and Leonie Simpson. A framework for security analysis of key derivation functions. Lecture Notes in Computer Science, 7232:199–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29101-2_14/.

Guo:2012:ETD

[616] Teng Guo, Feng Liu, and ChuanKun Wu. On the equivalence of two definitions of visual cryptography scheme. Lecture Notes in Computer Science, 7232:217–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29101-2_15/.

Shinohara:2012:KLE

[617] Naoyuki Shinohara, Takeshi Shimoyama, Takuya Hayashi, and Tsuyoshi Takagi. Key length estimation of pairing-based cryptosystems using η_{T} pairing. Lecture Notes in Computer Science, 7232:228–244, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29101-2_16/.

Izawa:2012:LIX

[618] Kazuya Izawa, Atsuko Miyaji, and Kazumasa Omote. Lightweight integrity for XOR network coding in wireless sensor networks. Lecture Notes in Computer Science, 7232:245–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29101-2_17/.

Braun:2012:IMS

[619] Johannes Braun, Moritz Horsch, and Alexander Wiesmaier. iPIN and mTAN for secure eID applications. Lecture Notes in Computer Science, 7232:259–276, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29101-2_18/.

Liedel:2012:SDC

[620] Manuel Liedel. Secure distributed computation of the square root and applications. Lecture Notes in Computer Science, 7232:277–288, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29101-2_19/.

Shuo:2012:PKR

[621] Tian Shuo, He Yeping, and Ding Baozeng. Prevent kernel return-oriented programming attacks using hardware virtualization. Lecture Notes in Computer Science, 7232:289–300, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29101-2_20/.

Anonymous:2012:FMbm

[622] Anonymous. Front matter. Lecture Notes in Computer Science, 7232: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29101-2/1.

deSouza:2012:KST

[623] Paulo de Souza. Keynote speech: Technological breakthroughs in environmental sensor networks. Lecture Notes in Computer Science, 7234:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-29426-6_1.

Yang:2012:NFD

[624] Li Yang, Xiedong Cao, Jie Li, Cundang Wei, Shiyong Cao, Dan Zhang, and Zhidi Chen. A new formal description model of network attacking and defence knowledge of oil and gas field SCADA system. Lecture Notes in Computer Science, 7234:2–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29426-6_2/.

Sun:2012:EOC

[625] Xiling Sun, Jiajie Xu, Zhiming Ding, Xu Gao, and Kuien Liu. An efficient overload control strategy in cloud. Lecture Notes in Computer Science, 7234:11–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29426-6_3/.

Cao:2012:GDD

[626] Xiedong Cao, Cundang Wei, Jie Li, Li Yang, Dan Zhang, and Gang Tang. The geological disasters defense expert system of the massive pipeline network SCADA system based on FNN. Lecture Notes in Computer Science, 7234:19–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29426-6_4/.

Yang:2012:TTP

[627] Zhuoluo Yang, Jinguo You, and Min Zhou. TDSC: a two-phase duplicate string compression algorithm. Lecture Notes in Computer Science, 7234:27–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29426-6_5/.

Deng:2012:TPM

[628] Ziqiang Deng, Husheng Liao, and Hongyu Gao. Twig pattern matching running on XML streams. Lecture Notes in Computer Science, 7234: 35–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29426-6_6/.

Li:2012:NME

[629] Xingsen Li, Zhongbiao Xiang, Haolan Zhang, and Zhengxiang Zhu. A novel method for extension transformation knowledge discovering. Lecture Notes in Computer Science, 7234:43–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29426-6_7/.

Liu:2012:FPR

[630] Keyan Liu, Shaohua Song, Ningnan Zhou, and Yanyu Ma. A flexible parallel runtime for large scale block-based matrix multiplication. Lecture Notes in Computer Science, 7234:51–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29426-6_8/.

Li:2012:GDR

[631] Yang Li, Kefu Xu, Jianlong Tan, and Li Guo. Group detection and relation analysis research for Web social network. Lecture Notes in Computer Science, 7234:60–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29426-6_9/.

Guo:2012:WPC

[632] Wei Guo, Lidong Zhai, Li Guo, and Jinqiao Shi. Worm propagation control based on spatial correlation in wireless sensor network. Lecture Notes in Computer Science, 7234:68–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29426-6_10/.

Wu:2012:PTT

[633] Hongchen Wu, Xinjun Wang, Zhaohui Peng, Qingzhong Li, and Lin Lin. PointBurst: Towards a trust-relationship framework for improved social recommendations. Lecture Notes in Computer Science, 7234: 78–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29426-6_11/.

Lee:2012:AFE

[634] Junghoon Lee, Hye-Jin Kim, Gyung-Leen Park, Ho-Young Kwak, and Moo Yong Lee. Analysis framework for electric vehicle sharing systems using vehicle movement data stream. Lecture Notes in Computer Science, 7234:89–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29426-6_12/.

Yang:2012:RCS

[635] Chunyan Yang, Xiaomei Li, and Weihua Li. Research on customer segmentation based on extension classification. Lecture Notes in Computer Science, 7234:95–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29426-6_13/.

Chiou:2012:IMA

[636] Chuang-Kai Chiou and Judy C. R. Tseng. An incremental mining algorithm for association rules based on minimal perfect hashing and pruning. Lecture Notes in Computer Science, 7234:106–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29426-6_14/.

Yokomoto:2012:LBT

[637] Daisuke Yokomoto, Kensaku Makita, Hiroko Suzuki, Daichi Koike, and Takehito Utsuro. LDA-based topic modeling in labeling blog posts with Wikipedia entries. Lecture Notes in Computer Science, 7234:114–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29426-6_15/.

Leung:2012:CBS

[638] Clement. H. C. Leung and Yuanxi Li. The correlation between semantic visual similarity and ontology-based concept similarity in effective Web image search. Lecture Notes in Computer Science, 7234: 125–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29426-6_16/.

Mannava:2012:ADP

[639] Vishnuvardhan Mannava and T. Ramesh. An adaptive design pattern for invocation of synchronous and asynchronous Web services in autonomic computing systems. Lecture Notes in Computer Science, 7234:

131–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29426-6_17/.

Ding:2012:MTL

[640] Zhaoyun Ding, Yan Jia, Bin Zhou, and Yi Han. Mining tribe-leaders based on the frequent pattern of propagation. Lecture Notes in Computer Science, 7234:143–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29426-6_18/.

Guo:2012:RSL

[641] Jianyi Guo, Jun Zhao, Zhengtao Yu, Lei Su, and Nianshu Jiang. Research on semantic label extraction of domain entity relation based on CRF and rules. Lecture Notes in Computer Science, 7234:154–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29426-6_19/.

Lvexing:2012:RCP

[642] Zheng Lvexing, Lv Xueqiang, Liu Kun, and Du Yuncheng. Recognition of Chinese personal names based on CRFs and law of names. Lecture Notes in Computer Science, 7234:163–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29426-6_20/.

Anonymous:2012:FMbn

[643] Anonymous. Front matter. Lecture Notes in Computer Science, 7234: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29426-6/1.

Moyer:2012:SIG

[644] Thomas Moyer, Trent Jaeger, and Patrick McDaniel. Scalable integrityguaranteed AJAX. Lecture Notes in Computer Science, 7235:1–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29253-8_1/.

Varadharajan:2012:STW

[645] Vijay Varadharajan. Security and trust in the Web. Lecture Notes in Computer Science, 7235:20–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-29253-8_2.

Hu:2012:EUD

[646] Haoji Hu, Chaofeng Sha, Xiaoling Wang, and Aoying Zhou. Estimate unlabeled-data-distribution for semi-supervised PU learning. Lecture Notes in Computer Science, 7235:22–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29253-8_3/.

Yamamoto:2012:DSS

[647] Yusuke Yamamoto. Disputed sentence suggestion towards credibilityoriented Web search. Lecture Notes in Computer Science, 7235:34–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29253-8_4/.

Yousaf:2012:EVA

[648] Jamal Yousaf, Juanzi Li, Haisu Zhang, and Lei Hou. Exploration and visualization of administrator network in Wikipedia. Lecture Notes in Computer Science, 7235:46–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29253-8_5/.

Liu:2012:EAA

[649] Jianquan Liu, Hanxiong Chen, Kazutaka Furuse, and Hiroyuki Kitagawa. An efficient algorithm for arbitrary reverse furthest neighbor queries. Lecture Notes in Computer Science, 7235:60–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29253-8_6/.

Wang:2012:MAB

[650] Jiaying Wang, Xiaochun Yang, Bin Wang, and Huaijie Zhu. Memoryaware BWT by segmenting sequences to support subsequence search. Lecture Notes in Computer Science, 7235:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29253-8_7/.

Zhang:2012:MFA

[651] Lijun Zhang, Zhanhuai Li, Qun Chen, Xia Li, Ning Li, and Ying Lou. Mining frequent association tag sequences for clustering XML documents. Lecture Notes in Computer Science, 7235:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29253-8_8/.

Xie:2012:CAP

[652] Haoran Xie, Qing Li, and Xudong Mao. Context-aware personalized search based on user and resource profiles in folksonomies. Lecture Notes in Computer Science, 7235:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29253-8_9/.

Shaikh:2012:DBO

[653] Salman Ahmed Shaikh and Hiroyuki Kitagawa. Distance-based outlier detection on uncertain data of Gaussian distribution. Lecture Notes in Computer Science, 7235:109–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29253-8_10/.

Ni:2012:EKS

[654] Weijian Ni, Tong Liu, and Qingtian Zeng. Extracting keyphrase set with high diversity and coverage using structural SVM. Lecture Notes in Computer Science, 7235:122–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29253-8_11/.

Peng:2012:DMP

[655] Zhuo Peng, Chaokun Wang, Lu Han, Jingchao Hao, and Xiaoping Ou. Discovering the most potential stars in social networks with infra-skyline queries. Lecture Notes in Computer Science, 7235:134–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29253-8_12/.

Cheng:2012:FBI

[656] Bing Cheng, Tianqi Chen, Diyi Yang, Weinan Zhang, Yongqiang Wang, and Yong Yu. Feature based informative model for discriminating favorite items from unrated ones. Lecture Notes in Computer Science, 7235: 146–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29253-8_13/.

Hattori:2012:SMI

[657] Yuki Hattori and Akiyo Nadamoto. Search for minority information from Wikipedia based on similarity of majority information. Lecture Notes in Computer Science, 7235:158–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29253-8_14/.

Du:2012:IAA

[658] Yanhua Du and Xitong Li. An incremental approach to analyzing temporal constraints of workflow processes. Lecture Notes in Computer Science, 7235:170–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29253-8_15/.

Zhang:2012:KFP

[659] Chenjing Zhang, Le Chang, Chaofeng Sha, Xiaoling Wang, and Aoying Zhou. Keywords filtering over probabilistic XML data. Lecture Notes in Computer Science, 7235:183–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29253-8_16/.

Zhang:2012:THS

[660] Qian Zhang, Biao Qin, and Shan Wang. H-tree: a hybrid structure for confidence computation in probabilistic databases. Lecture Notes in Computer Science, 7235:195–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29253-8_17/.

Yokotani:2012:POA

[661] Takuya Yokotani, Hung-Hsuan Huang, and Kyoji Kawagoe. Predicting online auction final prices using time series splitting and clustering. Lecture Notes in Computer Science, 7235:207–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29253-8_18/.

Li:2012:TOR

[662] Lei Li, Wenxing Hong, and Tao Li. Taxonomy-oriented recommendation towards recommendation with stage. Lecture Notes in Computer Science, 7235:219–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29253-8_19/.

Gu:2012:GBA

[663] Xiwu Gu, Ruixuan Li, Kunmei Wen, Bei Peng, and Weijun Xiao. A GPU-based accelerator for Chinese word segmentation. Lecture Notes in Computer Science, 7235:231–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29253-8_20/.

Anonymous:2012:FMbo

[664] Anonymous. Front matter. Lecture Notes in Computer Science, 7235: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29253-8/1.

Yin:2012:EVS

[665] Junjun Yin and James D. Carswell. Effects of variations in 3D spatial search techniques on mobile query speed vs. accuracy. Lecture Notes in Computer Science, 7236:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29247-7_1/.

Aydin:2012:ACA

[666] Betül Aydın, Jérôme Gensel, Sylvie Calabretto, and Bruno Tellez. ARCAMA-3D — a context-aware augmented reality mobile platform for environmental discovery. Lecture Notes in Computer Science, 7236: 17–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29247-7_2/.

Cutugno:2012:SMI

[667] Francesco Cutugno, Vincenza Anna Leano, Gianluca Mignini, and Roberto Rinaldi. Speech and multimodal interaction in mobile GIS search: a case of study. Lecture Notes in Computer Science, 7236: 27–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29247-7_3/.

Filonenko:2012:AUT

[668] Viacheslav Filonenko, Charlie Cullen, and James D. Carswell. Asynchronous ultrasonic trilateration for indoor positioning of mobile phones. Lecture Notes in Computer Science, 7236:33–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29247-7_4/.

Misev:2012:SRC

[669] Dimitar Misev, Mihaela Rusu, and Peter Baumann. A semantic resolver for coordinate reference systems. Lecture Notes in Computer Science, 7236:47–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29247-7_5/.

Cerasuolo:2012:VNV

[670] Francesco Cerasuolo, Francesco Cutugno, and Vincenza Anna Leano. Visualization with a new visual metaphor for hierarchical and stratified temporal domain. Lecture Notes in Computer Science, 7236:57– 71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29247-7_6/.

DeChiara:2012:TWB

[671] Davide De Chiara, Vincenzo Del Fatto, Monica Sebillo, and Genoveffa Tortora. Tag@Map: a Web-based application for visually analyzing geographic information through georeferenced tag clouds. Lecture Notes in Computer Science, 7236:72–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29247-7_7/.

Braga:2012:CUT

[672] Reinaldo Bezerra Braga, Ali Tahir, Michela Bertolotto, and Hervé Martin. Clustering user trajectories to find patterns for social interaction applications. Lecture Notes in Computer Science, 7236:82–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29247-7_8/.

Lerin:2012:MPV

[673] Pablo Martinez Lerin, Daisuke Yamamoto, and Naohisa Takahashi. Making a pictorial and verbal travel trace from a GPS trace. Lecture Notes in Computer Science, 7236:98–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29247-7_9/.

Patroumpas:2012:EPR

[674] Kostas Patroumpas and Timos Sellis. Event processing and real-time monitoring over streaming traffic data. Lecture Notes in Computer Science, 7236:116–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29247-7_10/.

Braga:2012:CAW

[675] Reinaldo Bezerra Braga, Sócrates de Moraes Medeiros da Costa, and Windson Viana de Carvalho. A context-aware Web content generator based on personal tracking. Lecture Notes in Computer Science, 7236: 134–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29247-7_11/.

Ballatore:2012:HSS

[676] Andrea Ballatore, David C. Wilson, and Michela Bertolotto. A holistic semantic similarity measure for viewports in interactive maps. Lecture Notes in Computer Science, 7236:151–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29247-7_12/.

McArdle:2012:RBR

[677] Gavin McArdle, Mathieu Petit, Cyril Ray, and Christophe Claramunt. Recommendations based on region and spatial profiles. Lecture Notes in Computer Science, 7236:167–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29247-7_13/.

Gabillon:2012:EPM

[678] Alban Gabillon and Patrick Capolsini. Enforcing protection mechanisms for geographic data. Lecture Notes in Computer Science, 7236: 185–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29247-7_14/.

Ciampi:2012:ITC

[679] Anna Ciampi, Annalisa Appice, Pietro Guccione, and Donato Malerba. Integrating trend clusters for spatio-temporal interpolation of missing sensor data. Lecture Notes in Computer Science, 7236:203–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29247-7_15/.

Bakillah:2012:DSS

[680] Mohamed Bakillah and Steve H. L. Liang. Discovering sensor services with social network analysis and expanded SQWRL querying. Lecture Notes in Computer Science, 7236:221–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29247-7_16/.

Cinque:2012:LPT

[681] Marcello Cinque, Christian Esposito, and Flavio Frattini. Leveraging power of transmission for range-free localization of tiny sensors. Lecture Notes in Computer Science, 7236:239–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29247-7_17/.

Anonymous:2012:BMi

[682] Anonymous. Back matter. Lecture Notes in Computer Science, 7236: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29247-7/1.

Anonymous:2012:FMbp

[683] Anonymous. Front matter. Lecture Notes in Computer Science, 7236: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29247-7/1.

Srivastava:2012:ERT

[684] Divesh Srivastava. Enabling real time data analysis. Lecture Notes in Computer Science, 7238:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-29038-1_1.

Cha:2012:NPT

[685] Sang Kyun Cha. A new paradigm of thinking and architecture for realtime information processing at fingertips. Lecture Notes in Computer Science, 7238:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-29038-1_2.

Mai:2012:IAH

[686] Hai Thanh Mai, Jaeho Kim, and Myoung Ho Kim. Improving the accuracy of histograms for geographic data objects. Lecture Notes in Computer Science, 7238:3–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29038-1_3/.

Wang:2012:IOA

[687] Yuxiang Wang, Junzhou Luo, Aibo Song, Jiahui Jin, and Fang Dong. Improving online aggregation performance for skewed data distribution. Lecture Notes in Computer Science, 7238:18–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29038-1_4/.

Le:2012:RBA

[688] Thanh-Huy Le, Haytham Elghazel, and Mohand-Saíd Hacid. A relationalbased approach for aggregated search in graph databases. Lecture Notes in Computer Science, 7238:33–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29038-1_5/.

Le:2012:DKS

[689] Van Bao Tran Le, Sebastian Link, and Mozhgan Memari. Discovery of keys from SQL tables. Lecture Notes in Computer Science, 7238: 48–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29038-1_6/.

Ngamakeur:2012:FRA

[690] Kan Ngamakeur, Sira Yongchareon, and Chengfei Liu. A framework for realizing artifact-centric business processes in service-oriented architecture. Lecture Notes in Computer Science, 7238:63–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29038-1_7/.

Ding:2012:AOB

[691] Guohui Ding, Han Dong, and Guoren Wang. Appearance-order-based schema matching. Lecture Notes in Computer Science, 7238:79–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29038-1_8/.

Zhou:2012:FRE

[692] Junfeng Zhou, Zhifeng Bao, Ziyang Chen, and Tok Wang Ling. Fast result enumeration for keyword queries on XML data. Lecture Notes in Computer Science, 7238:95–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29038-1_9/.

Leonardi:2012:SSF

[693] Erwin Leonardi, Sourav S. Bhowmick, and Fengrong Li. Stars on steroids: Fast evaluation of multi-source star twig queries in RDBMS. Lecture Notes in Computer Science, 7238:110–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29038-1_10/.

Liu:2012:UTX

[694] Jixue Liu, Chengfei Liu, Theo Haerder, and Jeffery Xu Yu. Updating typical XML views. Lecture Notes in Computer Science, 7238: 126–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29038-1_11/.

Du:2012:PIE

[695] Fang Du, Yueguo Chen, and Xiaoyong Du. Partitioned indexes for entity search over RDF knowledge bases. Lecture Notes in Computer Science, 7238:141–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29038-1_12/.

Truong:2012:STS

[696] Ba Quan Truong, Sourav S. Bhowmick, and Curtis Dyreson. SIN-BAD: Towards structure-independent querying of common neighbors in XML databases. Lecture Notes in Computer Science, 7238:156–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29038-1_13/.

Zhou:2012:TSC

[697] Junfeng Zhou, Zhifeng Bao, Ziyang Chen, Guoxiang Lan, Xudong Lin, and Tok Wang Ling. Top-down SLCA computation based on list partition. Lecture Notes in Computer Science, 7238:172–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29038-1_14/.

Chang:2012:EIC

[698] Ya-Hui Chang, Po-Hsien Chien, and Yu-Kai Chang. Efficiently identifying contributors for XML keyword search. Lecture Notes in Computer Science, 7238:185–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29038-1_15/.

Huang:2012:SSC

[699] Xin Huang, Hong Cheng, Jiong Yang, Jeffery Xu Yu, Hongliang Fei, and Jun Huan. Semi-supervised clustering of graph objects: a subgraph mining approach. Lecture Notes in Computer Science, 7238: 197–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29038-1_16/.

Xia:2012:PLU

[700] Huan Xia, Juanzi Li, Jie Tang, and Marie-Francine Moens. Plink-LDA: Using link as prior information in topic modeling. Lecture Notes in Computer Science, 7238:213–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29038-1_17/.

Assent:2012:AAO

[701] Ira Assent, Philipp Kranen, Corinna Baldauf, and Thomas Seidl. AnyOut: Anytime outlier detection on streaming data. Lecture Notes in Computer Science, 7238:228–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29038-1_18/.

Nguyen:2012:EBP

[702] Minh Nhut Nguyen, Xiao-Li Li, and See-Kiong Ng. Ensemble based positive unlabeled learning for time series classification. Lecture Notes in Computer Science, 7238:243–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29038-1_19/.

Rashid:2012:EMR

[703] Md. Mamunur Rashid, Md. Rezaul Karim, Byeong-Soo Jeong, and Ho-Jin Choi. Efficient mining regularly frequent patterns in transactional databases. Lecture Notes in Computer Science, 7238:258–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29038-1_20/.

Anonymous:2012:FMbq

[704] Anonymous. Front matter. Lecture Notes in Computer Science, 7238: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29038-1/1.

Le:2012:TBP

[705] Trieu Minh Nhut Le and Jinli Cao. Top-k best probability queries on probabilistic data. Lecture Notes in Computer Science, 7239:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29035-0_1/.

Bai:2012:PRS

[706] Mei Bai, Junchang Xin, and Guoren Wang. Probabilistic reverse skyline query processing over uncertain data stream. Lecture Notes in Computer Science, 7239:17–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29035-0_2/.

Lofi:2012:MAS

[707] Christoph Lofi, Ulrich Güntzer, and Wolf-Tilo Balke. Malleability-aware skyline computation on linked open data. Lecture Notes in Computer Science, 7239:33–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29035-0_3/.

Yap:2012:ENI

[708] Ghim-Eng Yap, Xiao-Li Li, and Philip S. Yu. Effective next-items recommendation via personalized sequential pattern mining. Lecture Notes in Computer Science, 7239:48–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29035-0_4/.

Xu:2012:STK

[709] Yanwei Xu, Jihong Guan, and Yoshiharu Ishikawa. Scalable top- k keyword search in relational databases. Lecture Notes in Computer Science, 7239:65–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29035-0_5/.

Roocks:2012:CEE

[710] Patrick Roocks, Markus Endres, Stefan Mandl, and Werner Kießling. Composition and efficient evaluation of context-aware preference queries. Lecture Notes in Computer Science, 7239:81–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29035-0_6/.

Sun:2012:ABI

[711] Weiwei Sun, Peng Liu, Jingjing Wu, Yongrui Qin, and Baihua Zheng. An automaton-based index scheme for on-demand XML data broadcast. Lecture Notes in Computer Science, 7239:96–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29035-0_7/.

Bellam:2012:CRS

[712] Haritha Bellam, Saladi Rahul, and Krishnan Rajan. Colored range searching on internal memory. Lecture Notes in Computer Science, 7239: 111–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29035-0_8/.

Liu:2012:CFQ

[713] Weimo Liu, Weiwei Sun, Chunan Chen, Yan Huang, Yinan Jing, and Kunjie Chen. Circle of friend query in geo-social networks. Lecture Notes in Computer Science, 7239:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29035-0_9/.

Hikida:2012:PSS

[714] Satoshi Hikida, Hieu Hanh Le, and Haruo Yokota. A power saving storage method that considers individual disk rotation. Lecture Notes in Computer Science, 7239:138–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29035-0_10/.

Ding:2012:CIM

[715] Linlin Ding, Junchang Xin, Guoren Wang, and Shan Huang. ComMapReduce: An improvement of MapReduce with lightweight communication mechanisms. Lecture Notes in Computer Science, 7239:150–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29035-0_11/.

Shi:2012:HCE

[716] Yingjie Shi, Xiaofeng Meng, and Bingbing Liu. Halt or continue: Estimating progress of queries in the cloud. Lecture Notes in Computer Science, 7239:169–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29035-0_12/.

Eavis:2012:TSP

[717] Todd Eavis and Ahmad Taleb. Towards a scalable, performance-oriented OLAP storage engine. Lecture Notes in Computer Science, 7239:185– 202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29035-0_13/.

Nishii:2012:HSS

[718] Shunsuke Nishii and Toyotaro Suzumura. Highly scalable speech processing on data stream management system. Lecture Notes in Computer Science, 7239:203–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29035-0_14/.

Tago:2012:EFS

[719] Shinichiro Tago, Tatsuya Asai, Takashi Katoh, Hiroaki Morikawa, and Hiroya Inakoshi. EVIS: a fast and scalable episode matching engine for massively parallel data streams. Lecture Notes in Computer Science, 7239: 213–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29035-0_15/.

Hong:2012:RTA

[720] Seokjin Hong, Rana Prasad Sahu, M. R. Srikanth, Supriya Mandal, and Kyoung-Gu Woo. Real-time analysis of ECG data using mobile data stream management system. Lecture Notes in Computer Science, 7239: 224–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29035-0_16/.

Han:2012:MMB

[721] Hyeok Han and Seong il Jin. A main memory based spatial DBMS: Kairos. Lecture Notes in Computer Science, 7239:234–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29035-0_17/.

Lim:2012:SIS

[722] Sungjoon Lim, Taesul Seo, Changhan Lee, and Soungsoo Shin. Study on the international standardization for the semantic metadata mapping procedure. Lecture Notes in Computer Science, 7239:243–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29035-0_18/.

Thollot:2012:SUS

[723] Raphaël Thollot, Nicolas Kuchmann-Beauger, and Marie-Aude Aufaure. Semantics and usage statistics for multi-dimensional query expansion. Lecture Notes in Computer Science, 7239:250–260, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29035-0_19/.

Feng:2012:HBU

[724] Haitang Feng, Nicolas Lumineau, Mohand-Saíd Hacid, and Richard Domps. Hierarchy-based update propagation in decision support systems. Lecture Notes in Computer Science, 7239:261–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29035-0_20/.

Anonymous:2012:FMbr

[725] Anonymous. Front matter. Lecture Notes in Computer Science, 7239: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29035-0/1.

Kim:2012:CSP

[726] Bumsoo Kim. Commercial SSD products— status quo and next. Lecture Notes in Computer Science, 7240:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-29023-7_1.

Ou:2012:IDP

[727] Yi Ou and Theo Härder. Improving database performance using a flash-based write cache. Lecture Notes in Computer Science, 7240:2– 13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29023-7_2/.

Wang:2012:BAB

[728] Rui Wang, Lihua Yue, Peiquan Jin, and Junjie Wang. h-Buffer: An adaptive buffer management scheme for flash-based storage devices. Lecture Notes in Computer Science, 7240:14–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29023-7_3/.

Fan:2012:SSR

[729] Yulei Fan, Wei Cao, and Xiaofeng Meng. A study of space reclamation on flash-based append-only storage management. Lecture Notes in Computer Science, 7240:28–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29023-7_4/.

Wang:2012:DGF

[730] Junjie Wang, Lihua Yue, Peiquan Jin, and Rui Wang. A dual-grained FTL for flash memory. Lecture Notes in Computer Science, 7240:40– 52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29023-7_5/.

Gudmundsson:2012:IST

[731] Gylfi Thór Gudmundsson, Laurent Amsaleg, and Björn Thór Jónsson. Impact of storage technology on the efficiency of cluster-based highdimensional index creation. Lecture Notes in Computer Science, 7240: 53–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29023-7_6/.

Pawlik:2012:IAT

[732] Maciej Pawlik and Wojciech Macyna. Implementation of the aggregated R-tree over flash memory. Lecture Notes in Computer Science, 7240: 65–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29023-7_7/.

Cao:2012:FBD

[733] Qingling Cao, Zhichao Liang, Yulei Fan, and Xiaofeng Meng. A flashbased decomposition storage model. Lecture Notes in Computer Science, 7240:73–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29023-7_8/.

Chaze:2012:CBN

[734] Xavier Chaze, Amal Bouejla, Aldo Napoli, Franck Guarnieri, and Thibaut Eude. The contribution of Bayesian networks to manage risks of maritime piracy against oil offshore fields. Lecture Notes in Computer Science, 7240:81–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29023-7_9/.

Kim:2012:SOB

[735] Gihong Kim, Bonghee Hong, and Joonho Kwon. A scalable object based discovery service for global tracing of RFID products. Lecture Notes in Computer Science, 7240:92–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29023-7_10/.

Xu:2012:SAC

[736] Guandong Xu and Zongda Wu. On smart and accurate contextual advertising. Lecture Notes in Computer Science, 7240:104, 2012.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-29023-7_11.

Park:2012:RSB

[737] Kyung-Mi Park, Hogun Park, Hyoung-Gon Kim, and Heedong Ko. Review summarization based on linguistic knowledge. Lecture Notes in Computer Science, 7240:105–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29023-7_12/.

Li:2012:FRM

[738] Lin Li, Huifan Xiao, and Guandong Xu. Finding related micro-blogs based on WordNet. Lecture Notes in Computer Science, 7240:115–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29023-7_13/.

Park:2012:EPN

[739] Yonghun Park, Kyoungsoo Bok, and Jaesoo Yoo. An efficient path nearest neighbor query processing scheme for location based services. Lecture Notes in Computer Science, 7240:123–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29023-7_14/.

Jeong:2012:SCB

[740] Ok-Ran Jeong and Jehwan Oh. Social community based blog search framework. Lecture Notes in Computer Science, 7240:130–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29023-7_15/.

Chen:2012:LGT

[741] Xian Chen, Hyoseop Shin, and Minsoo Lee. LSA as ground truth for recommending “Flickr-aware” representative tags. Lecture Notes in Computer Science, 7240:142–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29023-7_16/.

Cuppens:2012:AAC

[742] Frédéric Cuppens, Nora Cuppens-Boulahia, and Eduardo Pena Viña. Adaptive access control enforcement in social network using aspect weaving. Lecture Notes in Computer Science, 7240:154–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29023-7_17/.

Wakamiya:2012:ERU

[743] Shoko Wakamiya, Ryong Lee, and Kazutoshi Sumiya. Exploring reflection of urban society through cyber-physical crowd behavior on locationbased social network. Lecture Notes in Computer Science, 7240:168– 179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29023-7_18/.

Leung:2012:MSN

[744] Carson Kai-Sang Leung and Syed K. Tanbeer. Mining social networks for significant friend groups. Lecture Notes in Computer Science, 7240: 180–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29023-7_19/.

Arora:2012:RSP

[745] Nidhi R. Arora, Wookey Lee, and Simon Soon-Hyoung Park. Ranking structural parameters for social networks. Lecture Notes in Computer Science, 7240:193–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29023-7_20/.

Anonymous:2012:FMbs

[746] Anonymous. Front matter. Lecture Notes in Computer Science, 7240: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29023-7/1.

Jeuring:2012:PTH

[747] Johan Jeuring, Alex Gerdes, and Bastiaan Heeren. A programming tutor for Haskell. Lecture Notes in Computer Science, 7241:1–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32096-5_1/.

Plasmeijer:2012:DMU

[748] Rinus Plasmeijer, Peter Achten, and Bas Lijnse. Defining multi-user Web applications with iTasks. Lecture Notes in Computer Science, 7241: 46–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32096-5_2/.

Butterfield:2012:RAF

[749] Andrew Butterfield. Reasoning about I/O in functional programs. Lecture Notes in Computer Science, 7241:93–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32096-5_3/.

Loogen:2012:EPF

[750] Rita Loogen. Eden — parallel functional programming with Haskell. Lecture Notes in Computer Science, 7241:142–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32096-5_4/.

Grelck:2012:SAC

[751] Clemens Grelck. Single Assignment C (SAC) high productivity meets high performance. Lecture Notes in Computer Science, 7241:207–278, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32096-5_5/.

Michaelson:2012:RAM

[752] Greg Michaelson and Gudmund Grov. Reasoning about multi-process systems with the box calculus. Lecture Notes in Computer Science, 7241: 279–338, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32096-5_6/.

Marlow:2012:PCP

[753] Simon Marlow. Parallel and concurrent programming in Haskell. Lecture Notes in Computer Science, 7241:339–401, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32096-5_7/.

Axelsson:2012:FAI

[754] Emil Axelsson and Mary Sheeran. Feldspar: Application and implementation. Lecture Notes in Computer Science, 7241:402–439, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32096-5_8/.

Toth:2012:SAC

[755] Melinda Tóth and István Bozó. Static analysis of complex software systems implemented in Erlang. Lecture Notes in Computer Science, 7241: 440–498, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32096-5_9/.

Pali:2012:ELL

[756] Gábor Páli. Extending little languages into big systems. Lecture Notes in Computer Science, 7241:499–516, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32096-5_10/.

Slodicak:2012:SNA

[757] Viliam Slodičák and Pavol Macko. Some new approaches in functional programming based on categories. Lecture Notes in Computer Science, 7241:517–532, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32096-5_11/.

Anonymous:2012:BMj

[758] Anonymous. Back matter. Lecture Notes in Computer Science, 7241: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32096-5/1.

Anonymous:2012:FMbt

[759] Anonymous. Front matter. Lecture Notes in Computer Science, 7241: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32096-5/1.

Branco:2012:VAH

[760] António Branco and Filipe Nunes. Verb analysis in a highly inflective language with an MFF algorithm. Lecture Notes in Computer Science, 7243:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28885-2_1/.

Vasilevski:2012:AAP

[761] Vera Vasilévski, Leonor Scliar-Cabral, and Márcio José Araújo. Automatic analysis of Portuguese verb morphology. Lecture Notes in Computer Science, 7243:12–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28885-2_2/.

Baptista:2012:CME

[762] Jorge Baptista, Lucas Nunes Vieira, Cláudio Diniz, and Nuno Mamede. Coordination of-mente ending adverbs in Portuguese: an integrated solution. Lecture Notes in Computer Science, 7243:24–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28885-2_3/.

Reis:2012:MAL

[763] Raquel Reis and António Teixeira. Morphosyntactic analysis of language in children with autism spectrum disorder. Lecture Notes in Computer Science, 7243:35–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28885-2_4/.

Ferreira:2012:LEU

[764] José Pedro Ferreira, António Lourinho, and Margarita Correia. Lince, an end user tool for the implementation of the spelling reform of Portuguese. Lecture Notes in Computer Science, 7243:46–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28885-2_5/.

Carvalho:2012:SMC

[765] Gracinda Carvalho, Isabel Falé, David Martins de Matos, and Vitor Rocio. Searching a mixed corpus in the light of the new Portuguese orthographic norm. Lecture Notes in Computer Science, 7243:56–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28885-2_6/.

Gamallo:2012:EBC

[766] Pablo Gamallo and Marcos Garcia. Extraction of bilingual cognates from Wikipedia. Lecture Notes in Computer Science, 7243:63–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28885-2_7/.

Zeller:2012:CBA

[767] Britta D. Zeller and Sebastian Padó. Corpus-based acquisition of support verb constructions for Portuguese. Lecture Notes in Computer Science, 7243:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28885-2_8/.

Lopes:2012:IPT

[768] Lucelene Lopes and Renata Vieira. Improving Portuguese term extraction. Lecture Notes in Computer Science, 7243:85–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28885-2_9/.

Xavier:2012:MAE

[769] Clarissa Castellã Xavier and Vera Lúcia Strube de Lima. A method for automatically extracting domain semantic networks from Wikipedia. Lecture Notes in Computer Science, 7243:93–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28885-2_10/.

Costa:2012:ETI

[770] Francisco Costa and António Branco. Extracting temporal information from Portuguese texts. Lecture Notes in Computer Science, 7243: 99–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28885-2_11/.

Craveiro:2012:ITP

[771] Olga Craveiro, Joaquim Macedo, and Henrique Madeira. It is the time for Portuguese texts! Lecture Notes in Computer Science, 7243: 106–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-28885-2_12/.

Genereux:2012:LPC

[772] Michel Généreux, Iris Hendrickx, and Amália Mendes. A large Portuguese corpus on-line: Cleaning and preprocessing. Lecture Notes in Computer Science, 7243:113–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28885-2_13/.

Simoes:2012:DAS

[773] Alberto Simões, Álvaro Iriarte Sanromán, and José Joáo Almeida. Dicionário–Aberto: a source of resources for the Portuguese language processing. Lecture Notes in Computer Science, 7243:121–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28885-2_14/.

Pinheiro:2012:TCS

[774] Vládia Pinheiro, Vasco Furtado, Tarcisio Pequeno, and Caio Ferreira. Towards a common sense base in Portuguese for the linked open data cloud. Lecture Notes in Computer Science, 7243:128–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28885-2_15/.

Luis:2012:WOP

[775] Ana R. Luís. Weak object pronouns in Brazilian Portuguese: An LFG analysis. Lecture Notes in Computer Science, 7243:139–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28885-2_16/.

Fernandes:2012:EGF

[776] Eraldo R. Fernandes and Ruy L. Milidiú. Entropy-guided feature generation for structured learning of Portuguese dependency parsing. Lecture Notes in Computer Science, 7243:146–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28885-2_17/.

Beck:2012:BIS

[777] Daniel Emilio Beck and Helena de Medeiros Caseli. Bayesian induction of syntactic language models for Brazilian Portuguese. Lecture Notes in Computer Science, 7243:157–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28885-2_18/.

Correia:2012:AGC

[778] Rui Correia, Jorge Baptista, Maxine Eskenazi, and Nuno Mamede. Automatic generation of cloze question stems. Lecture Notes in Computer Science, 7243:168–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28885-2_19/.

Batista:2012:TDU

[779] David S. Batista, João D. Ferreira, Francisco M. Couto, and Mário J. Silva. Toponym disambiguation using ontology-based semantic similarity. Lecture Notes in Computer Science, 7243:179–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-28885-2_20/.

Anonymous:2012:FMbu

[780] Anonymous. Front matter. Lecture Notes in Computer Science, 7243: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28885-2/1.

Castle:2012:EHL

[781] Tom Castle and Colin G. Johnson. Evolving high-level imperative program trees with strongly formed genetic programming. Lecture Notes in Computer Science, 7244:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29139-5_1/.

Cotillon:2012:AGP

[782] Alban Cotillon, Philip Valencia, and Raja Jurdak. Android genetic programming framework. Lecture Notes in Computer Science, 7244: 13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29139-5_2/.

Dracopoulos:2012:GPG

[783] Dimitris C. Dracopoulos and Dimitrios Effraimidis. Genetic programming for generalised helicopter hovering control. Lecture Notes in Computer Science, 7244:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29139-5_3/.

Howard:2012:CGP

[784] Gerard David Howard, Larry Bull, and Andrew Adamatzky. Cartesian genetic programming for memristive logic circuits. Lecture Notes in Computer Science, 7244:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29139-5_4/.

Jackson:2012:NNF

[785] David Jackson. A new, node-focused model for genetic programming. Lecture Notes in Computer Science, 7244:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29139-5_5/.

Krawiec:2012:MCG

[786] Krzysztof Krawiec. Medial crossovers for genetic programming. Lecture Notes in Computer Science, 7244:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29139-5_6/.

Machado:2012:IFD

[787] Penousal Machado, João Correia, and Juan Romero. Improving face detection. Lecture Notes in Computer Science, 7244:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29139-5_7/.

Murphy:2012:GBI

[788] Eoin Murphy, Erik Hemberg, Miguel Nicolau, Michael O’Neill, and Anthony Brabazon. Grammar bias and initialisation in grammar based genetic programming. Lecture Notes in Computer Science, 7244:85– 96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29139-5_8/.

Neshatian:2012:IRM

[789] Kourosh Neshatian and Mengjie Zhang. Improving relevance measures using genetic programming. Lecture Notes in Computer Science, 7244: 97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29139-5_9/.

Nguyen:2012:IFS

[790] Quang Uy Nguyen, Xuan Hoai Nguyen, Michael O’Neill, and Alexandros Agapitos. An investigation of fitness sharing with semantic and syntactic distance metrics. Lecture Notes in Computer Science, 7244: 109–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29139-5_10/.

Nguyen:2012:ERO

[791] Su Nguyen, Mengjie Zhang, Mark Johnston, and Kay Chen Tan. Evolving reusable operation-based due-date assignment models for job shop scheduling with genetic programming. Lecture Notes in Computer Science, 7244:121–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29139-5_11/.

Nicolau:2012:EIM

[792] Miguel Nicolau, Matthew Saunders, Michael O’Neill, Bruce Osborne, and Anthony Brabazon. Evolving interpolating models of net ecosystem CO_{2} exchange using grammatical evolution. Lecture Notes in Computer Science, 7244:134–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29139-5_12/.

Olmo:2012:MOA

[793] Juan Luis Olmo, José Raúl Romero, and Sebastián Ventura. Multiobjective ant programming for mining classification rules. Lecture Notes in Computer Science, 7244:146–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29139-5_13/.

Parkes:2012:MAG

[794] Andrew J. Parkes, Ender Özcan, and Matthew R. Hyde. Matrix analysis of genetic programming mutation. Lecture Notes in Computer Science, 7244:158–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29139-5_14/.

Seaton:2012:EAM

[795] Tom Seaton, Julian F. Miller, and Tim Clarke. An ecological approach to measuring locality in linear genotype to phenotype maps. Lecture Notes in Computer Science, 7244:170–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29139-5_15/.

Sikulova:2012:CCG

[796] Michaela Šikulová and Lukáš Sekanina. Coevolution in Cartesian genetic programming. Lecture Notes in Computer Science, 7244:182–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29139-5_16/.

Smolka:2012:EDM

[797] Tobiáš Smolka, Petr Švenda, Lukáš Sekanina, and Vashek Matyáš. Evolutionary design of message efficient secrecy amplification protocols. Lecture Notes in Computer Science, 7244:194–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29139-5_17/.

Tavares:2012:ADA

[798] Jorge Tavares and Francisco B. Pereira. Automatic design of ant algorithms with grammatical evolution. Lecture Notes in Computer Science, 7244:206–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29139-5_18/.

Goncalves:2012:RST

[799] Ivo Gonçalves, Sara Silva, Joana B. Melo, and João M. B. Carreiras. Random sampling technique for overfitting control in genetic programming. Lecture Notes in Computer Science, 7244:218–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29139-5_19/.

Kim:2012:EOS

[800] MinHyeok Kim, Robert Ian (Bob) McKay, Dong-Kyun Kim, and Xuan Hoai Nguyen. Evolutionary operator self-adaptation with diverse operators. Lecture Notes in Computer Science, 7244:230–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29139-5_20/.

Anonymous:2012:FMbv

[801] Anonymous. Front matter. Lecture Notes in Computer Science, 7244: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29139-5/1.

Dominguez:2012:MCE

[802] Julián Domínguez and Enrique Alba. A methodology for comparing the execution time of metaheuristics running on different hardware. Lecture Notes in Computer Science, 7245:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29124-1_1/.

Schwengerer:2012:VNS

[803] Martin Schwengerer and Sandro Pirkwieser. A variable neighborhood search approach for the two-echelon location-routing problem. Lecture Notes in Computer Science, 7245:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29124-1_2/.

Avila:2012:IBM

[804] Thais Ávila, Ángel Corberán, and Isaac Plana. An ILS-based metaheuristic for the stacker crane problem. Lecture Notes in Computer Science, 7245:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29124-1_3/.

Jemai:2012:NIA

[805] Jaber Jemai, Manel Zekri, and Khaled Mellouli. An NSGA-II algorithm for the green vehicle routing problem. Lecture Notes in Computer Science, 7245:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29124-1_4/.

deOliveira:2012:CSH

[806] Rudinei Martins de Oliveira. Clustering search heuristic for solving a continuous berth allocation problem. Lecture Notes in Computer Science, 7245:49–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29124-1_5/.

Holborn:2012:CHE

[807] Penny L. Holborn and Jonathan M. Thompson. Combining heuristic and exact methods to solve the vehicle routing problem with pickups, deliveries and time Windows. Lecture Notes in Computer Science, 7245: 63–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29124-1_6/.

AlMoubayed:2012:MMO

[808] Noura Al Moubayed and Andrei Petrovski. D^{2} MOPSO: Multi-objective particle swarm optimizer based on decomposition and dominance. Lecture Notes in Computer Science, 7245:75–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29124-1_7/.

Rahmani:2012:DRU

[809] Mohsen Rahmani and Ruben A. Romero. Domain reduction using GRASP construction phase for transmission expansion planning problem. Lecture Notes in Computer Science, 7245:87–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29124-1_8/.

Allerding:2012:ELM

[810] Florian Allerding and Marc Premm. Electrical load management in smart homes using evolutionary algorithms. Lecture Notes in Computer Science, 7245:99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29124-1_9/.

Chicano:2012:ECF

[811] Francisco Chicano and Enrique Alba. Exact computation of the fitnessdistance correlation for pseudoboolean functions with one global optimum. Lecture Notes in Computer Science, 7245:111–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29124-1_10/.

Gavanelli:2012:GAS

[812] Marco Gavanelli and Maddalena Nonato. Genetic algorithms for scheduling devices operation in a water distribution system in response to contamination events. Lecture Notes in Computer Science, 7245:124–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29124-1_11/.

Ochoa:2012:HBF

[813] Gabriela Ochoa, Matthew Hyde, and Tim Curtois. HyFlex: a benchmark framework for cross-domain heuristic search. Lecture Notes in Computer Science, 7245:136–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29124-1_12/.

Kubalik:2012:HHB

[814] Jiří Kubalík. Hyper-heuristic based on iterated local search driven by evolutionary algorithm. Lecture Notes in Computer Science, 7245: 148–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29124-1_13/.

Loudni:2012:IDD

[815] Samir Loudni. Intensification/diversification-driven ILS for a graph coloring problem. Lecture Notes in Computer Science, 7245:160–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29124-1_14/.

Rodriguez:2012:IGA

[816] Francisco J. Rodriguez and Christian Blum. Iterated greedy algorithms for the maximal covering location problem. Lecture Notes in Computer Science, 7245:172–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29124-1_15/.

Garza-Fabre:2012:MHM

[817] Mario Garza-Fabre and Eduardo Rodriguez-Tello. Multiobjectivizing the HP model for protein structure prediction. Lecture Notes in Computer Science, 7245:182–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29124-1_16/.

Abdou:2012:MPR

[818] Wahabou Abdou and Christelle Bloch. Multi-pareto-ranking evolutionary algorithm. Lecture Notes in Computer Science, 7245:194–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29124-1_17/.

Dubois-Lacoste:2012:PLS

[819] Jérémie Dubois-Lacoste and Manuel López-Ibáñez. Pareto local search algorithms for anytime bi-objective optimization. Lecture Notes in Computer Science, 7245:206–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29124-1_18/.

He:2012:PSM

[820] Jun He, Feidun He, and Hongbin Dong. Pure strategy or mixed strategy? Lecture Notes in Computer Science, 7245:218–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29124-1_19/.

Fakeih:2012:RGA

[821] Adnan Fakeih and Ahmed Kattan. Recurrent genetic algorithms: Sustaining evolvability. Lecture Notes in Computer Science, 7245:230–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29124-1_20/.

Anonymous:2012:FMbw

[822] Anonymous. Front matter. Lecture Notes in Computer Science, 7245: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29124-1/1.

Manning:2012:ATD

[823] Timmy Manning and Paul Walsh. Automatic task decomposition for the NeuroEvolution of Augmenting Topologies (NEAT) algorithm. Lecture Notes in Computer Science, 7246:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29066-4_1/.

Manzoni:2012:ERS

[824] Luca Manzoni, Mauro Castelli, and Leonardo Vanneschi. Evolutionary reaction systems. Lecture Notes in Computer Science, 7246:13–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29066-4_2/.

Rosenbaum:2012:OEW

[825] Lars Rosenbaum, Andreas Jahn, and Andreas Zell. Optimizing the edge weights in optimal assignment methods for virtual screening with particle swarm optimization. Lecture Notes in Computer Science, 7246: 26–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29066-4_3/.

Darabos:2012:LFG

[826] Christian Darabos, Mario Giacobini, Ting Hu, and Jason H. Moore. Lévy– Flight genetic programming: Towards a new mutation paradigm. Lecture Notes in Computer Science, 7246:38–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29066-4_4/.

Marini:2012:UZL

[827] Simone Marini and Alessandra Conversi. Understanding zooplankton long term variability through genetic programming. Lecture Notes in Computer Science, 7246:50–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29066-4_5/.

Franken:2012:IDR

[828] Holger Franken, Alexander Seitz, Rainer Lehmann, Hans-Ulrich Häring, and Norbert Stefan. Inferring disease-related metabolite dependencies with a Bayesian optimization algorithm. Lecture Notes in Computer Science, 7246:62–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29066-4_6/.

Nobile:2012:GBM

[829] Marco S. Nobile, Daniela Besozzi, Paolo Cazzaniga, Giancarlo Mauri, and Dario Pescini. A GPU-based multi-swarm PSO method for parameter estimation in stochastic biological systems exploiting discrete-time target series. Lecture Notes in Computer Science, 7246:74–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29066-4_7/.

Pinheiro:2012:TEC

[830] Flávio L. Pinheiro, Francisco C. Santos, and Jorge M. Pacheco. Tracking the evolution of cooperation in complex networked populations. Lecture Notes in Computer Science, 7246:86–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29066-4_8/.

Vanneschi:2012:GGB

[831] Leonardo Vanneschi, Matteo Mondini, Martino Bertoni, Alberto Ronchi, and Mattia Stefano. GeNet: a graph-based genetic programming framework for the reverse engineering of gene regulatory networks. Lecture Notes in Computer Science, 7246:97–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29066-4_9/.

Gonzalez-Alvarez:2012:CMA

[832] David L. González-Álvarez, Miguel A. Vega-Rodríguez, and Juan A. Gómez-Pulido. Comparing multiobjective artificial bee colony adaptations for discovering DNA motifs. Lecture Notes in Computer Science, 7246:110–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29066-4_10/.

Pan:2012:RMW

[833] Qinxin Pan, Christian Darabos, and Jason H. Moore. The role of mutations in whole genome duplication. Lecture Notes in Computer Science, 7246:122–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29066-4_11/.

Holzinger:2012:CMM

[834] Emily R. Holzinger, Scott M. Dudek, Alex T. Frase, Brooke Fridley, and Prabhakar Chalise. Comparison of methods for meta-dimensional data analysis using in silico and biological data sets. Lecture Notes in Computer Science, 7246:134–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29066-4_12/.

Santander-Jimenez:2012:IPT

[835] Sergio Santander-Jiménez, Miguel A. Vega-Rodríguez, and Juan A. Gómez-Pulido. Inferring phylogenetic trees using a multiobjective artificial bee colony algorithm. Lecture Notes in Computer Science, 7246: 144–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29066-4_13/.

Asencio-Cortes:2012:PMM

[836] Gualberto Asencio-Cortés, Jesús S. Aguilar-Ruiz, and Alfonso E. Márquez-Chamorro. Prediction of mitochondrial matrix protein structures based on feature selection and fragment assembly. Lecture Notes in Computer Science, 7246:156–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29066-4_14/.

Kancherla:2012:FSL

[837] Kesav Kancherla and Srinivas Mukkamala. Feature selection for lung cancer detection using SVM based recursive feature elimination method. Lecture Notes in Computer Science, 7246:168–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29066-4_15/.

Spirov:2012:MGE

[838] Alexander V. Spirov, Nina E. Golyandina, David M. Holloway, and Theodore Alexandrov. Measuring gene expression noise in early Drosophila embryos: The highly dynamic compartmentalized microenvironment of the blastoderm is one of the main sources of noise. Lecture Notes in Computer Science, 7246:177–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29066-4_16/.

Granizo-Mackenzie:2012:AIS

[839] Delaney Granizo-Mackenzie and Jason H. Moore. Artificial immune systems perform valuable work when detecting epistasis in human genetic datasets. Lecture Notes in Computer Science, 7246:189–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29066-4_17/.

Buchanan:2012:BIM

[840] Carrie C. Buchanan, John R. Wallace, Alex T. Frase, and Eric S. Torstenson. A biologically informed method for detecting associations with rare variants. Lecture Notes in Computer Science, 7246:201–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29066-4_18/.

Pizzuti:2012:CDP

[841] Clara Pizzuti, Simona E. Rombo, and Elena Marchiori. Complex detection in protein-protein interaction networks: a compact overview for researchers and practitioners. Lecture Notes in Computer Science, 7246: 211–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29066-4_19/.

Santiesteban-Toca:2012:SRI

[842] Cosme E. Santiesteban-Toca, Gualberto Asencio-Cortés, and Alfonso E. Márquez-Chamorro. Short-range interactions and decision tree-based protein contact map predictor. Lecture Notes in Computer Science, 7246: 224–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29066-4_20/.

Anonymous:2012:FMbx

[843] Anonymous. Front matter. Lecture Notes in Computer Science, 7246: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29066-4/1.

Antunes:2012:GCA

[844] Rui Filipe Antunes and Frederic Fol Leymarie. Generative choreography: Animating in real-time dancing avatars. Lecture Notes in Computer Science, 7247:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29142-5_1/.

Bergen:2012:AME

[845] Steve Bergen and Brian J. Ross. Aesthetic 3D model evolution. Lecture Notes in Computer Science, 7247:11–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29142-5_2/.

Byrne:2012:LSI

[846] Jonathan Byrne and Erik Hemberg. A local search interface for interactive evolutionary architectural design. Lecture Notes in Computer Science, 7247:23–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29142-5_3/.

Colton:2012:ELA

[847] Simon Colton. Evolving a library of artistic scene descriptors. Lecture Notes in Computer Science, 7247:35–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29142-5_4/.

denHeijer:2012:EPA

[848] E. den Heijer and A. E. Eiben. Evolving pop art using scalable vector graphics. Lecture Notes in Computer Science, 7247:48–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29142-5_5/.

denHeijer:2012:MPD

[849] E. den Heijer and A. E. Eiben. Maintaining population diversity in evolutionary art. Lecture Notes in Computer Science, 7247:60–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29142-5_6/.

Eigenfeldt:2012:PPC

[850] Arne Eigenfeldt and Philippe Pasquier. Populations of populations: Composing with multiple evolutionary algorithms. Lecture Notes in Computer Science, 7247:72–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29142-5_7/.

Fernandez:2012:FRB

[851] Francisco Fernández and Francisco Chávez. Fuzzy rule based system ensemble for music genre classification. Lecture Notes in Computer Science, 7247:84–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29142-5_8/.

Fornari:2012:CEE

[852] José Fornari. A computational environment for the evolutionary sound synthesis of birdsongs. Lecture Notes in Computer Science, 7247:96– 107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29142-5_9/.

Greenfield:2012:PEC

[853] Gary Greenfield. A platform for evolving controllers for simulated drawing robots. Lecture Notes in Computer Science, 7247:108–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29142-5_10/.

Hornby:2012:LCU

[854] Gregory S. Hornby and Josh Bongard. Learning comparative user models for accelerating human-computer collaborative search. Lecture Notes in Computer Science, 7247:117–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29142-5_11/.

Johnson:2012:FEA

[855] Colin G. Johnson. Fitness in evolutionary art and music: What has been used and what could be used? Lecture Notes in Computer Science, 7247: 129–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29142-5_12/.

Kaliakatsos–Papakostas:2012:IEB

[856] Maximos A. Kaliakatsos-Papakostas. Interactive evolution of 8-bit melodies with genetic programming towards finding aesthetic measures for sound. Lecture Notes in Computer Science, 7247:141–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29142-5_13/.

Li:2012:IAF

[857] Yang Li, Changjun Hu, Ming Chen, and Jingyuan Hu. Investigating aesthetic features to model human preference in evolutionary art. Lecture Notes in Computer Science, 7247:153–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29142-5_14/.

Lopes:2012:TP

[858] Phil Lopes and Paulo Urbano. The traveling percussionist. Lecture Notes in Computer Science, 7247:165–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29142-5_15/.

Loughran:2012:GPM

[859] Róisín Loughran and Jacqueline Walker. Genetic programming for musical sound analysis. Lecture Notes in Computer Science, 7247:176– 186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29142-5_16/.

Machado:2012:EBE

[860] Penousal Machado, João Correia, and Juan Romero. Expression-based evolution of faces. Lecture Notes in Computer Science, 7247:187–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29142-5_17/.

McDermott:2012:GGR

[861] James McDermott. Graph grammars as a representation for interactive evolutionary 3D design. Lecture Notes in Computer Science, 7247: 199–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29142-5_18/.

Anonymous:2012:PTE

[862] Anonymous. Polyphonic transcription: Exploring a hybrid of tone models and particle swarm optimisation. Lecture Notes in Computer Science, 7247:211–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29142-5_19/.

Smith:2012:RLC

[863] Benjamin D. Smith and Guy E. Garnett. Reinforcement learning and the creative, automated music improviser. Lecture Notes in Computer Science, 7247:223–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29142-5_20/.

Anonymous:2012:BMk

[864] Anonymous. Back matter. Lecture Notes in Computer Science, 7247: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29142-5/1.

Anonymous:2012:FMby

[865] Anonymous. Front matter. Lecture Notes in Computer Science, 7247: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29142-5/1.

Lanza-Gutierrez:2012:OEC

[866] José Manuel Lanza-Gutiérrez, Juan Antonio Gómez-Pulido, and Miguel A. Vega-Rodríguez. Optimizing energy consumption in heterogeneous wireless sensor networks by means of evolutionary algorithms. Lecture Notes in Computer Science, 7248:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29178-4_1/.

LaRoche:2012:NPD

[867] Patrick LaRoche, A. Nur Zincir-Heywood, and Malcolm I. Heywood. Network protocol discovery and analysis via live interaction. Lecture Notes in Computer Science, 7248:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29178-4_2/.

Limmer:2012:EDA

[868] Steffen Limmer, Dietmar Fey, Ulrich Lohmann, and Jürgen Jahns. Evolutionary design of active free space optical networks based on digital mirror devices. Lecture Notes in Computer Science, 7248:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29178-4_3/.

Tabia:2012:FRO

[869] Nourredine Tabia, Alexandre Gondran, Oumaya Baala, and Alexandre Caminada. Frequency robustness optimization with respect to traffic distribution for LTE system. Lecture Notes in Computer Science, 7248: 31–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29178-4_4/.

Arsuaga-Rios:2012:SWO

[870] María Arsuaga-Ríos, Francisco Prieto-Castrillo, and Miguel A. Vega-Rodríguez. Small-world optimization applied to job scheduling on grid environments from a multi-objective perspective. Lecture Notes in Computer Science, 7248:42–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29178-4_5/.

Garcia-Sanchez:2012:TDE

[871] Pablo García-Sánchez, A. E. Eiben, Evert Haasdijk, and Berend Weel. Testing diversity-enhancing migration policies for hybrid on-line evolution of robot controllers. Lecture Notes in Computer Science, 7248: 52–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29178-4_6/.

Kuyucu:2012:EOP

[872] Tüze Kuyucu, Ivan Tanev, and Katsunori Shimohara. Evolutionary optimization of pheromone-based stigmergic communication. Lecture Notes in Computer Science, 7248:63–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29178-4_7/.

Pacula:2012:HTB

[873] Maciej Pacula, Jason Ansel, Saman Amarasinghe, and Una-May O’Reilly. Hyperparameter tuning in bandit-based adaptive operator selection. Lecture Notes in Computer Science, 7248:73–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29178-4_8/.

Richter:2012:ADF

[874] Hendrik Richter. Analyzing dynamic fitness landscapes of the targeting problem of chaotic systems. Lecture Notes in Computer Science, 7248:83–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29178-4_9/.

Trueba:2012:SOS

[875] Pedro Trueba, Abraham Prieto, Francisco Bellas, Pilar Caamaño, and Richard J. Duro. Self-organization and specialization in multiagent systems through open-ended natural evolution. Lecture Notes in Computer Science, 7248:93–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29178-4_10/.

Turkey:2012:ETA

[876] Mikdam Turkey and Riccardo Poli. An empirical tool for analysing the collective behaviour of population-based algorithms. Lecture Notes in Computer Science, 7248:103–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29178-4_11/.

Wang:2012:SPO

[877] Crown Guan Wang and Kwok Yip Szeto. Sales potential optimization on directed social networks: a quasi-parallel genetic algorithm approach. Lecture Notes in Computer Science, 7248:114–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29178-4_12/.

Weel:2012:EMC

[878] Berend Weel, Evert Haasdijk, and A. E. Eiben. The emergence of multicellular robot organisms through on-line on-board evolution. Lecture Notes in Computer Science, 7248:124–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29178-4_13/.

Agapitos:2012:ESF

[879] Alexandros Agapitos, Michael O’Neill, and Anthony Brabazon. Evolving seasonal forecasting models with genetic programming in the context of pricing weather-derivatives. Lecture Notes in Computer Science, 7248: 135–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29178-4_14/.

Arriaga:2012:SAH

[880] Jonathan Arriaga and Manuel Valenzuela-Rendón. Steepest ascent Hill climbing for portfolio selection. Lecture Notes in Computer Science, 7248: 145–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29178-4_15/.

Azzini:2012:NEA

[881] Antonia Azzini, Mauro Dragoni, and Andrea G. B. Tettamanzi. A neuroevolutionary approach to intraday financial modeling. Lecture Notes in Computer Science, 7248:155–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29178-4_16/.

Duran:2012:CSM

[882] Feijoo E. Colomine Duran, Carlos Cotta, and Antonio J. Fernández-Leiva. A comparative study of multi-objective evolutionary algorithms to optimize the selection of investment portfolios with cardinality constraints. Lecture Notes in Computer Science, 7248:165–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29178-4_17/.

Contreras:2012:GCT

[883] Iván Contreras, José Ignacio Hidalgo, and Laura Núñez-Letamendia. A GA combining technical and fundamental analysis for trading the stock market. Lecture Notes in Computer Science, 7248:174–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-29178-4_18/.

Mayo:2012:EDS

[884] Michael Mayo. Evolutionary data selection for enhancing models of intraday Forex time series. Lecture Notes in Computer Science, 7248: 184–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-29178-4_19/.

Cook:2012:IRC

[885] Michael Cook, Simon Colton, and Jeremy Gow. Initial results from cooperative co-evolution for automated platformer design. Lecture Notes in Computer Science, 7248:194–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29178-4_20/.

Anonymous:2012:FMbz

[886] Anonymous. Front matter. Lecture Notes in Computer Science, 7248: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29178-4/1.

Anastasi:2012:MTN

[887] Giuseppe Anastasi, Emilio Bellini, Elisabetta Di Nitto, Carlo Ghezzi, Letizia Tanca, and Eugenio Zimeo, editors. Methodologies and Technologies for Networked Enterprises: ArtDeco: Adaptive Infrastructures for Decentralised Organisations, volume 7200 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31738-3 (print), 3-642-31739-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31739-2.

Schmitt:2012:MME

[888] Jens B. Schmitt, editor. Measurement, Modelling, and Evaluation of Computing Systems and Dependability and Fault Tolerance: 16th International GI/ITG Conference, MMB & DFT 2012, Kaiserslautern, Germany, March 19–21, 2012. Proceedings, volume 7201 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28539-2 (print), 3-642-28540-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-28540-0.

Zhang:2012:ISI

[889] Yanning Zhang, Zhi-Hua Zhou, Changshui Zhang, and Ying Li, editors. Intelligent Science and Intelligent Data Engineering: Second Sinoforeign-interchange Workshop, IScIDE 2011, Xi’an, China, October 23– 25, 2011, Revised Selected Papers, volume 7202 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31918-1 (print), 3-642-31919-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-31919-8.

Wyrzykowski:2012:PPAa

[890] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Waśniewski, editors. Parallel Processing and Applied Mathematics: 9th International Conference, PPAM 2011, Torun, Poland, September 11–14, 2011. Revised Selected Papers, Part I, volume 7203 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-31463-5 (print), 3-642-31464-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-31464-3.

Wyrzykowski:2012:PPAb

[891] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Waśniewski, editors. Parallel Processing and Applied Mathematics: 9th International Conference, PPAM 2011, Torun, Poland, September 11–14, 2011. Revised Selected Papers, Part II, volume 7204 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-31499-6 (print), 3-642-31500-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-31500-8.

Croitoru:2012:GSK

[892] Madalina Croitoru, Sebastian Rudolph, Nic Wilson, John Howse, and Olivier Corby, editors. Graph Structures for Knowledge Representation and Reasoning: Second International Workshop, GKR 2011, Barcelona, Spain, July 16, 2011. Revised Selected Papers, volume 7205 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29448-0 (print), 3-642-29449-9 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-29449-5.

Efthimiou:2012:GSL

[893] Eleni Efthimiou, Georgios Kouroupetroglou, and Stavroula-Evita Fotinea, editors. Gesture and Sign Language in Human-Computer Interaction and Embodied Communication: 9th International Gesture Workshop, GW 2011, Athens, Greece, May 25–27, 2011, Revised Selected Papers, volume 7206 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34181-0 (print), 3-642-34182-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
34182-3.

Muggleton:2012:ILP

[894] Stephen H. Muggleton, Alireza Tamaddoni-Nezhad, and Francesca A. Lisi, editors. Inductive Logic Programming: 21st International Conference, ILP 2011, Windsor Great Park, UK, July 31–August 3, 2011, Revised Selected Papers, volume 7207 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31950-5 (print), 3-642-31951-3 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-31951-8.

Corchado:2012:HAIa

[895] Emilio Corchado, Václav Snášel, Ajith Abraham, Michał Woźniak, Manuel Graña, and Sung-Bae Cho, editors. Hybrid Artificial Intelligent Systems: 7th International Conference, HAIS 2012, Salamanca, Spain, March 28–30th, 2012. Proceedings, Part I, volume 7208 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28941-X (print), 3-642-28942-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-28942-2.

Corchado:2012:HAIb

[896] Emilio Corchado, Václav Snášel, Ajith Abraham, Michał Woźniak, Manuel Graña, and Sung-Bae Cho, editors. Hybrid Artificial Intelligent Systems: 7th International Conference, HAIS 2012, Salamanca, Spain, March 28–30th, 2012. Proceedings, Part II, volume 7209 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28930-4 (print), 3-642-28931-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-28931-6.

OBoyle:2012:CCI

[897] Michael O’Boyle, editor. Compiler Construction: 21st International Conference, CC 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24–April 1, 2012. Proceedings, volume 7210 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28651-8 (print), 3-642-28652-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-28652-0.

Seidl:2012:PLS

[898] Helmut Seidl, editor. Programming Languages and Systems: 21st European Symposium on Programming, ESOP 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7211 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28868-5 (print), 3-642-28869-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28869-2.

deLara:2012:FAS

[899] Juan de Lara and Andrea Zisman, editors. Fundamental Approaches to Software Engineering: 15th International Conference, FASE 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7212 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28871-5 (print), 3-642-28872-3 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-28872-2.

Birkedal:2012:FSS

[900] Lars Birkedal, editor. Foundations of Software Science and Computational Structures: 15th International Conference, FOSSACS 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24–April 1, 2012. Proceedings, volume 7213 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28728-X (print), 3-642-28729-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
28729-9.

Flanagan:2012:TAC

[901] Cormac Flanagan and Barbara König, editors. Tools and Algorithms for the Construction and Analysis of Systems: 18th International Conference, TACAS 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24–April 1, 2012. Proceedings, volume 7214 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28755-7 (print), 3-642-28756-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-28756-5.

Degano:2012:PST

[902] Pierpaolo Degano and Joshua D. Guttman, editors. Principles of Security and Trust: First International Conference, POST 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012, Proceedings, volume 7215 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28640-2 (print), 3-642-28641-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-28641-4.

Hadjiantonis:2012:TES

[903] Antonis M. Hadjiantonis and Burkhard Stiller, editors. Telecommunication Economics: Selected Results of the COST Action ISO605 Econ@Tel, volume 7216 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30381-1 (print), 3-642-30382-X (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-30382-1.

Dennis:2012:PMA

[904] Louise Dennis, Olivier Boissier, and Rafael H. Bordini, editors. Programming Multi-Agent Systems: 9th International Workshop, ProMAS 2011, Taipei, Taiwan, May 3, 2011, Revised Selected Papers, volume 7217 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31914-9 (print), 3-642-31915-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-31915-0.

Aloni:2012:LLM

[905] Maria Aloni, Vadim Kimmelman, Floris Roelofsen, Galit W. Sassoon, Katrin Schulz, and Matthijs Westera, editors. Logic, Language and Meaning: 18th Amsterdam Colloquium, Amsterdam, The Netherlands, December 19–21, 2011, Revised Selected Papers, volume 7218 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31481-3 (print), 3-642-31482-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31482-7.

Pan:2012:TEV

[906] Zhigeng Pan, Adrian David Cheok, Wolfgang Müller, Maiga Chang, and Mingmin Zhang, editors. Transactions on Edutainment VIII, volume 7220 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31438-4 (print), 3-642-31439-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
31439-1.

Pallis:2012:SOC

[907] George Pallis, Mohamed Jmaiel, Anis Charfi, Sven Graupner, Yücel Karabulut, Sam Guinea, Florian Rosenberg, Quan Z. Sheng, Cesare Pautasso, and Sonia Ben Mokhtar, editors. Service-Oriented Computing — ICSOC 2011 Workshops: ICSOC 2011, International Workshops WE-SOA, NFPSLAM-SOC, and Satellite Events, Paphos, Cyprus, December 5–8, 2011. Revised Selected Papers, volume 7221 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-31874-6 (print), 3-642-31875-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-31875-7.

Chen:2012:TST

[908] Liqun Chen, Moti Yung, and Liehuang Zhu, editors. Trusted Systems: Third International Conference, INTRUST 2011, Beijing, China, November 27–29, 2011, Revised Selected Papers, volume 7222 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32297-2 (print), 3-642-32298-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-32298-3.

Lones:2012:IPC

[909] Michael A. Lones, Stephen L. Smith, Sarah Teichmann, Felix Naef, James A. Walker, and Martin A. Trefzer, editors. Information Processign in Cells and Tissues: 9th International Conference, IPCAT 2012, Cambridge, UK, March 31–April 2, 2012. Proceedings, volume 7223 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28791-3 (print), 3-642-28792-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28792-3.

Baeza-Yates:2012:AIR

[910] Ricardo Baeza-Yates, Arjen P. de Vries, Hugo Zaragoza, B. Barla Cambazoglu, Vanessa Murdock, Ronny Lempel, and Fabrizio Silvestri, editors. Advances in Information Retrieval: 34th European Conference on IR Research, ECIR 2012, Barcelona, Spain, April 1–5, 2012. Proceedings, volume 7224 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28996-7 (print), 3-642-28997-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
28997-2.

Vidal:2012:LBP

[911] Germán Vidal, editor. Logic-Based Program Synthesis and Transformation: 21st International Symposium, LOPSTR 2011, Odense, Denmark, July 18–20, 2011. Revised Selected Papers, volume 7225 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32210-7 (print), 3-642-32211-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-32211-2.

Goodloe:2012:NFM

[912] Alwyn E. Goodloe and Suzette Person, editors. NASA Formal Methods: 4th International Symposium, NFM 2012, Norfolk, VA, USA, April 3– 5, 2012. Proceedings, volume 7226 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28890-1 (print), 3-642-28891-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-28891-3.

Yang:2012:SCB

[913] Shanchieh Jay Yang, Ariel M. Greenberg, and Mica Endsley, editors. Social Computing, Behavioral — Cultural Modeling and Prediction: 5th International Conference, SBP 2012, College Park, MD, USA, April 3–5, 2012. Proceedings, volume 7227 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29046-9 (print), 3-642-29047-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-29047-3.

Shi:2012:TDH

[914] Yun Q. Shi and Stefan Katzenbeisser, editors. Transactions on Data Hiding and Multimedia Security VIII: Special Issue on Pattern Recognition for IT Security, volume 7228 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31970-X (print), 3-642-31971-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31971-6.

Constable:2012:LPS

[915] Robert L. Constable and Alexandra Silva, editors. Logic and Program Semantics: Essays Dedicated to Dexter Kozen on the Occasion of His 60th Birthday, volume 7230 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29484-7 (print), 3-642-29485-5 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-29485-3.

He:2012:HIS

[916] Jing He, Xiaohui Liu, Elizabeth A. Krupinski, and Guandong Xu, editors. Health Information Science: First International Conference, HIS 2012, Beijing, China, April 8–10, 2012. Proceedings, volume 7231 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29360-3 (print), 3-642-29361-1 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-29361-0.

Ryan:2012:ISP

[917] Mark D. Ryan, Ben Smyth, and Guilin Wang, editors. Information Security Practice and Experience: 8th International Conference, ISPEC 2012, Hangzhou, China, April 9–12, 2012. Proceedings, volume 7232 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29100-7 (print), 3-642-29101-5 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-29101-2.

Wang:2012:WTA

[918] Hua Wang, Lei Zou, Guangyan Huang, Jing He, Chaoyi Pang, Hao Lan Zhang, Dongyan Zhao, and Zhuang Yi, editors. Web Technologies and Applications: APWeb 2012 International Workshops: SenDe, IDP, IEKB, MBC, Kunming, China, April 11–13, 2012. Proceedings, volume 7234 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29425-1 (print), 3-642-29426-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-29426-6.

Sheng:2012:WTA

[919] Quan Z. Sheng, Guoren Wang, Christian S. Jensen, and Guandong Xu, editors. Web Technologies and Applications: 14th Asia–Pacific Web Conference, APWeb 2012, Kunming, China, April 11–13, 2012. Proceedings, volume 7235 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29252-6 (print), 3-642-29253-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
29253-8.

DiMartino:2012:WWG

[920] Sergio Di Martino, Adriano Peron, and Taro Tezuka, editors. Web and Wireless Geographical Information Systems: 11th International Symposium, W2GIS 2012, Naples, Italy, April 12–13, 2012. Proceedings, volume 7236 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29246-1 (print), 3-642-29247-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
29247-7.

Lee:2012:DSAa

[921] Sang goo Lee, Zhiyong Peng, Xiaofang Zhou, Yang-Sae Moon, Rainer Unland, and Jaesoo Yoo, editors. Database Systems for Advanced Applications: 17th International Conference, DASFAA 2012, Busan, South Korea, April 15–19, 2012, Proceedings, Part I, volume 7238 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29037-X (print), 3-642-29038-8 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-29038-1.

Lee:2012:DSAb

[922] Sang goo Lee, Zhiyong Peng, Xiaofang Zhou, Yang-Sae Moon, Rainer Unland, and Jaesoo Yoo, editors. Database Systems for Advanced Applications: 17th International Conference, DASFAA 2012, Busan, South Korea, April 15–19, 2012, Proceedings, Part II, volume 7239 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29034-5 (print), 3-642-29035-3 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-29035-0.

Yu:2012:DSA

[923] Hwanjo Yu, Ge Yu, Wynne Hsu, Yang-Sae Moon, Rainer Unland, and Jaesoo Yoo, editors. Database Systems for Advanced Applications: 17th International Conference, DASFAA 2012, International Workshops: FlashDB, ITEMS, SNSM, SIM 3, DQDI, Busan, South Korea, April 15–19, 2012. Proceedings, volume 7240 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29022-1 (print), 3-642-29023-X (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-29023-7.

Zsok:2012:CEF

[924] Viktória Zsók, Zoltán Horváth, and Rinus Plasmeijer, editors. Central European Functional Programming School: 4th Summer School, CEFP 2011, Budapest, Hungary, June 14–24, 2011, Revised Selected Papers, volume 7241 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32095-3 (print), 3-642-32096-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
32096-5.

Caseli:2012:CPP

[925] Helena Caseli, Aline Villavicencio, António Teixeira, and Fernando Perdigão, editors. Computational Processing of the Portuguese Language: 10th International Conference, PROPOR 2012, Coimbra, Portugal, April 17–20, 2012. Proceedings, volume 7243 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28884-7 (print), 3-642-28885-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-28885-2.

Moraglio:2012:GPE

[926] Alberto Moraglio, Sara Silva, Krzysztof Krawiec, Penousal Machado, and Carlos Cotta, editors. Genetic Programming: 15th European Conference, EuroGP 2012, Málaga, Spain, April 11–13, 2012. Proceedings, volume 7244 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29138-4 (print), 3-642-29139-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
29139-5.

Hao:2012:ECC

[927] Jin-Kao Hao and Martin Middendorf, editors. Evolutionary Computation in Combinatorial Optimization: 12th European Conference, EvoCOP 2012, Málaga, Spain, April 11–13, 2012. Proceedings, volume 7245 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29123-6 (print), 3-642-29124-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-29124-1.

Giacobini:2012:ECM

[928] Mario Giacobini, Leonardo Vanneschi, and William S. Bush, editors. Evolutionary Computation, Machine Learning and Data Mining in Bioinformatics: 10th European Conference, EvoBIO 2012, Málaga, Spain, April 11–13, 2012. Proceedings, volume 7246 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29065-5 (print), 3-642-29066-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-29066-4.

Machado:2012:EBI

[929] Penousal Machado, Juan Romero, and Adrian Carballal, editors. Evolutionary and Biologically Inspired Music, Sound, Art and Design: First International Conference, EvoMUSART 2012, Málaga, Spain, April 11– 13, 2012. Proceedings, volume 7247 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29141-4 (print), 3-642-29142-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-29142-5.

DiChio:2012:AEC

[930] Cecilia Di Chio, Alexandros Agapitos, Stefano Cagnoni, Carlos Cotta, Francisco Fernández de Vega, Gianni A. Di Caro, Rolf Drechsler, Anikó Ekárt, Anna I. Esparcia-Alcázar, Muddassar Farooq, William B. Langdon, Juan J. Merelo-Guervós, Mike Preuss, Hendrik Richter, Sara Silva, Anabela Simões, Giovanni Squillero, Ernesto Tarantino, Andrea G. B. Tettamanzi, Julian Togelius, Neil Urquhart, A. Şima Uyar, and Georgios N. Yannakakis, editors. Applications of Evolutionary Computation: EvoApplications 2012: EvoCOMNET, EvoCOMPLEX, EvoFIN, EvoGAMES, EvoHOT, EvoIASP, EvoNUM, EvoPAR, EvoRISK, EvoSTIM, and EvoS-TOC, Málaga, Spain, April 11–13, 2012, Proceedings, volume 7248 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29177-5 (print), 3-642-29178-REFERENCES 180

3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-29178-4.